### v3

# On uniqueness in structured model learning

Abstract
This paper addresses the problem of uniqueness in learning physical laws for systems of partial differential equations (PDEs). Contrary to most existing approaches, it considers a framework of structured model learning, where existing, approximately correct physical models are augmented with components that are learned from data. The main results of the paper are a uniqueness and a convergence result that cover a large class of PDEs and a suitable class of neural networks used for approximating the unknown model components. The uniqueness result shows that, in the limit of full, noiseless measurements, a unique identification of the unknown model components as functions is possible as classical regularization-minimizing solutions of the PDE system. This result is complemented by a convergence result showing that model components learned as parameterized neural networks from incomplete, noisy measurements approximate the regularization-minimizing solutions of the PDE system in the limit. These results are possible under specific properties of the approximating neural networks and due to a dedicated choice of regularization. With this, a practical contribution of this analytic paper is to provide a class of model learning frameworks different to standard settings where uniqueness can be expected in the limit of full measurements.
Keywords: Model learning, partial differential equations, neural networks, unique identifiability, inverse problems.
1 Introduction
Learning nonlinear differential equation based models from data is a highly active field of research. Its general goal is to gain information on a (partially) unknown differential-equation-based physical model from measurements of its state. Information on the model here means to either directly learn a parametrized version of the model or to learn a corresponding parametrized solution map. In both cases, neural networks are used as parametrized approximation classes in most of the existing recent works. Important examples, reviewed in [14], are physics informed neural operators [73], DeepONets [84], Fourier Neural Operators [77], Graph Neural Networks [78], Wavelet Neural Operators [108], DeepGreen [44] and model reduction [12], among others. The comprehensive reviews [7, 13, 16, 28, 74, 89] and the references therein, provide an overview of the state of the art.
Scope.
The above works all focus on full model learning, i.e., learning the entire differential-equation-based model from data. In contrast to this, the approach considered here is focused on structured model learning, where we assume that an approximately correct physical model is available, and only extensions of the model (corresponding to fine-scale hidden physics not present in the approximate model) are learned from data. Specifically, we are concerned with the problem of identifying an unknown nonlinear term $f$ together with physical parameters $\varphi$ of a system of partial differential equations (PDEs)
$\partial_{t}u=F(t,u,\varphi)+f(t,u),\qquad(t,x)\in(0,T)\times\Omega,$ (1)
from indirect, noisy measurements of the state $u$. Here, $T>0$, $\Omega$ is a domain, $F$ is the known physical model and all involved quantities can potentially be vector valued such that systems of PDEs are covered. Also note that the terms $F$ and $f$ can act on values and higher order derivatives of the state. Given this, even though we focus on non-trivial physical models $F$, our work covers also the setting of full model learning by setting $F(t,u,\varphi)=0$.
The main question considered in this work is to what extent measurements $Ku^{l}$ of system states $u^{l}$ corresponding to (unknown) parameters $\varphi^{l}$, $l=1,\ldots,L$, allow to uniquely identify the nonlinearity $f$. Already in the simple setting that $f$ acts pointwise, i.e., $f(\cdot,u)(t,x)=f(u(t,x))$, it is clear that, without further specification, this question only has a trivial answer: Even if $(u^{l},\varphi^{l})_{l}$ is known entirely, $f$ is only determined on $\bigcup_{l=1}^{L}\{u^{l}(t,x)\,|\,(t,x)\in(0,T)\times\Omega\}$.
A natural way to overcome this, as done in [102] (and [103]) for full model learning, is to consider particular types of functions $f$: Specifying to the case $F(t,u,\varphi)=0$, a result of [102] is that a linear or algebraic function $f$ is uniquely identifiable from full state measurements if and only if the state variables (and their derivatives in case $f$ acts also on derivatives) are linearly or algebraically independent, respectively. Similarly, [102] shows that a smooth $f$ is uniquely reconstructable from full state measurements if the values of the state variables (and their derivatives) are dense in the underlying Euclidean vector space. Consistent with this result, [104] shows that equation discovery typically requires chaotic behavior. While these results provide answers in rather general settings, the conditions on $u$ that guarantee unique recovery are difficult to verify exactly in practice ([102] provides an SVD-based algorithm that classifies unique identifiability via thresholding).
A different possibility to address the uniqueness problem would be to consider a specific parametrized class of functions $\{f_{\theta}\,|\,\theta\in\Theta\}$ for approximating $f$, and to investigate uniqueness of the parameters. In case of simple approximation classes such as polynomials, this would indeed provide a simple solution (e.g., parameters of a $n$-degree polynomial are uniquely determined by $n+1$ different values of the state). In case of more complex approximation classes such as neural networks however, this even introduces an additional difficulty, namely that different sets of parameters might represent the same function.
The approach we take in this work to address the uniqueness problem in model learning follows classical inverse-problems techniques for unique parameter identification via regularization-minimizing solutions. Specifically, covering also the setting of non-trivial physical $F$, additional, unknown parameters $(\varphi^{l})_{l}$ and non-trivial forward models, we consider uniqueness of the function $f$ (and the corresponding parameters $\varphi=(\varphi^{l})_{l}$ and states $u=(u^{l})_{l}$) as solutions to the full measurement/vanishing noise limit problem
$\min_{\varphi,u,f}\mathcal{R}^{\dagger}(\varphi,u,f)\qquad\text{s.t. }\forall l:\quad\partial_{t}u^{l}=F(t,u^{l},\varphi^{l})+f(t,u^{l}),\quad K^{\dagger}u^{l}=\hat{y}^{l}$ ($p^{\dagger}$)
where $K^{\dagger}$ is the injective full measurement operator and $y=(\hat{y}^{l})_{l}$ is the corresponding full-measurement data. With this, we allow $\mathcal{R}^{\dagger}$ to incorporate prior information on $f$ that can be used to resolve unique identifiability. In view of the above-described alternative works on uniqueness in model learning, this is related to considering $f$ to be out of a class of functions for which the measurements are sufficient for unique identifiability, only that we do not explicitly rely on such as setting, but rather provide a general framework that will always approximate a desired solution $f^{\dagger}$ as unique regularization-minimizing solution consistent with the measurement data. Of course, if the measurement data is sufficient to uniquely identify a ground-truth $f$ without the use of any additional prior information, our setting will recover this ground-truth. In addition to the question of recovering $(\varphi^{\dagger},u^{\dagger},f^{\dagger})$ as unique solution to ($p^{\dagger}$), it is necessary to analyze in what sense parametrized solutions $(\varphi,u,f_{\theta})$ of the regularized problem
$\min_{\varphi,u,\theta}\mathcal{R}_{m}(\varphi,u,\theta)+\sum_{l=1}^{L}(\lambda^{m}\|\partial_{t}u^{l}-F(t,u^{l},\varphi^{l})-f_{\theta}(t,u^{l})\|^{q}+\mu^{m}\|K^{m}u^{l}-y^{m,l}\|^{r})$ ($p^{m}$)
converge to solutions of ($p^{\dagger}$) for some $1\leq q,r<\infty$. Here, $(K^{m})_{m}$ is a sequence of measurement operators suitably approaching $K^{\dagger}$, $(y^{m,l})_{m}$ with $y^{m,l}\approx K^{m}u^{\dagger,l}$ is a sequence of (noisy) measured data and $\lambda^{m},\mu^{m}>0$ are regularization parameters. More concretely, we suppose the measured data $(y^{m,l})_{m}$ to fulfill the noise estimation given by
$\|y^{m,l}-K^{m}u^{\dagger,l}\|\leq\delta(m)$ (2)
such that $\delta(m)\to 0$ as $m\to\infty$. This is in fact our only requirement on the noise model, i.e., our analytic results do not require assumptions on the noise distribution nor the nature of the noise such as homo-/heteroscedasticity. Note further that although (2) anticipates comparable noise levels on the different measurements $y^{m,l}$ for $l=1,\dots,L$, an extension to different noise levels for different measurements $l$ is straightforward by introducing different weightings in ($p^{m}$) for the data fidelities.
In order to obtain our convergence- and uniqueness results, a suitable regularity of $f$, approximation properties of the parametrized approximation class $\mathcal{F}=\{f_{\theta}|\,\theta\in\Theta\}$ (such as neural networks) as well as a suitable choice of the regularization functionals $\mathcal{R}_{m}$ and $\mathcal{R}^{\dagger}$ are necessary. It turns out from our analysis that the class of locally $W^{1,\infty}$-regular functions is suitable for $f$ and that parameter-growth estimates and local $W^{1,\infty}$ approximation capacities are required for $\mathcal{F}$. We refer to Assumption 5, iv) below for precise requirements on $\mathcal{F}$ which are, as we argue in our work, satisfied for example by certain classes of neural networks. Regarding the regularization functionals, a suitable choice is
$\mathcal{R}_{m}(\varphi,u,\theta)$ $=\mathcal{R}_{0}(\varphi,u)+\|f_{\theta}\|_{L^{\rho}}^{\rho}+\|\nabla f_{\theta}\|_{L^{\infty}}+\nu^{m}\|\theta\|,$ (3)
$\mathcal{R}^{\dagger}(\varphi,u,f)$ $=\mathcal{R}_{0}(\varphi,u)+\|f\|_{L^{\rho}}^{\rho}+\|\nabla f\|_{L^{\infty}},$
with the parameters $\nu^{m}$ appropriately converging to zero as $m\rightarrow\infty$ and $1<\rho<\infty$. Here, the norms $\|\cdot\|_{L^{\rho}}^{\rho}+\|\nabla(\cdot)\|_{L^{\infty}}$ (as opposed to, e.g., a standard $L^{p}$ norm) are necessary to ensure convergence of $f_{\theta}$ to $f$ as functions in $W^{1,\infty}$, which in turn is necessary for convergence of the PDE model. We showcase the necessity of choosing this $W^{1,\infty}$-type norm in contrast to only using the $\|\cdot\|_{L^{\rho}}^{\rho}$-norm in an example in Appendix E: In this example, it is possible to recover the underlying unique hidden physics when using the suggested $W^{1,\infty}$-type regularization approach, while a standard $\|\cdot\|_{L^{\rho}}^{\rho}$-type approach fails to do so.
The norm $\|\theta\|$ on the finite dimensional parameters $\theta$ in (3) is necessary for well-posedness of ($p^{m}$), but will vanish in the limit as $m\rightarrow\infty$. The choice $1<\rho<\infty$ is necessary for ensuring uniqueness of a regularization-minimizing solution ($p^{\dagger}$) via strict convexity, and $\mathcal{R}_{0}(\varphi,u)$ can be any problem-dependent regularization. Note that here, the main ingredients for obtaining uniqueness are that $f$ is no longer parametrized by $\theta$ in the limit (e.g. can be any $W^{1,\infty}$ function) together with a classical strict convexity argument.
While our results on using $W^{1,\infty}$ regularization are formulated for the structured-model-learning-setting here, we note that similar requirements and results can also be expected when learning entire PDE models e.g. with neural operators.
An important question from the computational perspective is how the $W^{1,\infty}$-norm can be approximated in practice or whether there exist scalable surrogates. In fact, a direct way to estimate the $\|\nabla(\cdot)\|_{L^{\infty}}$ term is to draw a certain number of uniformly random samples of the gradient over the considered domain and determine a global extremum over the samples. Aside from this direct approach, this question has been broadly considered in literature in the context of ensuring Lipschitz-stability of neural networks, see for example [76], where the Lipschitz constant of neural networks is estimated via sparse polynomial optimization using linear or semidefinite programming, [59], where the local Lipschitz constant of ReLU networks is computed exactly using mixed-integer programming and by providing upper bound in case of an early stop of the algorithm, [45] where a regularization of the Lipschitz constant is carried out by computing upper bounds during the training process using layerwise constants, [55] which applies a least squares regression algorithm for estimating the Lipschitz constant and also provides lower bounds on the sample complexity of the underlying problem and [41], where the problem of estimating the Lipschitz constant is interpreted as a semidefinite program. See also [19] which considers variational regularization for controlling the Lipschitz constant of a neural network.
Contributions. Following the above concept, we provide a comprehensive analysis of structured model learning in a general setting. Our main contribution is a precise mathematical setup under which we prove the above-mentioned uniqueness and approximation results. Notably, this setup differs from standard model-learning frameworks commonly used in practice, in particular with respect to the choice of regularization for the approximating functions. In view of this, a practical consequence of our work can be a suggestion of appropriate regularization functionals for model learning that ensure unique recovery in the full-measurement/ vanishing noise limit. Indeed, as an example in Appendix E shows, without appropriate regularization a unique recovery in the limit may fail. Besides our main uniqueness result and the corresponding general framework to which it applies, we provide a well-posedness analysis and concrete examples to which our results apply. The latter includes linear and nonlinear (in the state) examples for the physical term $F$ as well as classes of neural networks for $\mathcal{F}$ to which our assumptions apply.
The following proposition, which is a consequence of Proposition 29 and Theorem 30 below, showcases our main results for a specific, linear example.
Proposition 1.
Let the space setup be given by the state space $V=H^{1}(\Omega)$, the image space $W=L^{2}(\Omega)$, the measurement space $Y=L^{2}(\Omega)$ and parameter space $X_{\varphi}=H^{1}(\Omega)$ for a bounded interval $\Omega\subseteq\mathbb{R}$ with the time extended spaces
$\mathcal{V}=W^{1,2,2}(0,T;V),\quad\mathcal{W}=L^{2}(0,T;W),\quad\mathcal{Y}=L^{2}(0,T;Y).$
Consider the one dimensional convection equation with unknown reaction term
$\partial_{t}u^{l}=\varphi^{l}\cdot\nabla u^{l}+f(u^{l})$ (4)
where $\varphi^{l}\in X_{\varphi}$ for $1\leq l\leq L$ subject to $K^{\dagger}u^{l}=\hat{y}^{l}$ with $K^{\dagger}:\mathcal{V}\to\mathcal{Y}$ an injective, linear, bounded operator and $(\hat{y}^{l})_{l}\subseteq\mathcal{Y}$ full measurement data. Suppose that there exist admissible $\hat{f}\in W^{1,\infty}(\mathbb{R})$, $\hat{u}\in\mathcal{V}^{L}$ and $\hat{\varphi}\in X_{\varphi}^{L}$ solving (4) such that $K^{\dagger}\hat{u}^{l}=\hat{y}^{l}$ for $1\leq l\leq L$. Assume that $f$ is approximated by neural networks $f_{\theta}$ of the form in [11, Theorem 1] parameterized by $\theta\in\Theta^{m}$ with $m\in\mathbb{N}$ a scale of approximation. Suppose that $(K^{m})_{m}$ is a sequence of bounded linear operators strongly converging to $K^{\dagger}$ and $(y^{m,l})_{m}\subseteq\mathcal{Y}$ a sequence of measurement data converging to $y^{l}$. Assume further that $U\subseteq\mathbb{R}$ is a sufficiently large interval.
Then there exists a unique solution $(\varphi^{\dagger},u^{\dagger},f^{\dagger})$ to the vanishing noise limit problem
$\min_{\begin{subarray}{c}\varphi\in H^{1}(\Omega)^{L},u\in\mathcal{V}^{L},\\ f\in W^{1,\infty}(U)\end{subarray}}\sum_{l=1}^{L}(\|\varphi^{l}\|_{H^{1}(\Omega)}^{2}+\|u^{l}\|_{\mathcal{V}}^{2})+\|f\|_{L^{2}(U)}^{2}+\|\nabla f\|_{L^{\infty}(U)}$ ($p^{\dagger}$)
$\text{s.t. }\forall l:\quad\partial_{t}u^{l}=\varphi^{l}\cdot\nabla u^{l}+f(u^{l}),\quad K^{\dagger}u^{l}=\hat{y}^{l}.$
Furthermore, for $\lambda^{m},\mu^{m}\to\infty,\nu^{m}\to 0$ as $m\to\infty$ at certain rate depending on the neural network architectures and $(y^{m,l})_{m}$, let $(\varphi_{m},u_{m},\theta_{m})$ be a solution to
$\min_{\varphi\in H^{1}(\Omega)^{L},u\in\mathcal{V}^{L},\theta\in\Theta^{m}}\sum_{l=1}^{L}(\|\varphi^{l}\|_{H^{1}(\Omega)}^{2}+\|u^{l}\|_{\mathcal{V}}^{2})+\|f_{\theta}\|_{L^{2}(U)}^{2}+\|\nabla f_{\theta}\|_{L^{\infty}(U)}$ ($p^{m}$)
$\qquad+\nu^{m}\|\theta\|+\sum_{l=1}^{L}(\lambda^{m}\|\partial_{t}u^{l}-\varphi^{l}\cdot\nabla u^{l}-f_{\theta}(u^{l})\|^{2}_{\mathcal{W}}+\mu^{m}\|K^{m}u^{l}-y^{m,l}\|^{2}_{\mathcal{Y}})$
for each $m\in\mathbb{N}$. Then if $f^{\dagger}\in\mathcal{C}^{1}(U)$ it holds true that $\varphi_{m}\rightharpoonup\varphi^{\dagger}$ in $H^{1}(\Omega)^{L}$, $u_{m}\rightharpoonup u^{\dagger}$ in $\mathcal{V}^{L}$ and $f_{\theta_{m}}\overset{*}{\rightharpoonup}f^{\dagger}$ in $W^{1,\infty}(U)$.
Proof.
See Appendix D. ∎
It is important to emphasize that, among the assumptions stated in Proposition 1, the most restrictive one from a practical perspective is the one on existence of an admissible solution $\hat{u}$ with sufficiently high regularity. This can be viewed as an implicit assumption on the existence of a sufficiently regular transport field $\hat{\varphi}$ that is compatible with $\hat{u}$. Indeed, the regularity of $\hat{u}$ as a solution to the transport equation (4) depends not only on the source term $\hat{f}$, which is essentially Lipschitz continuous, but crucially on the regularity of the transport field $\hat{\varphi}$. The relationship between the smoothness of the transport field and the well-posedness and regularity of solutions has been extensively studied. The foundational work [30] establishes well-posedness for transport equations when the transport field has Sobolev regularity, connecting the regularity of solutions to that of the flow generated by the field. A more recent survey of these results is given in [5]. Further developments, including [17, 25, 26, 27, 86], study how the smoothness of the transport field affects the stability and regularity of solutions, and demonstrate that, at critical levels of regularity, the solutions may lose uniqueness and smoothness. In fact, for Lipschitz continuous $\hat{\varphi}$ and $\hat{f}$ together with initial condition $u_{0}\in H^{2}(0,1)$ one can show, following [17, Theorem 1.2], that $u\in W^{1,2,2}(0,T;H^{1}(0,1))$. Thus, the regularity assumption on the admissible state above can be interpreted as a regularity assumption on the parameters $\hat{\varphi}$, $\hat{f}$ and $\hat{u}_{0}$.
Related works. This work is mainly motivated by [1] on data-driven structured model learning which proposes an all-at-once approach for learning-informed parameter identification, i.e., determining the state simultaneously with the nonlinearity and the input parameters. Note that [1] considers single PDEs, while our work generalizes to PDE systems where the unknown term may additionally depend on higher order derivatives of the state variable. Besides this fundamental difference, we derive wellposedness of the learning problem under slightly different conditions, where higher regularity assumptions on the state space stated in [1] can be omitted if the activation function of the neural networks approximating the nonlinearities is globally Lipschitz continuous. Moreover, we treat the cases of linear and nonlinear physical terms separately. Finally, the main difference of our work to [1] is that we focus on unique reconstructability, whereas [1] is mostly focused on well-posedness of the learning problem and the resulting PDE.
The main reason for choosing an all-at-once approach (see e.g. [63, 64]) in general is the possibility to account for practically realistic, incomplete and indirectly measured state data, which may be polluted by noise. It also circumvents the use of the parameter-to-state map, which requires regularity conditions that may not be feasible in practice (see e.g. [50, 61, 62, 91]).
In contrast to the all-at-once setting pursued here, works that use a learning-informed control-to-state map to study the optimal control of certain PDEs are [32, 33, 34]. There it is assumed that the nonlinear constituents are only accessible through data-driven techniques e.g. arising from neural networks. Another related work in the field of optimal control is [24] on nonlinearity identification in the monodomain model via neural network parameterization. We also mention the recent paper [22] which deals with the identification of semilinear elliptic PDEs in a low-regularity control regime. In the context of approximating nonlinearities for elliptic state equations see [105]. We also mention the recent work [96] which establishes global convergence guarantees for adjoint-based training of infinite-width neural networks embedded in nonlinear parabolic PDEs. For structured model learning for ODEs we refer to [36, 46]. See also [90] on regularized inversion for hidden reaction law discovery.
Recent work incorporates conservation laws and symmetries into machine-learning models to improve physical fidelity and data efficiency. Soft-constraint methods enforce approximate conservation through regularized loss terms [57, 79, 110], whereas exact-conservation approaches leverage integral forms [51], or apply adaptive correction mechanisms [21, 43, 80]. Complementary strategies embed conservation directly into the architecture, via hard constraints in output layers [107], projection onto admissible solution spaces [93], or symmetry encoding [87], with additional architectural designs explored in [54, 81, 82, 95].
An important aspect of model learning is interpretability, which seeks representations that are accurate, parsimonious, and physically consistent. A primary goal is to recover simple laws that faithfully describe the underlying data [6, 94, 99, 101]. For a comprehensive overview of related methods, see [75].
From the perspective of inverse problems, model learning is, at its core, an identification problem, requiring that the inferred model is uniquely determined by the data to represent the true system dynamics rather than an equivalent alternative. Foundational contributions [10, 23, 31, 85] formalize structural and parameter identifiability in dynamical systems. For the motivation of uniqueness results for parameter identification, we refer to the works [20, 37, 97], which derive uniqueness from stability estimates. Uniqueness has also been established for semilinear parabolic equations [56] and for the recovery of nonlinear diffusion coefficients [67]. Further results on parameter identification for elliptic equations include [2, 4, 72]. Foundational overviews of parameter identification and related inverse problems are provided in [39] on deterministic regularization theory, [71] on coefficient inverse problems, [8] on PDE parameter estimation, and [60] on Bayesian inverse problems. Beyond coefficients, model identification targets entirely unknown PDE components, with applications to reaction-diffusion systems [35, 65, 66, 68], semilinear equations [42, 69, 70] and hyperbolic inverse sources [58, 111].
Nonetheless, there is little hope to obtain results of this kind for the general system (1), even if the known physical term is linear in its physical input parameters due to the ambiguity of shift perturbations. In this respect, it seems indispensable to exploit the structural/regularity properties of the unknown term $f$ and the input parameter $\varphi$, as it is in this work and in [102], which was already discussed above. For the sake of completeness we also mention the recent preprint [52], extending the results of [102] on identifiability for symbolic recovery of differential equations to the noisy regime. Note that both works [52, 102] focus on unique identifiability per se, i.e. the classification of uniqueness, whereas our work provides an analysis-based guideline guaranteeing unique reconstructability in the limit of a practical PDE-based model learning setup.
Structure of the paper.
In Section 2 we present the problem setting under consideration. The necessary assumptions are outlined in detail in Subsection 2.1. In Subsection 2.2, applicability of our general assumptions for $\mathcal{F}$ being a certain class of neural networks are discussed. Applicability of the assumptions on the known physical term are discussed in Subsection 2.3, with examples both for the linear and nonlinear case. Our main result on unique reconstructability in the limit problem is presented in Section 3. To ensure a concise presentation of our results, most proofs are covered in the appendix. The results of Subsection 2.2 are proven in Appendix A and those of Subsection 2.3 are given in Appendix B. In Appendix C wellposedness of the main minimization problem is verified under our general assumptions. In Appendix D a proof of Proposition 1, showcasing our main results for a specific, linear example, is sketched. Finally, in Appendix E an example is presented covering the necessity of the proposed regularization for unique recovery in the limit.
2 Problem setting
In the general case, we are interested in obtaining nonlinearities $(f_{n})_{n}$, states $(u^{l}_{n})_{n,l}$, parameters $(\varphi^{l}_{n})_{n,l}$, initial conditions $(u_{0,n}^{l})_{n,l}$ and boundary conditions $(g^{l}_{n})_{n,l}$ as solutions of the following system of nonlinear PDEs:
$\frac{\partial}{\partial t}u_{n}^{l}$ $=F_{n}(t,u^{l}_{1},\dots,u^{l}_{N},\varphi_{n}^{l})+f_{n}(t,\mathcal{J}_{\kappa}u^{l}_{1},\dots,\mathcal{J}_{\kappa}u^{l}_{N}),$ ($S$)
$u^{l}_{n}(0)$ $=u^{l}_{0,n},$
$\gamma(u^{l}_{n})$ $=g_{n}^{l}$
Here, $n=1,\ldots,N$ denotes the number of PDEs and $l=1,\ldots,L$ the number of measurements of different states (with different parameters) that we will have at our disposal for obtaining the $f_{n}$.
In the above system, the states $u^{l}_{n}\in\mathcal{V}$ are given as $u^{l}_{n}:(0,T)\rightarrow V$ with $T>0$ and $V$ a static state space of functions $v:\Omega\rightarrow\mathbb{R}$ with $d\in\mathbb{N}$ and $\Omega\subset\mathbb{R}^{d}$ a bounded Lipschitz domain, $X_{\varphi}\ni\varphi_{n}^{l}$ is a static parameter space, $H\ni u^{l}_{n}(0),u^{l}_{0,n}$ is a static initial trace space, and $\mathcal{B}\ni g_{n}^{l}$ is a boundary trace space with $g_{n}^{l}:(0,T)\rightarrow B$, $B$ the static boundary trace space and $\gamma:\mathcal{V}\rightarrow\mathcal{B}$ the boundary trace map. The (known) physical terms $F_{n}$ are given as Nemytskii operators of
$F_{n}:(0,T)\times V^{N}\times X_{\varphi}$ $\to W$ (5)
$(t,u_{1},\dots,u_{N},\varphi)$ $\mapsto F_{n}(t,u_{1},\dots,u_{N},\varphi)$
with $W$ a static image space and $\mathcal{W}$ the corresponding dynamic version. The $\mathcal{J}_{\kappa}$ are derivative operators given as
$\mathcal{J}_{\kappa}:V$ $\to\otimes_{k=0}^{\kappa}V_{k}^{\times}$ (6)
$v$ $\mapsto(v,J^{1}v,\dots,J^{\kappa}v)$
with the Jacobian mappings $J^{k}$ given as
$J^{k}:V\to V_{k}^{\times},~~v\mapsto(D^{\beta}v)_{|\beta|=k}.$ (7)
[…]
$f_{n}:(0,T)\times(\otimes_{k=0}^{\kappa}V_{k}^{\times})^{N}$ $\to W$
$(t,(v_{1}^{k})_{0\leq k\leq\kappa},\dots,(v_{N}^{k})_{0\leq k\leq\kappa})$ $\mapsto f_{n}(t,(v_{1}^{k})_{0\leq k\leq\kappa},\dots,(v_{N}^{k})_{0\leq k\leq\kappa})$
where $f_{n}:(0,T)\times(\otimes_{k=0}^{\kappa}\mathbb{R}^{p_{k}})^{N}\rightarrow\mathbb{R}$ is extended to $f_{n}:(0,T)\times(\otimes_{k=0}^{\kappa}V_{k}^{\times})^{N}\to W$ via $f_{n}(t,v)(x):=f_{n}(t,v(x))$. We will approximate them with parameterized approximation classes
$\mathcal{F}_{n}^{m}=\{f_{\theta_{n},n}:(0,T)\times(\otimes_{k=0}^{\kappa}\mathbb{R}^{p_{k}})^{N}\to\mathbb{R}~|~\theta_{n}\in\Theta_{n}^{m}\}$ (8)
where $m\in\mathbb{N}$ is the scale of approximation and $\Theta_{n}^{m}$ are parameter sets. Here, we further define $\Theta^{m}=\otimes_{n=1}^{N}\Theta_{n}^{m}$ and $\mathcal{F}^{m}=\otimes_{n=1}^{N}\mathcal{F}_{n}^{m}$.
Approximation of the $f_{n}$ via the $f_{\theta_{n},n}$ will be achieved on the basis of noisy measurements $y^{l}\approx K^{m}u^{l}$, with the $K^{m}:\mathcal{V}^{N}\rightarrow\mathcal{Y}$ being measurement operators (for scale $m\in\mathbb{N}$) and $\mathcal{Y}$ a space of functions $y:(0,T)\rightarrow Y$ with $Y$ a static measurement space. To this aim, we will analyze the following minimization problem
$\min_{\begin{subarray}{c}\varphi\in X_{\varphi}^{N\times L},\theta\in\Theta^{m},\\ u\in\mathcal{V}^{N\times L},u_{0}\in H^{N\times L},\\ g\in\mathcal{B}^{N\times L}\end{subarray}}\sum_{1\leq l\leq L}\lambda\|\frac{\partial}{\partial t}u^{l}-F(t,u^{l},\varphi^{l})-f_{\theta}(t,\mathcal{J}_{\kappa}u^{l})\|_{\mathcal{W}}^{q}+\mathcal{R}(\varphi,u,\theta,u_{0},g)$
$+\sum_{1\leq l\leq L}\bigg[\lambda\|u^{l}(0)-u_{0}^{l}\|_{H}^{2}+\lambda\mathcal{D}_{\text{BC}}(\gamma(u^{l})-g^{l})+\mu\|K^{m}u^{l}-y^{l}\|_{\mathcal{Y}}^{r}\bigg]$ ($\mathcal{P}$)
where $\mathcal{D}_{\text{BC}}$ and $\mathcal{R}$ are suitable discrepancy and regularization functionals, respectively. Note that here, notation wise, we use a direct vectorial extension over $n=1,\ldots,N$ of all involved spaces and quantities, e.g., $F(t,u^{l},\varphi^{l})=(F_{n}(t,u^{l},\varphi^{l}_{n}))_{n=1}^{N}$.
2.1 Assumptions
The following assumptions, motivated by [1, Assumption 1], encompass all requirements necessary to tackle the goals of this work. Under Assumption 2, 3 and 4 we verify wellposedness of (2). Additionally, under Assumption 5, we will establish our results on unique reconstructability in the limit $m\rightarrow\infty$.
Assumption 2 (Functional analytic setup).
Spaces/Embeddings:
  1. i)
    For $\kappa\in\mathbb{N}$, suppose that the state space $V$, the spaces $V_{k}$ for $1\leq k\leq\kappa$, the image space $W$, the observation space $Y$, the initial trace space $H$, the boundary trace space $B$ and the space $\tilde{V}$ are separable, reflexive Banach spaces. Further assume that the parameter space $X_{\varphi}$ is a reflexive Banach space and let $\Theta^{m}_{n}$, for $n=1,\ldots,N$ and $m\in\mathbb{N}$ be closed parameter sets, each contained in a finite-dimensional space.
  2. ii)
    Let $\Omega\subset\mathbb{R}^{d}$ with $d\in\mathbb{N}$ be a bounded Lipschitz domain and assume the following embeddings to hold:
[…]
    $L^{\hat{p}}(\Omega)\hookrightarrow V_{k}\hookrightarrow L^{\hat{q}}(\Omega)~\text{for}~1\leq k\leq\kappa,~~V\hookrightarrow Y,~~L^{\hat{q}}(\Omega)\hookrightarrow W$
    and either $W^{\kappa,\hat{p}}(\Omega)\hookrightarrow\tilde{V}$ or $\tilde{V}\hookrightarrow W^{\kappa,\hat{p}}(\Omega)$ for some $1\leq\hat{q}\leq\hat{p}$$\infty$.
  3. iii)
    Let $T>0$ and the extended spaces be defined by $\mathcal{W}=L^{q}(0,T;W),$
    $\mathcal{V}=L^{p}(0,T;V)\cap W^{1,p,p}(0,T;\tilde{V}),\mathcal{Y}=L^{r}(0,T;Y),\mathcal{B}=L^{s}(0,T;B),$
    $\mathcal{V}_{0}=\mathcal{V}_{0}^{\times}:=\mathcal{V},\mathcal{V}_{k}=L^{p}(0,T;V_{k}),~~\mathcal{V}_{k}^{\times}=L^{p}(0,T;V_{k}^{\times})~~\text{for}~1\leq k\leq\kappa$
    for some $1\leq p,q,r,s<\infty$ with $p\geq q$, $p\geq s$. We refer to [98, Chapter 7] for the definition and properties of (Sobolev-)Bochner spaces.
    Trace map:
  4. iv)
    Assume that the boundary trace map $\gamma:\mathcal{V}\rightarrow\mathcal{B}$ is linear and continuous.
    Measurement operator:
  5. v)
    Suppose that the operator $K^{m}:\mathcal{V}^{N}\to\mathcal{Y}$ is weak-weak continuous for $m\in\mathbb{N}$.
    Energy functionals:
  6. vi)
    Assume that the discrepancy term $\mathcal{D}_{\text{BC}}:\mathcal{B}^{N}\to[0,\infty]$ is weakly lower semicontinuous, coercive and fulfills $\mathcal{D}_{\text{BC}}(z)=0$ iff $z=0$. Suppose that the regularization functional $\mathcal{R}:X_{\varphi}^{N\times L}\times\mathcal{V}^{N\times L}\times\Theta^{m}\times H^{N\times L}\times\mathcal{B}^{N\times L}\to[0,\infty]$ is coercive in its first three components and weakly lower semicontinuous. Further suppose that there exists $(\varphi,u,\theta,u_{0},g)\in\mathbf{D}(\mathcal{R})$ with $(\gamma(u^{l})-g^{l})_{l}\subseteq\mathbf{D}(\mathcal{D}_{\text{BC}})$ where $\mathbf{D}(\mathcal{D}_{\text{BC}})$ and $\mathbf{D}(\mathcal{R})$ denote the domains of the respective functionals.
The next assumption concerns general properties on the parameterized nonlinearities that will be needed for wellposedness.
Assumption 3 (Parameterized approximation classes $(\mathcal{F}^{m}_{n})_{n}$).
Nemytskii operators:
  1. i)
    Assume that $f_{\theta_{n},n}\in\mathcal{F}^{m}_{n}$ with $\mathcal{F}^{m}_{n}$ defined as in (8) induce well-defined Nemytskii operators $f_{\theta_{n},n}:(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}\to\mathcal{W}$ via
    $f_{\theta_{n},n}((v^{k})_{0\leq k\leq\kappa})](t)(x)=f_{\theta_{n},n}(t,(v^{k}(t,x))_{0\leq k\leq\kappa}).\vskip 5.69046pt$
    Strong-weak continuity:
  2. ii)
    Suppose that for each $f_{\theta_{n},n}\in\mathcal{F}^{m}_{n}$ the map
    $\Theta_{n}^{m}\times(\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}}))^{N}\ni(\theta_{n},v)\mapsto f_{\theta_{n},n}(v)\in L^{q}(0,T;L^{\hat{q}}(\Omega))$
    is strongly-weakly continuous.
We require an analogous assumption for the physical PDE-term.
Assumption 4 (Known physical term).
Nemytskii operators:
  1. i)
    Assume that the $F_{n}$ induce well-defined Nemytskii operators
    $F_{n}:\mathcal{V}^{N}\times X_{\varphi}\to\mathcal{W}~~\text{with}~~[F_{n}(v,\varphi)](t)=F_{n}(t,v(t),\varphi).$
    Weak-closedness:
  2. ii)
    Suppose that the $F_{n}:\mathcal{V}^{N}\times X_{\varphi}\to\mathcal{W}$ are weakly closed.
Finally, to obtain our uniqueness results, we need to impose more regularity both on the state space and the approximation class. For that, recall the definition of the differential operator $\mathcal{J}_{\kappa}$ in (6) and note that, as we will show in Lemma 42, it follows from Assumption 2 that the $\mathcal{J}_{\kappa}$ induce suitable Nemytskii operators such that the following assumption makes sense notationally.
Assumption 5 (Uniqueness).
Regularity:
  1. i)
    Assume that there exists a constant $c_{\mathcal{V}}>0$ such that
    $\|\mathcal{J}_{\kappa}v\|_{L^{\infty}((0,T)\times\Omega)}\leq c_{\mathcal{V}}\|v\|_{\mathcal{V}}~~\text{for all}~v\in\mathcal{V}.$
  2. ii)
    For $D=1+N\sum_{k=0}^{\kappa}p_{k}$, $1\leq n\leq N$, $m\in\mathbb{N}$ suppose that $\mathcal{F}_{n}^{m}\subseteq W^{1,\infty}_{\text{loc}}(\mathbb{R}^{D})$.
  3. iii)
    Suppose that the full measurement data $\hat{y}\in\mathcal{Y}^{L}$ is such that there exist admissible functions $\hat{f}\in W^{1,\infty}(\mathbb{R}^{D})^{N}$, $\hat{u}\in\mathcal{V}^{N\times L}$, $\hat{\varphi}\in X_{\varphi}^{N\times L}$, $\hat{u}_{0}\in H^{N\times L}$ and $\hat{g}\in\mathcal{B}^{N\times L}$ solving ($S$) such that $K^{\dagger}\hat{u}^{l}=\hat{y}^{l}$ for all $l=1,\ldots,L$.
    Approximation capacity of $\mathcal{F}^{m}$ for $f\in W^{1,\infty}_{\text{loc}}(\mathbb{R}^{D})^{N}$:
  4. iv)
    The approximation capacity condition is considered to be satisfied for a fixed $f\in W^{1,\infty}_{\text{loc}}(\mathbb{R}^{D})^{N}$ if for any bounded domain $U\subseteq\mathbb{R}^{D}$ there exist a monotonically increasing $\psi:\mathbb{N}\to\mathbb{R}$ and $c,\beta>0$ such that for $\|\cdot\|$ denoting some $l^{p}$-Norm for $1\leq p\leq\infty$ there exist parameters $\theta^{m}\in\Theta^{m}$ with
    $\|f-f_{\theta^{m}}\|_{L^{\infty}(U)}\leq cm^{-\beta},\qquad\|\theta^{m}\|\leq\psi(m)$ (9)
    and $\|\nabla f_{\theta^{m}}\|_{L^{\infty}(U)}\to\|\nabla f\|_{L^{\infty}(U)}$ as $m\to\infty$.
    Measurement operator:
  5. v)
    Suppose that for any weakly convergent sequence $(u^{m})_{m}\subset\mathcal{V}^{N}$ it holds true that
    $K^{m}u^{m}-K^{\dagger}u^{m}\to 0\quad\text{in}~\mathcal{Y}\quad\text{as}~m\to\infty.$ (10)
    Assume that $K^{\dagger}$ is injective and weak-strong continuous.
    Regularization functional:
  6. vi)
    Let $\mathcal{R}_{0}:X_{\varphi}^{N\times L}\times\mathcal{V}^{N\times L}\times H^{N\times L}\times\mathcal{B}^{N\times L}\to[0,\infty]$ be strictly convex in its first component. Assume that there exists a monotonically increasing function $\pi:[0,\infty)\to[0,\infty)$ (e.g. the $p$-th root) such that for $v\in\mathcal{V}^{N\times L}$
    $\|v\|_{\mathcal{V}}\leq\pi(\mathcal{R}_{0}(\cdot,v,\cdot,\cdot)).$
    Let $\mathcal{R}:X_{\varphi}^{N\times L}\times\mathcal{V}^{N\times L}\times\otimes_{n}\Theta_{n}^{m}\times H^{N\times L}\times\mathcal{B}^{N\times L}\rightarrow[0,\infty]$ be given as
    $\mathcal{R}(\varphi,u,\theta,u_{0},g)=\mathcal{R}_{0}(\varphi,u,u_{0},g)+\nu\|\theta\|+\|f_{\theta}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f_{\theta}\|_{L^{\infty}(U)},$
    for $1<\rho<\infty$ and $U\subset\mathbb{R}^{D}$ a bounded Lipschitz domain.
  7. vii)
[…]
    Physical term:
  8. viii)
    Suppose that $X_{\varphi}\ni\varphi\mapsto F(t,u,\varphi)\in W^{N}$ is affine for $u\in V^{N}$ and $t\in(0,T)$. Assume that $F:\mathcal{V}^{N}\times X_{\varphi}\to\mathcal{W}^{N}$ is weakly continuous.
The following remarks discuss some aspects of the above assumptions.
Remark 6 (Examples).
In the next two subsections we provide examples of approximation classes $\mathcal{F}_{n}^{m}$ and physical terms $F$ where Assumptions 2 to 5 hold. In particular, we show that Assumption 3 together with ii) and iv) in Assumption 5 hold in case $\mathcal{F}_{n}^{m}$ is chosen as a suitable class of neural networks and $f$ in Assumption 5, iv) has a suitable regularity.
Remark 7 (Compact embedding of state space).
A possible choice of the space $V$ satisfying the compact embedding in Assumption 2 is $V=W^{\kappa+\tilde{\kappa},p_{0}}(\Omega)$ for $1<p_{0}<\infty,\tilde{\kappa}\in\mathbb{N}$ fulfilling either $\tilde{\kappa}p_{0}<d$ with $1\leq\hat{p}<\frac{dp_{0}}{d-\tilde{\kappa}p_{0}}$ or $\tilde{\kappa}p_{0}=d$ with $1\leq\hat{p}$$\infty$ due to the Rellich-Kondrachov Theorem (see e.g. [3, Theorem 6.3] and [40, §5.7]). The spaces $V_{k}$ can be chosen as $V_{k}=L^{\hat{p}}(\Omega)$ for $1\leq k\leq\kappa$.
Remark 8 (Role of operator $\mathcal{J}_{\kappa}$).
As the nonlinearities $f_{\theta_{n},n}$ operate pointwise in space and time, the operator $\mathcal{J}_{\kappa}$ is needed to allow for a dependence of $f_{\theta_{n},n}$ also on derivatives of the state. For the physical term $F$ on the other hand, an explicit incorporation of derivatives is not necessary, as $F$ does not act pointwise in space but rather directly on $V$.
Remark 9 (Regularity condition extended state space).
[…]
If $\tilde{V}$ is sufficiently regular, e.g. fulfills some embedding of the form
$\tilde{V}\hookrightarrow W^{\kappa+\tilde{\kappa},\eta}(\Omega)$ (12)
with $\tilde{\kappa}\eta>d=\dim(\Omega)$, then
$\mathcal{C}(0,T;W^{\tilde{\kappa},\eta}(\Omega))\hookrightarrow L^{\infty}((0,T)\times\Omega).$ (13)
Combining the embeddings (11), (12) and (13) together with $D^{\beta}v(t)\in W^{\tilde{\kappa},\eta}(\Omega)$ for $v\in\mathcal{V}$ and $t\in(0,T)$ yields Assumption 5, i).
Remark 10 (Convergence of measurement operators).
Note that the required convergence in (10) is rather weak in practice. In fact it holds for (potentially nonlinear) operators $(K^{m})_{m}$ converging to $K^{\dagger}$ uniformly on bounded sets in $\mathcal{V}$ (since weakly convergent sequences are bounded by [15, Proposition 3.5 (iii)]. This, in particular encompasses bounded linear operators converging in the operator norm.
Remark 11 (Regularity of admissible function).
The assumption $\hat{f}\in W^{1,\infty}(\mathbb{R}^{D})^{N}$ in Assumption 5, iii), seems to be restrictive. However, since an admissible state $\hat{u}$ attains uniformly bounded $\mathcal{J}_{\kappa}\hat{u}$ by Assumption 5, i), the term $\hat{f}$ only acts on a compact subset of $\mathbb{R}^{D}$ in ($S$) due to its composition to $\mathcal{J}_{\kappa}\hat{u}$. Thus, whenever a $\hat{f}\in W^{1,\infty}_{loc}(\mathbb{R}^{D})^{N}$ solving ($S$) exists, there exists w.l.o.g. a solution which is globally $W^{1,\infty}(\mathbb{R}^{D})^{N}$-regular as a consequence of the following extension argument. For a sufficiently large and regular subset $U\subset\mathbb{R}^{D}$ meeting the regularity conditions in the references below and containing $(t,\mathcal{J}_{\kappa}\hat{u}(t,x))$ for a.e. $(t,x)\in(0,T)\times\Omega$ define $\hat{f}_{0}:\mathbb{R}^{D}\to\mathbb{R}^{N}$ with $\hat{f}_{0}=\hat{f}$ on $U$. The function $\hat{f}_{0}\in W^{1,\infty}(U)^{N}$ is then extendable to some $\hat{f}_{0}\in W^{1,\infty}(\mathbb{R}^{D})^{N}$ due to regularity of $U$. The result in [106, Chapter VI, Theorem 5] treats this in a more general framework that includes general Sobolev spaces and minimal smoothness conditions on the domain $U$. We refer to [3, Theorem 5.24] for an outline of the proof. A proof of the extension result, but for first-order Sobolev spaces and stricter smoothness assumptions on the underlying domain, can be found in [15, Theorem 9.7].
Remark 12 (Regularity of admissible state).
The existence of an admissible solution to ($S$) as required in Assumption 5, iii), with state regularity $\mathcal{V}$ may in practice be difficult to guarantee in view of the regularity typically expected from the underlying equation. For a concrete example we refer to the discussion of the transport equation following Proposition 1 in the introduction. Nevertheless, it is important to note that this regularity requirement can be interpreted as an implicit assumption on the parameter space $X_{\varphi}$, as briefly highlighted for the transport equation, since the regularity of the state is generally inherited from the regularity of the model and that of the input parameters.
Remark 13 (Choice of $U$).
In view of Assumption 5, vii) one can circumvent choosing a sufficiently large radius $\delta$ depending on all admissible functions as in Assumption 5, iii) as follows. For a closed and convex set $U\subset\mathbb{R}^{D}$ containing $(t,\mathcal{J}_{\kappa}\hat{u}(t,x))$ for a.e. $(t,x)\in(0,T)\times\Omega$ one can define $P_{U}$ to be the metric projector onto $U$. Now considering the term $f_{\theta}(P_{U}(t,\mathcal{J}_{\kappa}u^{l}))$ in (2) instead of $f_{\theta}(t,\mathcal{J}_{\kappa}u^{l})$ the results of this work still apply. Of course well-definedness of $f_{\theta}\circ P_{U}$ in function space has to be argued first. From a model perspective the composition $f_{\theta}\circ P_{U}$ can be interpreted as part of the parameterized approximation classes (8). For the specific case of neural networks this generalizes to applying a sufficiently regular sigmoidal-type function as activation function in the first layer.
Remark 14 (A priori bounded states).
It is possible to circumvent both the assumption $\hat{f}\in W^{1,\infty}(\mathbb{R}^{D})^{N}$ and the regularity condition in Assumption 5, i), if it is a priori known that the $\mathcal{J}_{\kappa}u$ are uniformly bounded.
For instance, in case $\kappa=0$, the state $u$ may model e.g. some chemical concentration which is a priori bounded in the interval $[0,1]$.
Remark 15 (Boundary trace map).
In view of Assumption 2, i) if $V\hookrightarrow W^{\kappa+1,\hat{p}}(\Omega)$, a possible choice of the trace map $\gamma:\mathcal{V}\to\mathcal{B}$ is the (pointwise in time) Dirichlet trace operator $\gamma_{0}:V\rightarrow B$ (see [3, Chapter 5]) with $B=L^{b}(\partial\Omega)$ for $b$ as follows. Following [3, Theorem 5.36] for instance, $\gamma_{0}:W^{\kappa,\hat{p}}(\Omega)\to L^{b}(\partial\Omega)$ (and hence $\gamma$) is weak-weak continuous if $\kappa\hat{p}\leq d$ and $\hat{p}\leq b\leq\frac{(d-1)\hat{p}}{d-\kappa\hat{p}}$ (with $\hat{p}\leq b<\infty$ if $\kappa\hat{p}=d$). The choice of the (pointwise in time) Neumann trace operator (see [88, Chapter 2])) may be treated similarly with the same conditions on $b$.
The discrepancy functional $\mathcal{D}_{\text{BC}}$ can for instance be given as the indicator functional by $\mathcal{D}_{\text{BC}}(w)=0$ if $w=0$ and $\mathcal{D}_{\text{BC}}(w)=\infty$ else, acting as a hard constraint, or as soft constraint via $\mathcal{D}_{\text{BC}}(w)=\sum_{n}\|w_{n}\|_{\mathcal{B}}^{s}$ for $w\in\mathcal{B}^{N}$. In both cases $\mathcal{D}_{\text{BC}}$ is weakly lower semicontinuous, coercive and fulfills $\mathcal{D}_{\text{BC}}(z)=0$ iff $z=0$.
2.2 Neural networks
In this section we discuss Assumption 3 together with ii) of Assumption 5 in case $(\mathcal{F}_{n}^{m})_{n}$ are chosen as suitable classes of feed forward neural networks. Furthermore, we provide results from literature that ensure Assumption 5, iv) for specific network architectures and suitably regular $f$. Moreover, we address also Assumption 5, vi).
Definition 16.
Let $L\in\mathbb{N}$, $(n_{l})_{0\leq l\leq L}\subseteq\mathbb{N}$, $\sigma\in\mathcal{C}(\mathbb{R},\mathbb{R})$ and $\theta_{l}=(w^{l},\beta^{l})$ with $w^{l}\in\mathcal{L}(\mathbb{R}^{n_{l-1}},\mathbb{R}^{n_{l}})\simeq\mathbb{R}^{n_{l}\times n_{l-1}}$ and $\beta^{l}\in\mathbb{R}^{n_{l}}$ for $1\leq l\leq L$. Furthermore, let $L_{\theta_{l}}:\mathbb{R}^{n_{l-1}}\to\mathbb{R}^{n_{l}}$ via $L_{\theta_{l}}(z):=\sigma(w^{l}z+\beta^{l})$ for $1\leq l\leq L-1$ together with $L_{\theta_{L}}(z):=w^{L}z+\beta^{L}$. Then a fully connected feed forward neural network $\mathcal{N}_{\theta}$ with activation function $\sigma$ is defined as $\mathcal{N}_{\theta}=L_{\theta_{L}}\circ\dots\circ L_{\theta_{1}}$. The input dimension of $\mathcal{N}_{\theta}$ is $n_{0}$ and the output dimension $n_{L}$. Moreover, we define the width of the network by $\mathcal{W}(\mathcal{N})=\max_{l}n_{l}$ and the depth by $\mathcal{D}(\mathcal{N})=L$.
Definition 17 (Model for $(\mathcal{F}_{n}^{m})_{n}$).
Let $\sigma:\mathbb{R}\to\mathbb{R}$ be locally Lipschitz continuous. Then we define for $L,(n_{l})_{l}$ depending on $m\in\mathbb{N}$ and $\Theta_{n}^{m}\subseteq\otimes_{l=1}^{L}\mathbb{R}^{n_{l}\times n_{l-1}}\times\mathbb{R}^{n_{l}}$ for $1\leq n\leq N$ with $n_{0}=1+N\sum_{k=0}^{\kappa}p_{k}$ and $n_{L}=1$ the class of parameterized approximation functions of the unknown terms,
$\mathcal{F}_{n}^{m}=\left\{\mathcal{N}_{\theta}~|~\theta\in\Theta_{n}^{m}\right\},$
for $n=1,\dots,N$ where each $\mathcal{N}_{\theta}:(0,T)\times(\otimes_{k=0}^{\kappa}\mathbb{R}^{p_{k}})^{N}\to\mathbb{R}$ is a fully connected feed forward neural network with activation function $\sigma$.
Remark 18.
Commonly used activation functions which are globally Lipschitz continuous include the softplus, saturated activation functions such as the sigmoid, hyperbolic tangent and Gaussian but also ReLU and some of its variations like the leaky ReLU and exponential linear unit amongst others. An example of a locally Lipschitz continuous activation function that is not globally Lipschitz continuous is the Rectified Quadratic Unit (ReQU).
Now as first step, we focus on the induction of well-defined Nemytskii operators and strong-weak continuity as specified in Assumption 3. Following [1, Lemma 4, Lemma 5], the former can be shown for general, continuous activation functions and the latter for locally Lipschitz continuous activation functions, both under the additional regularity assumption Assumption 5, i). Here, we focus on a different strategy that does not require Assumption 5, i), but assumes a globally Lipschitz continuous activation function. Note that in this section we write generically $\Theta$ instead of $\Theta_{n}^{m}$, as the results below on neural networks hold for general parameter sets as in Definition 16. The following result, whose proof can be found in Appendix A, shows that for $(\mathcal{F}_{n}^{m})_{n}$ as in Definition 17 and Lipschitz continuous $\sigma$ the properties in Assumption 3 follow.
Proposition 19.
Let Assumption 2 hold true. Suppose that $\sigma\in\mathcal{C}(\mathbb{R},\mathbb{R})$ is Lipschitz continuous with constant $L_{\sigma}$ (w.l.o.g. $L_{\sigma}\geq 1$). Then $\mathcal{N}_{\theta}:(0,T)\times(\otimes_{k=0}^{\kappa}\mathbb{R}^{p_{k}})^{N}\to\mathbb{R}$ induces a well-defined Nemytskii operator $\mathcal{N}_{\theta}:(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}\to L^{p}(0,T;L^{\hat{q}}(\Omega))$ via $\mathcal{N}_{\theta}(u)](t)=\mathcal{N}_{\theta}(u(t,\cdot))$. The same applies to $\mathcal{N}_{\theta}:(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}\to\mathcal{W}$. Furthermore,
$\mathcal{N}:\Theta\times(\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}}))^{N}\to L^{q}(0,T;L^{\hat{q}}(\Omega)),\quad(\theta,v)\mapsto\mathcal{N}_{\theta}(v)$
is strongly-strongly continuous.
Proof.
See Appendix A.1. ∎
Assuming a proper choice of the regularization functional $\mathcal{R}_{0}$, an important question is whether regularizing via
$\mathcal{R}(\varphi,u,\theta,u_{0},g)=\mathcal{R}_{0}(\varphi,u,u_{0},g)+\nu\|\theta\|+\|f_{\theta}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f_{\theta}\|_{L^{\infty}(U)}$ (14)
is justified for the class of parameterized approximation functions as introduced in Definition 17 based on feed forward neural networks. This includes both $W^{1,\infty}_{loc}-$regularity of the classes $\mathcal{F}_{n}^{m}$ and weak lower semicontinuity of (14) as required by Assumption 2. For the latter, in turn, it suffices to verify for fixed $n=1,\dots,N$ weak lower semicontinuity of the map
$\Theta\ni\theta\mapsto\|\mathcal{N}_{\theta}\|_{L^{\rho}(U)}+\|\nabla\mathcal{N}_{\theta}\|_{L^{\infty}(U)},$
again for a generic parameter set $\Theta$ in Definition 16. By weak lower semicontinuity of the $L^{\rho}$-norm and strong-strong continuity of $\Theta\ni\theta\mapsto\mathcal{N}_{\theta}\in L^{\infty}(U)$ (as follows from (41) in the proof of Proposition 19), for this, it remains to argue weak lower semicontinuity of
$\Theta\ni\theta\mapsto\|\nabla\mathcal{N}_{\theta}\|_{L^{\infty}(U)}.$ (15)
The next result, which is also proved in Appendix A, shows that ii) and vi) in Assumption 5 in fact hold true in this particular framework. In view of weak lower semicontinuity of (15) we restrict ourselves to the cases of Lipschitz continuous, $\mathcal{C}^{1}$-regular activation functions, and the Rectified Linear Unit (ReLU).
Proposition 20.
Assume that $\sigma\in\mathcal{C}(\mathbb{R},\mathbb{R})$ is locally Lipschitz continuous and let $(\mathcal{F}_{n}^{m})_{n}$ be given as in Definition 17. Then for $1\leq n\leq N$, $m\in\mathbb{N}$ it holds true that
$\mathcal{F}_{n}^{m}\subseteq W^{1,\infty}_{\text{loc}}(\mathbb{R}^{D}).$
Now fix some bounded subset $U\subseteq\mathbb{R}^{D}$. Let the activation function $\sigma$ either fulfill $\sigma\in\mathcal{C}^{1}(\mathbb{R},\mathbb{R})$ or let $\sigma$ be the Rectified Linear Unit. Then for $(\theta^{m})_{m}\subseteq\Theta$ with $\theta^{m}\to\theta\in\Theta$ as $m\to\infty$ it holds
$\|\nabla\mathcal{N}_{\theta}\|_{L^{\infty}(U)}\leq\liminf_{m\to\infty}\|\nabla\mathcal{N}_{\theta^{m}}\|_{L^{\infty}(U)}.$
Proof.
See Appendix A.2. ∎
We conclude the considerations on neural networks by discussing results from literature ensuring that Assumption 5, iv) holds true for suitably regular $f$. The corresponding estimate in (9) is closely related to universal approximation theory for neural networks, an active field of research which is presented e.g. in [29, 38, 47] and the references therein. Determining suitable functions $\psi$ regarding (9) for these approximation results is, however, not usually considered in works on neural network approximation theory and is in general not trivial. For an outline of state of the art results dealing with suitable estimates on $\psi$ we refer to the comparative overview presented in [53]. The result in [53] shows that a slight modification of the nearly optimal uniform approximation result of piecewise smooth functions by ReLU networks in [83] grows polynomially and in general yields a better bound than the other results providing polynomial bounds except for [11] which uses the ReQU activation function. As discussed in [53], the following (simplified) results hold true.
Proposition 21.
Let the parameterized classes in (8) be given by neural networks of the form in [53, Theorem 4] and $f\in\mathcal{C}^{q}(U)$ for some $q\geq 1$. Then (9) in Assumption 5, iv) holds true with $\beta=2q/D$ (with the networks attaining constant depth and width of order $m\log m$) and $\psi(m)=\tilde{c}m^{\frac{6q-3}{D}}$ for some constant $\tilde{c}>0$.
Proposition 22.
Let the parameterized classes in (8) be given by neural networks of the form in [11, Theorem 1] and $f\in\mathcal{C}^{q}(U)$ for some $q\geq 1$. Then (9) in Assumption 5, iv) holds true with $\beta=q/D$ (with the networks attaining constant depth and width of order $m$) and $\psi(m)=\tilde{c}$ for some constant $\tilde{c}>0$.
Recall that a similar result as in Proposition 19 and Proposition 20 holds true for locally Lipschitz continuous activations (e.g. the ReQU activating the networks in [11, Theorem 1]) under the additional regularity Assumption 5, i) as discussed above directly after Remark 18. It remains to discuss the convergence of $\|\nabla f_{\theta^{m}}\|_{L^{\infty}(U)}\to\|\nabla f\|_{L^{\infty}(U)}$ as $m\to\infty$. The result in [11, Theorem 1] realizes also the simultaneous approximation of higher order derivatives at the loss of a poorer approximation rate. The work by [112] considers approximation in $W^{m,p}$-Sobolev norms for integer $m\geq 2$. Note that both results are stronger than the previously stated convergence. The works [48, 49] cover $W^{1,\infty}$-approximation by ReLU neural networks, thus, in particular inferring this type of convergence. However, a parameter estimation as stated in Assumption 5, iv) is not covered. Alternatively, e.g. for the result in [53, Theorem 4], one might eventually apply a lifting technique as outlined in Appendix A.3. This is possible in case $f$ attains higher regularity such as $W^{2,\infty}$- or $\mathcal{C}^{2}$-regularity
2.3 Physical term
In the next subsections we verify Assumption 4 in the setup of affine linear physical terms and in the general setup of nonlinear physical terms, and provide examples.
2.3.1 Linear case
We assume that the physical term is given in linear form for some fixed $\omega\in\mathbb{N}_{0}$ by
$F(t,(u_{n})_{1\leq n\leq N},\varphi)$ $=\Psi(t,\varphi)+\sum_{n=1}^{N}\mathcal{J}_{\omega}u_{n}\cdot\Phi_{n}(t,\varphi)$ (16)
$\text{with}~~~\mathcal{J}_{\omega}u_{n}\cdot\Phi_{n}(t,\varphi):=\sum_{0\leq|\beta|\leq\omega}D^{\beta}u_{n}\cdot\Phi_{n,\beta}(t,\varphi)$
for $t\in(0,T),(u_{n})_{1\leq n\leq N}\in V^{N},\varphi\in X_{\varphi}$, where we suppose that $(V\hookrightarrow)\tilde{V}\hookrightarrow W^{\omega,\hat{p}}(\Omega)$. The functionals $\Psi$ and $(\Phi_{n,\beta})_{n,\beta}$ are given as $\Psi:(0,T)\times X_{\varphi}\to L^{\hat{q}}(\Omega)$ and $\Phi_{n,\beta}:(0,T)\times X_{\varphi}\to L^{s_{\beta}}(\Omega)$ for $1\leq n\leq N$, $0\leq|\beta|\leq\omega$ and some suitable $1\leq s_{\beta}\leq\infty$ (to be determined below). Since $\Psi(t,\varphi)\in W$ due to $L^{\hat{q}}(\Omega)\hookrightarrow W$, in order to show that $F(t,(u_{n})_{1\leq n\leq N},\varphi)\in W$ (i.e., that $F$ is well-defined) it suffices to choose the $s_{\beta}$ such that $\mathcal{J}_{\omega}u_{n}\cdot\Phi_{n}(t,\varphi)\in W$. This can be done as follows. For $(u_{n})_{1\leq n\leq N}\in V^{N}$ we have that $D^{\beta}u_{n}\in W^{\omega-|\beta|,\hat{p}}(\Omega)$ for $0\leq|\beta|\leq\omega$ and $D^{\beta}u_{n}\cdot\Phi_{n,\beta}(t,\varphi)\in L^{\hat{q}}(\Omega)\hookrightarrow W$, which shows welldefinedness of (16), if one of the following conditions on $s_{\beta}$ is fulfilled:
[…]
In the first two cases welldefinedness of (16) is a consequence of [9, Theorem 6.1] (see also [9, Remark 6.2, Corollary 6.3] for the generalization to bounded Lipschitz domains). In the last case (with $\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}}$$\infty$ for $\hat{p}=\hat{q}$ which is important as $s_{\beta}=\infty$ is excluded in the first two cases) welldefinedness follows by $V\hookrightarrow W^{\omega,\hat{p}}(\Omega)$ and Hölder’s inequality. To achieve Assumption 4, ii) we need stricter conditions than (17) in general as outlined below. The following result, proven in Appendix B, covers Assumption 4 in the linear setup.
Proposition 23.
Let Assumption 2 hold true and $\tilde{V}\hookrightarrow W^{\omega,\hat{p}}(\Omega)$. Suppose that $t\mapsto\Phi_{n}(t,\varphi)$ and $t\mapsto\Psi(t,\varphi)$ are measurable for all $\varphi\in X_{\varphi}$ and $s_{\beta}$ fulfill (17). Assume that there exist functions $\mathcal{B}_{1},\mathcal{B}_{2}:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ that map bounded sets to bounded sets and $\phi\in L^{\frac{pq}{p-q}}(0,T)$ (with $\phi\in L^{\infty}(0,T)$ if $p=q$), $\psi\in L^{q}(0,T)$ such that
$\|\Phi_{n,\beta}(t,\varphi)\|_{L^{s_{\beta}}(\Omega)}\leq\phi(t)\mathcal{B}_{1}(\|\varphi\|_{X_{\varphi}}),~~~\|\Psi(t,\varphi)\|_{L^{\hat{q}}(\Omega)}\leq\psi(t)\mathcal{B}_{2}(\|\varphi\|_{X_{\varphi}}).$ (18)
Then $F$ in (16) induces a well-defined Nemytskii operator $F:\mathcal{V}^{N}\times X_{\varphi}\to\mathcal{W}$ with
$F((u_{n})_{1\leq n\leq N},\varphi)](t)=F(t,(u_{n}(t))_{1\leq n\leq N},\varphi)$
for $(u_{n})_{1\leq n\leq N}\in\mathcal{V}^{N},\varphi\in X_{\varphi}$ and $t\in(0,T)$. Now suppose that $\Psi(t,\cdot):X_{\varphi}\to L^{\hat{q}}(\Omega)$ and $\Phi_{n,\beta}(t,\cdot):X_{\varphi}\to L^{s_{\beta}}(\Omega)$ are weakly continuous for almost every $t\in(0,T)$, additionally with $\frac{\omega-|\beta|}{d}>\frac{1}{\hat{p}}-\frac{1}{\hat{q}}+\frac{1}{s_{\beta}}$ if $\hat{q}=1$ or $s_{\beta}=\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}}$. Furthermore, suppose that either $\omega\leq\kappa$ or otherwise in case $\omega>\kappa$ the following additional conditions hold:
[…]
    $\|\Phi_{n,\beta}(t,\varphi)\|_{\frac{c_{\beta}\hat{q}}{c_{\beta}-\hat{q}}}\leq\phi(t)\mathcal{B}_{1}(\|\varphi\|_{X_{\varphi}}).$
  • For $|\beta|=\omega$ assume that $\Phi_{n,\beta}(t,\cdot):X_{\varphi}\to L^{\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}}}(\Omega)$ is well-defined and weak-strong continuous for a.e. $t\in(0,T)$.
Then $\mathcal{V}^{N}\times X_{\varphi}\ni(u,\varphi)\mapsto F(u,\varphi)\in\mathcal{W}$ induced by (16) is weak-weak continuous.
Proof.
See Appendix B.1. ∎
To conclude this subsection we give the following example which is motivated by the parabolic problem considered in [1, Chapter 4]. We restrict ourselves to a single equation which can be immediately generalized to general systems by introducing technical notation. Note that the space setup in the following example is consistent with Assumption 2, but we do not discuss it in order to not distract from the central conditions on the parameters. In the example we suppose that the physical term is governed by a reaction-diffusion equation such that $\omega=2$ in (16). Here, we assume that the general task consists in reconstructing additional unknown convection terms (which are of derivative order one) such that we can choose $\kappa=1$ in Assumption 2. If only additional unknown reaction terms need to be reconstructed, one could choose $\kappa=0$.
Example 24.
Let $V=\tilde{V}=W^{2,\hat{p}}(\Omega)$, $W=L^{\hat{p}}(\Omega)$, $\mathcal{V},\mathcal{W}$ as in Assumption 2, $\kappa=1$, $\hat{p}=\hat{q}=2$ and
$F(t,u,\varphi)=\nabla\cdot(a\nabla u)+cu$
for $t\in(0,T),u\in\mathcal{V}$ and $\varphi=(a,c)$ with $a\in W^{1,\gamma}(\Omega)$ for $3=d<\gamma<\infty$ and $c\in L^{2}(\Omega)$. Note that $X_{\varphi}=W^{1,\gamma}(\Omega)\times L^{2}(\Omega)$. Thus, the physical term $F$ attains a representation of the form in (16) with $\omega=2$, $\Psi\equiv 0$ and under abuse of notation
$\Phi_{\bar{0}}(t,\varphi)=c,~\Phi_{e_{k}}(t,\varphi)=\partial_{x_{k}}a,~\Phi_{2e_{k}}(t,\varphi)=a$
for $1\leq k\leq 3$ with $e_{k}$ the $k$-th unit vector in $\mathbb{R}^{3}$ and $\bar{0}=(0,0,0)$. Furthermore, we set $\Phi_{\beta}\equiv 0$ for $\beta\notin\left\{0e_{k},e_{k},2e_{k}\right\}_{1\leq k\leq 3}$. We verify the requirements on $\Phi$ in Proposition 23 based on the following case distinction for $0\leq|\beta|\leq 2$.
Case 1. $|\beta|=0$: For $s_{\bar{0}}=2$ (fulfills $\frac{\omega-|\beta|}{d}>\frac{1}{\hat{p}}-\frac{1}{\hat{q}}+\frac{1}{s_{\bar{0}}}$ and $\hat{q}\leq s_{\bar{0}}<\infty$) it holds
$\|\Phi_{\bar{0}}(t,\varphi)\|_{L^{s_{\bar{0}}}(\Omega)}=\|c\|_{L^{2}(\Omega)}\leq\|\varphi\|_{X_{\varphi}},$
[…]
Case 2. $|\beta|=1$: For $\hat{q}\leq d<s_{e_{k}}\leq\gamma<\infty$ (fulfills $\frac{\omega-|\beta|}{d}>\frac{1}{\hat{p}}-\frac{1}{\hat{q}}+\frac{1}{s_{\beta}}$) it holds
$\|\Phi_{e_{k}}(t,\varphi)\|_{L^{s_{e_{k}}}(\Omega)}=\|\partial_{x_{k}}a\|_{L^{s_{e_{k}}}(\Omega)}\leq|\Omega|^{\frac{\gamma-s_{e_{k}}}{\gamma s_{e_{k}}}}\|\partial_{x_{k}}a\|_{L^{\gamma}(\Omega)}\leq|\Omega|^{\frac{\gamma-s_{e_{k}}}{\gamma s_{e_{k}}}}\|\varphi\|_{X_{\varphi}},$
[…]
Case 3. $|\beta|=2$: We may choose $s_{2e_{k}}=\infty$. As $\gamma>d$ there exists some constant $c_{\gamma}>0$ such that $\|\cdot\|_{\mathcal{C}(\overline{\Omega})}\leq c_{\gamma}\|\cdot\|_{W^{1,\gamma}(\Omega)}$ yielding
$\|\Phi_{2e_{k}}(t,\varphi)\|_{L^{s_{2e_{k}}}(\Omega)}=\|a\|_{L^{\infty}(\Omega)}\leq c_{\gamma}\|a\|_{W^{1,\gamma}(\Omega)}\leq c_{\gamma}\|\varphi\|_{X_{\varphi}}$
[…]
Thus, the requirements on $\Phi$ and $\Psi$ in Proposition 23 are fulfilled.
2.3.2 Nonlinear case
The following result, proven in Appendix B.2, verifies Assumption 4 for general nonlinear physical terms under stronger conditions. Note that instead of weak closedness in Assumption 4, ii) we show weak-weak continuity which is stronger.
Proposition 25.
Let Assumption 2 and the extended state space embedding
$\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$
hold true. Suppose that the $F_{n}(\cdot,\cdot,\varphi):(0,T)\times V^{N}\to W$ satisfy the Carathéodory condition, i.e., $t\mapsto F_{n}(t,v,\varphi)$ is measurable for $v\in V^{N}$ and $v\mapsto F_{n}(t,v,\varphi)$ is continuous for a.e. $t\in(0,T)$. Further assume that the $F_{n}$ satisfy the growth condition
$\|F_{n}(t,(v_{n})_{1\leq n\leq N},\varphi)\|_{W}\leq\mathcal{B}_{0}(\|\varphi\|_{X_{\varphi}},\sum_{n=1}^{N}\|v_{n}\|_{H})(\Gamma(t)+\sum_{n=1}^{N}\|v_{n}\|_{V})$ (19)
for some $\Gamma\in L^{q}(0,T)$ and $\mathcal{B}_{0}:\mathbb{R}^{2}\to\mathbb{R}$, increasing in the second entry and, for fixed second entry, mapping bounded sets to bounded sets. Then the $F_{n}:(0,T)\times V^{N}\times X_{\varphi}\to W$ induce well-defined Nemytskii operators $F_{n}:\mathcal{V}^{N}\times X_{\varphi}\to\mathcal{W}$ with
$[F_{n}(v,\varphi)](t)=F_{n}(t,v(t),\varphi)$ (20)
for $v\in\mathcal{V}^{N}$ and $\varphi\in X_{\varphi}$. Now suppose weak-weak continuity of
$F_{n}(t,\cdot):H^{N}\times X_{\varphi}$ $\to W$
$(v_{1},\dots,v_{N},\varphi)$ $\mapsto F_{n}(t,v_{1},\dots,v_{N},\varphi)$
for a.e. $t\in(0,T)$. Further assume that the $F_{n}$ satisfy the stricter growth condition
$\|F_{n}(t,(v_{n})_{1\leq n\leq N},\varphi)\|_{W}\leq\mathcal{B}_{0}(\|\varphi\|_{X_{\varphi}},\sum_{n=1}^{N}\|v_{n}\|_{H})(\Gamma(t)+\sum_{n=1}^{N}\|v_{n}\|_{H})$ (21)
for some $\Gamma\in L^{q}(0,T)$ and $\mathcal{B}_{0}:\mathbb{R}^{2}\to\mathbb{R}$ as above. Then (20) is weak-weak continuous.
Proof.
See Appendix B.2. ∎
Remark 26.
A possible application case of the previous proposition is the following. Assume that there exists a reflexive, separable Banach space $V^{\prime}$ and $\lambda\in\mathbb{N}_{0}$ with
$H\hookrightarrow W^{\lambda,\hat{p}}(\Omega)\hookrightarrow V^{\prime}$ (22)
with the property that $F:(0,T)\times(V^{\prime})^{N}\times X_{\varphi}\to W$ is well-defined. One might think of physical terms which regarding the state space variable do not need all higher order derivative information provided by the space $V$ (eventually given by $V=W^{\kappa+m,p_{0}}(\Omega)$ as outlined in Remark 7) but only $\lambda<\kappa+m$ many. Then the growth condition in (19) with $\|\cdot\|_{V^{\prime}}$ instead of $\|\cdot\|_{V}$ implies condition (21) due to (22). Note that $H$ needs to be regular enough to be embeddable in $W^{\lambda,\hat{p}}(\Omega)$.
The condition in (22) can be also understood the other way around. That is for given $H$ one might determine the maximal $\lambda\in\mathbb{N}$ such that $H\hookrightarrow W^{\lambda,\hat{p}}(\Omega)$. Then the previous considerations cover physical terms which are well-defined regarding state space variables with highest derivative order given by $\lambda$.
To conclude this subsection we give the following example addressing the ideas in Remark 26 more concretely. We restrict ourselves to a single equation which can be immediately generalized to general systems by introducing technical notation. Note that the space setup in the following example is consistent with Assumption 2, but we do not discuss it in order to not distract from the central conditions on the parameters. For some preliminary ideas regarding the embedding $\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$ see Remark 9 where one might have $\tilde{V}=H$.
Example 27.
We consider a simple three-dimensional transport problem where it is assumed that the known physics are governed by the inviscid Burgers’ equation, i.e., we have $F(u)=-u\partial_{x}u-u\partial_{y}u-u\partial_{z}u$. Anticipating eventual viscosity effects we suppose that the unknown approximated term accounts for these effects. Let $V=W^{2,\hat{p}}(\Omega),\tilde{V}=H=W^{1,2}(\Omega)$, $W=L^{\hat{p}}(\Omega)$, $d=3$, $\kappa=1$ and $\hat{p}=\hat{q}=\frac{6-\epsilon}{4-\epsilon/2}$ for some small $0<\epsilon<1$. Then we have for $u\in V$ as $L^{3/2}(\Omega)\hookrightarrow W$ for some $c>0$ that
$\|F(u)\|_{W}\leq c\|u(\partial_{x}u+\partial_{y}u+\partial_{z}u)\|_{L^{3/2}(\Omega)}\leq c\|u\|_{L^{6}(\Omega)}\|\nabla u\|_{L^{2}(\Omega)}$
where the last inequality follows by the generalized Hölder’s inequality. Due to the embedding $W^{1,2}(\Omega)\hookrightarrow L^{6}(\Omega)$ (recall that $d=3$) we derive that $\|F(u)\|_{W}\leq c\|u\|_{H}^{2}$ and hence, a growth condition of the form in (21).
To see weak-weak continuity of $F:H\to W$ let $(u_{n})_{n}\subseteq H$ with $u_{n}\rightharpoonup u\in H$ as $n\to\infty$. Then for $w\in L^{\hat{p}^{*}}(\Omega)$ we have that
$\langle u_{n}(\partial_{x}u_{n}+\partial_{y}u_{n}+\partial_{z}u_{n})-u(\partial_{x}u+\partial_{y}u+\partial_{z}u),w\rangle_{L^{\hat{p}}(\Omega),L^{\hat{p}^{*}}(\Omega)}$
can be rewritten for $e=\begin{pmatrix}1&1&1\end{pmatrix}^{T}\in\mathbb{R}^{3}$ by
$\langle u~e\cdot(\nabla u_{n}-\nabla u),w\rangle_{L^{\hat{p}}(\Omega),L^{\hat{p}^{*}}(\Omega)}+\langle(u_{n}-u)~e\cdot\nabla u_{n},w\rangle_{L^{\hat{p}}(\Omega),L^{\hat{p}^{*}}(\Omega)}.$ (23)
For the first term in (23) note that $\nabla u_{n}\rightharpoonup\nabla u$ in $L^{2}(\Omega)$ as $n\to\infty$. As
$\langle u~e\cdot(\nabla u_{n}-\nabla u),w\rangle_{L^{\hat{p}}(\Omega),L^{\hat{p}^{*}}(\Omega)}=\int_{\Omega}u(x)~e\cdot(\nabla u_{n}(x)-\nabla u(x))w(x)\mathop{}\!\mathrm{d}x$
it suffices to show that $uw\in L^{2}(\Omega)$ to obtain the convergence $u~e\cdot(\nabla u_{n}-\nabla u)\rightharpoonup 0$ in $L^{\hat{p}}(\Omega)$ as $n\to\infty$. This follows by $u\in W^{1,2}(\Omega)\hookrightarrow L^{6}(\Omega)$, Hölder’s generalized inequality and $\hat{p}^{*}=\frac{6-\epsilon}{2-\epsilon/2}$ as
$\left(\frac{1}{6}+\frac{2-\epsilon/2}{6-\epsilon}\right)^{-1}=\frac{6-\epsilon}{3-2\epsilon/3}\geq 2.$
It remains to show that the second term in (23) approaches zero as $n\to\infty$. By
$\langle(u_{n}-u)~e\cdot\nabla u_{n},w\rangle_{L^{\hat{p}}(\Omega),L^{\hat{p}^{*}}(\Omega)}=\int_{\Omega}(u_{n}(x)-u(x))~e\cdot\nabla u_{n}(x)w(x)\mathop{}\!\mathrm{d}x$
it suffices to show that $(\nabla u_{n}w)_{n}$ is uniformly bounded in $L^{\frac{6-\epsilon}{5-\epsilon}}(\Omega)$ as $u_{n}\to u$ in $L^{6-\epsilon}(\Omega)$ by the Rellich-Kondrachov Theorem. This follows by boundedness of $(\nabla u_{n})_{n}$ in $L^{2}(\Omega)$ due to weak convergence and Hölder’s generalized inequality concluding weak-weak continuity of $F$.
Remark 28.
The choice of $V=W^{2,\hat{p}}(\Omega)$ in the previous example might seem unnecessarily strong for modeling the inviscid Burgers’ equation. The reason for this choice is that we suppose that the additional hidden physics that need to be reconstructed possibly include viscosity effects which are of order $\kappa=1$, i.e., the (un)known physical effects are of equal differential order. As a consequence, the state space $V$ needs to attain higher regularity to cover Assumption 2 (see also Remark 7). If one would anticipate only additional unknown reaction terms ($\kappa=0$) choosing $V$ of lower regularity would be possible. In any case, however, it is important to note that the stronger regularity requirement comes from the fact that we simultaneously recover both the state and a non-linear term acting on the state. During the learning process, this regularity can be enforced by using appropriate regularization. Afterwards, in applications of the learned model, the additional regularity is no longer required.
3 The uniqueness problem
The starting point of our considerations on uniqueness is Assumption 5, iii), where we assume for given full measurement data $(\hat{y}^{l})_{l}\in\mathcal{Y}^{L}$ and $F:\mathcal{V}^{N}\times X_{\varphi}\to\mathcal{W}^{N}$, to be understood as in Section 2, the existence of $\hat{f}:(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}\to\mathcal{W}^{N}$, a state $(\hat{u}_{n}^{l})_{\begin{subarray}{c}1\leq n\leq N\\ 1\leq l\leq L\end{subarray}}\in\mathcal{V}^{N\times L}$, an initial condition $(\hat{u}_{0,n}^{l})_{\begin{subarray}{c}1\leq n\leq N\\ 1\leq l\leq L\end{subarray}}\in H^{N\times L}$, a boundary condition $(\hat{g}_{n}^{l})_{\begin{subarray}{c}1\leq n\leq N\\ 1\leq l\leq L\end{subarray}}\in\mathcal{B}^{N\times L}$ and a source term $(\hat{\varphi}_{n}^{l})_{\begin{subarray}{c}1\leq n\leq N\\ 1\leq l\leq L\end{subarray}}\in X_{\varphi}^{N\times L}$ solving the system of partial differential equations ($S$), i.e.,
$\partial_{t}\hat{u}^{l}=F(t,\hat{u}^{l},\hat{\varphi}^{l})+\hat{f}(t,\mathcal{J}_{\kappa}\hat{u}^{l})$ ($S$)
$\text{s.t.}~~\hat{u}^{l}(0)=\hat{u}_{0}^{l},~\gamma(\hat{u}^{l})=\hat{g}^{l},$~~
together with the measurements
$K^{\dagger}\hat{u}^{l}=\hat{y}^{l}$ ($M$)
for $l=1,\ldots,L$. The results of this section are developed based on Assumption 2 to 5. Note that under these assumptions, due to ($M$) and injectivity of the full measurement operator $K^{\dagger}$ by Assumption 5, v), the state $\hat{u}$ is uniquely given in system ($S$) even if the term $\hat{f}$ is not.
We recall that the bounded Lipschitz domain $U$ is chosen and fixed according to Assumption 5, vii). Note that by Assumption 5, ii), it holds that $\mathcal{F}_{n}^{m}\subseteq W^{1,\infty}(U)$ for $1\leq n\leq N$, $m\in\mathbb{N}$.
Before we move on to the limit problem and question of uniqueness let us justify the choice of regularization for $f_{\theta}\in W^{1,\infty}(U)^{N}$. The problem of using the $W^{1,\infty}(U)$-norm directly is that its powers are not strictly convex which is necessary for uniqueness issues later. This is overcome by the well known equivalence of the norms $\|\cdot\|_{W^{1,\infty}(U)}$ and $\|\cdot\|_{L^{\rho}(U)}+|\cdot|_{W^{1,\infty}(U)}$ on $W^{1,\infty}(U)$ for bounded domains $U$, which follows by [15, 6.12 A lemma of J.-L. Lions] and [15, Theorem 9.16 (Rellich–Kondrachov)]. That is, the space $W^{1,\infty}(U)$ may be strictly convexified under the equivalent norm $\|\cdot\|_{L^{\rho}(U)}+|\cdot|_{W^{1,\infty}(U)}$ for $1<\rho<\infty$ with $|\cdot|_{W^{1,\infty}(U)}$ the seminorm in $W^{1,\infty}(U)$.
The following proposition introduces the limit problem and shows uniqueness:
Proposition 29.
Let Assumptions 2 to 5 without Assumption 5, iv) be satisfied. Then there exists a unique solution $(\varphi^{\dagger},u^{\dagger},u_{0}^{\dagger},g^{\dagger},f^{\dagger})\in X_{\varphi}^{N\times L}\times\mathcal{V}^{N\times L}\times H^{N\times L}\times\mathcal{B}^{N\times L}\times W^{1,\infty}(U)^{N}$ to
$\min_{\begin{subarray}{c}\varphi\in X_{\varphi}^{N\times L},u\in\mathcal{V}^{N\times L},\\ u_{0}\in H^{N\times L},g\in\mathcal{B}^{N\times L},\\ f\in W^{1,\infty}(U)^{N}\end{subarray}}$ $\mathcal{R}_{0}(\varphi,u,u_{0},g)+\|f\|_{L^{\rho}(U)}^{\rho}+\|\nabla f\|_{L^{\infty}(U)}$ ($\mathcal{P}^{\dagger}$)
s.t. $\frac{\partial}{\partial t}u_{n}^{l}-F_{n}(t,u^{l}_{1},\dots,u^{l}_{N},\varphi^{l}_{n})-f_{n}(t,\mathcal{J}_{\kappa}u^{l}_{1},\dots,\mathcal{J}_{\kappa}u^{l}_{N})=0,$
$K^{\dagger}u^{l}=y^{l},\,u_{n}^{l}(0)=u_{0,n}^{l},\,\gamma(u^{l})=g^{l}.$
Proof.
First of all, the constraint set of problem ($\mathcal{P}^{\dagger}$) is not empty by Assumption 5, iii), i.e., there exist admissible functions solving system ($S$) such that ($M$) holds true. Due to injectivity of the full measurement operator $K^{\dagger}$, for any element satisfying the constraint set of ($\mathcal{P}^{\dagger}$) the state is uniquely given by $u^{\dagger}=\hat{u}$. As a consequence, also the initial and boundary trace are uniquely determined by $u_{0}^{\dagger}=u^{\dagger}(0)=\hat{u}(0)=\hat{u}_{0}$ and $(g^{\dagger,l}_{n})_{n,l}=(\gamma(u^{\dagger,l}_{n}))_{n,l}=(\gamma(\hat{u}^{l}_{n}))_{n,l}=(\hat{g}^{l}_{n})_{n,l}$, respectively. By Assumption 5, i), iii) and vi) it follows that
$)\displaystyle\|\mathcal{J}_{\kappa}u^{\dagger}\|_{L^{\infty}((0,T)\times\Omega)}=\|\mathcal{J}_{\kappa}\hat{u}\|_{L^{\infty}((0,T)\times\Omega)}\leq c_{\mathcal{V}}\|\hat{u}\|_{\mathcal{V}}\leq c_{\mathcal{V}}\pi(\mathcal{R}_{0}(\hat{\varphi},\hat{u},\hat{u}_{0},\hat{g}$)) (24)
and hence, that $(t,\mathcal{J}_{\kappa}u^{\dagger,l}(t,x))\in U$ for $(t,x)\in(0,T)\times\Omega$ by Assumption 5, vii).
Thus, problem ($\mathcal{P}^{\dagger}$) may be rewritten equivalently by
$\min_{\begin{subarray}{c}\varphi\in X_{\varphi}^{N\times L},\\ f\in W^{1,\infty}(U)^{N}\end{subarray}}$ $\mathcal{R}_{0}(\varphi,u^{\dagger},u^{\dagger}_{0},g^{\dagger})+\|f\|_{L^{\rho}(U)}^{\rho}+\|\nabla f\|_{L^{\infty}(U)}$ (25)
s.t. $\frac{\partial}{\partial t}u^{\dagger,l}-F(t,u^{\dagger,l},\varphi^{l})-f(t,\mathcal{J}_{\kappa}u^{\dagger,l})=0.$
The existence of a solution to (25) follows by the direct method: In the following, w.l.o.g., we omit a relabelling of sequences to convergent subsequences. Using the norm equivalence of $\|\cdot\|_{W^{1,\infty}(U)}$, $\|\cdot\|_{L^{\rho}(U)}+|\cdot|_{W^{1,\infty}(U)}$ and coercivity of $\mathcal{R}_{0}$ a minimizing sequence $(\varphi^{k},f^{k})_{k}\subseteq X_{\varphi}^{N\times L}\times W^{1,\infty}(U)^{N}$ to (25) is bounded. Thus, there exist $\varphi^{\prime}\in X_{\varphi}^{N\times L}$ and $f^{\prime}\in W^{1,\infty}(U)^{N}$ such that $\varphi^{k}\rightharpoonup\varphi^{\prime}$ in $X_{\varphi}^{N\times L}$ and $f^{k}\overset{*}{\rightharpoonup}f^{\prime}$ in $W^{1,\infty}(U)^{N}$ as $k\to\infty$ by reflexivity of $X_{\varphi}^{N\times L}$ and $W^{1,\infty}(U)^{N}$ being the dual of a separable space. By $f^{k}\overset{*}{\rightharpoonup}f^{\prime}$ in $L^{\infty}(U)^{N}$ and $\nabla f^{k}\overset{*}{\rightharpoonup}\nabla f^{\prime}$ in $L^{\infty}(U)^{N}$ as $k\to\infty$ together with $L^{\infty}(U)\hookrightarrow L^{\rho}(U)$, $1<\rho<\infty$ and weak lower semicontinuity of $\mathcal{R}_{0}$ it follows that $(\varphi^{\prime},f^{\prime})\in X_{\varphi}^{N\times L}\times W^{1,\infty}(U)^{N}$ minimizes the objective functional of (25). We argue that also
$\partial_{t}u^{\dagger,l}=F(t,u^{\dagger,l},\varphi^{\prime l})+f^{\prime}(t,\mathcal{J}_{\kappa}u^{\dagger,l})$ (26)
concluding that $(\varphi^{\prime},f^{\prime})$ is indeed a solution of (25). For that note that $f^{k}\to f^{\prime}$ in $\mathcal{C}(\overline{U})^{N}$ as $k\to\infty$ by the Rellich-Kondrachov Theorem. Thus, by $L^{q}(0,T;L^{\hat{p}}(\Omega))\hookrightarrow\mathcal{W}$ and boundedness of $U$ together with (24) and $u^{\dagger}=\hat{u}$ we have for some $c>0$
$\|f^{k}(\mathcal{J}_{\kappa}u^{\dagger,l})-f^{\prime}(\mathcal{J}_{\kappa}u^{\dagger,l})\|_{\mathcal{W}}\leq c\|f^{k}-f^{\prime}\|_{L^{\infty}(U)},$
and conclude that $f^{k}(\mathcal{J}_{\kappa}u^{\dagger,l})\to f^{\prime}(\mathcal{J}_{\kappa}u^{\dagger,l})$ in $\mathcal{W}^{N}$ as $k\to\infty$. Using this, as a consequence of boundedness of $\|\partial_{t}u^{\dagger,l}-f^{k}(\mathcal{J}_{\kappa}u^{\dagger,l})\|_{\mathcal{W}}$ for $k\in\mathbb{N}$ it follows by Assumption 4, ii) that $F(u^{\dagger,l},\varphi^{k,l})\rightharpoonup F(u^{\dagger,l},\varphi^{\prime l})$ in $\mathcal{W}^{N}$ as $k\to\infty$ and we recover (26).
Finally, uniqueness of $(\varphi^{\dagger},f^{\dagger})=(\varphi^{\prime},f^{\prime})$ as solution to (25) follows from strict convexity of the objective functional in $(\varphi,f)\in X_{\varphi}^{N\times L}\times W^{1,\infty}(U)^{N}$ and from $F$ being affine with respect to $\varphi$. ∎
Now recall that, under Assumption 5, the minimization problem (2) reduces to the following specific case:
$\min_{\begin{subarray}{c}\varphi\in X_{\varphi}^{N\times L},\theta\in\otimes_{n}\Theta^{m}_{n},\\ u\in\mathcal{V}^{N\times L},u_{0}\in H^{N\times L},\\ g\in\mathcal{B}^{N\times L}\end{subarray}}\sum_{1\leq l\leq L}\bigg[\lambda^{m}\bigg(\|\frac{\partial}{\partial t}u^{l}-F(t,u^{l},\varphi^{l})-f_{\theta}(t,\mathcal{J}_{\kappa}u^{l})\|_{\mathcal{W}}^{q}$ ($\mathcal{P}^{m}$)
$+\|u^{l}(0)-u_{0}^{l}\|_{H}^{2}+\mathcal{D}_{\text{BC}}(\gamma(u^{l})-g^{l})\bigg)+\mu^{m}\|K^{m}u^{l}-y^{m,l}\|_{\mathcal{Y}}^{r}\bigg]$
$+\mathcal{R}_{0}(\varphi,u,u_{0},g)+\nu^{m}\|\theta\|+\|f_{\theta}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f_{\theta}\|_{L^{\infty}(U)}$
for a sequence of measured data $\mathcal{Y}\ni y^{m,l}\approx K^{m}u^{\dagger,l}$ for $m\in\mathbb{N}$ and $1\leq l\leq L$ with $u^{\dagger}$ as in Proposition 29. More concretely the measured data $(y^{m,l})_{m}\subseteq\mathcal{Y}$ is supposed to be given under some noise estimation
$\|y^{m,l}-K^{m}u^{\dagger,l}\|_{\mathcal{Y}}\leq\delta(m)$ (27)
for $1\leq l\leq L$ with $\delta:\mathbb{N}\to\mathbb{R}_{\geq 0}$ the noise estimator such that $\lim_{m\to\infty}\delta(m)=0$.
Our main result on approximating the unique solution of ($\mathcal{P}^{\dagger}$) is now the following:
Theorem 30.
Let Assumptions 2 to 5 hold true with the approximation capacity condition in Assumption 5, iv) being satisfied for $f^{\dagger}$ where $(\varphi^{\dagger},u^{\dagger},u_{0}^{\dagger},g^{\dagger},f^{\dagger})$ is the unique solution to ($\mathcal{P}^{\dagger}$).
  • Let $(\varphi^{m},\theta^{m},u^{m},u_{0}^{m},g^{m})$ be a solution to ($\mathcal{P}^{m}$) for each $m\in\mathbb{N}$.
  • Let further the parameters $\lambda^{m},\mu^{m},\nu^{m}>0$ be chosen such that $\lambda^{m}\to\infty$, $\mu^{m}\to\infty$ and $\nu^{m}\to 0$ with $\lambda^{m}m^{-\beta q}=o(1)$, $\mu^{m}\delta(m)^{r}=o(1)$ and $\nu^{m}\psi(m)=o(1)$ as $m\to\infty$.
Then $\varphi^{m}\rightharpoonup\varphi^{\dagger}$ in $X_{\varphi}^{N\times L}$, $u^{m}\rightharpoonup u^{\dagger}$ in $\mathcal{V}^{N\times L}$, $u_{0}^{m}\rightharpoonup u_{0}^{\dagger}$ in $H^{N\times L}$, $g^{m}\rightharpoonup g^{\dagger}$ in $\mathcal{B}^{N\times L}$ and $f_{\theta^{m}}\overset{*}{\rightharpoonup}f^{\dagger}$ in $W^{1,\infty}(U)^{N}$.
Proof.
First of all, the existence of solutions $(\varphi^{m},\theta^{m},u^{m},u_{0}^{m},g^{m})$ to ($\mathcal{P}^{m}$) for each $m\in\mathbb{N}$ follows by the direct method which is discussed in all details in Appendix C on the existence of minimizers. Let now $c>0$ be a generic constant used throughout the following estimations. By Assumption 5, iv) being satisfied for $f^{\dagger}$, there exist $\tilde{\theta}^{m}\in\otimes_{n=1}^{N}\Theta_{n}^{m}$ such that $\|f^{\dagger}-f_{\tilde{\theta}^{m}}\|_{L^{\infty}(U)}\leq cm^{-\beta}$ and $\|\tilde{\theta}^{m}\|\leq\psi(m)$ for $m\in\mathbb{N}$ together with $\|\nabla f_{\tilde{\theta}^{m}}\|_{L^{\infty}(U)}\to\|\nabla f^{\dagger}\|_{L^{\infty}(U)}$ as $m\to\infty$. As $(\varphi^{m},\theta^{m},u^{m},u_{0}^{m},g^{m})$ is a solution to Problem ($\mathcal{P}^{m}$) we may estimate its objective functional value using the noise estimate (27) by
$\sum_{1\leq l\leq L}\bigg[\lambda^{m}\bigg(\|\frac{\partial}{\partial t}u^{m,l}-F(t,u^{m,l},\varphi^{m,l})-f_{\theta^{m}}(t,\mathcal{J}_{\kappa}u^{m,l})\|_{\mathcal{W}}^{q}$
$~~+\|u^{m,l}(0)-u_{0}^{m,l}\|_{H}^{2}+\mathcal{D}_{\text{BC}}(\gamma(u^{m,l})-g^{m,l})\bigg)+\mu^{m}\|K^{m}u^{m,l}-y^{m,l}\|_{\mathcal{Y}}^{r}\bigg]$
$~~+\mathcal{R}_{0}(\varphi^{m},u^{m},u_{0}^{m},g^{m})+\nu^{m}\|\theta^{m}\|+\|f_{\theta^{m}}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f_{\theta^{m}}\|_{L^{\infty}(U)}$
$\leq$ $\sum_{1\leq l\leq L}\left[\lambda^{m}\|\frac{\partial}{\partial t}u^{\dagger,l}-F(t,u^{\dagger,l},\varphi^{\dagger,l})-f_{\tilde{\theta}^{m}}(t,\mathcal{J}_{\kappa}u^{\dagger,l})\|_{\mathcal{W}}^{q}+\mu^{m}\delta(m)^{r}\right]$
$~~+\mathcal{R}_{0}(\varphi^{\dagger},u^{\dagger},u_{0}^{\dagger},g^{\dagger})+\nu^{m}\|\tilde{\theta}^{m}\|+\|f_{\tilde{\theta}^{m}}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f_{\tilde{\theta}^{m}}\|_{L^{\infty}(U)}.$ (28)
We may further estimate one part of the sum on the right hand side of (3) by
$\sum_{1\leq l\leq L}\|\frac{\partial}{\partial t}u^{\dagger,l}-F(t,u^{\dagger,l},\varphi^{\dagger,l})-f_{\tilde{\theta}^{m}}(t,\mathcal{J}_{\kappa}u^{\dagger,l})\|_{\mathcal{W}}^{q}\\ =\sum_{1\leq l\leq L}\|f^{\dagger}(t,\mathcal{J}_{\kappa}u^{\dagger,l})-f_{\tilde{\theta}^{m}}(t,\mathcal{J}_{\kappa}u^{\dagger,l})\|_{\mathcal{W}}^{q}\leq c\|f^{\dagger}-f_{\tilde{\theta}^{m}}\|_{L^{\infty}(U)}^{q}\leq cm^{-\beta q}$
where in the penultimate estimation we have used $(t,\mathcal{J}_{\kappa}u^{\dagger,l})\in U$ which follows by Proposition 29 together with (24), and in the last step Assumption 5, iv). By
$\lim_{m\to\infty}\|f_{\tilde{\theta}^{m}}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f_{\tilde{\theta}^{m}}\|_{L^{\infty}(U)}=\|f^{\dagger}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f^{\dagger}\|_{L^{\infty}(U)},$
due to Assumption 5, iv), and the choice of the $\lambda^{m},\mu^{m},\nu^{m}$ we derive that the right hand side of (3) converges to
$\mathcal{R}_{0}(\varphi^{\dagger},u^{\dagger},u_{0}^{\dagger},g^{\dagger})+\|f^{\dagger}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f^{\dagger}\|_{L^{\infty}(U)}$ (29)
as $m\to\infty$ which is exactly the objective functional of problem ($\mathcal{P}^{\dagger}$). Using that $(\varphi^{\dagger},u^{\dagger},u_{0}^{\dagger},g^{\dagger},f^{\dagger})$ is the minimizer to ($\mathcal{P}^{\dagger}$) we can estimate (29) from above by
$\mathcal{R}_{0}(\hat{\varphi},\hat{u},\hat{u}_{0},\hat{g})+\|\hat{f}\|_{L^{\rho}(U)}^{\rho}+\|\nabla\hat{f}\|_{L^{\infty}(U)}$
for admissible $\hat{f}\in W^{1,\infty}(\mathbb{R}^{D})^{N}$, $\hat{u}\in\mathcal{V}^{N\times L}$, $\hat{\varphi}\in X_{\varphi}^{N\times L}$, $\hat{u}_{0}\in H^{N\times L}$, $\hat{g}\in\mathcal{B}^{N\times L}$ according to Assumption 5, iii). As a consequence, for $m$ sufficiently large it follows by Assumption 5, i) and vi), that
$\|\mathcal{J}_{\kappa}u^{m}\|_{L^{\infty}((0,T)\times\Omega)}$ $\leq c_{\mathcal{V}}\|u^{m}\|_{\mathcal{V}}\leq c_{\mathcal{V}}\pi(\mathcal{R}_{0}(\varphi^{m},u^{m},u_{0}^{m},g^{m}))$
$\leq c_{\mathcal{V}}\pi(\mathcal{R}_{0}(\hat{\varphi},\hat{u},\hat{u}_{0},\hat{g})+\|\hat{f}\|_{L^{\rho}(U)}^{\rho}+\|\nabla\hat{f}\|_{L^{\infty}(U)}+1).$ (30)
Hence, we derive that $(t,\mathcal{J}_{\kappa}u^{m,l})\in U$ for $m$ sufficiently large by monotonicity of $\pi$ and $\|\hat{f}\|_{L^{\rho}(U)}^{\rho}+\|\nabla\hat{f}\|_{L^{\infty}(U)}\leq\|\hat{f}\|_{L^{\rho}(\mathbb{R}^{D})}^{\rho}+\|\nabla\hat{f}\|_{L^{\infty}(\mathbb{R}^{D})}$. By convergence of the right hand side of (3) the terms $\|\varphi^{m}\|_{X_{\varphi}}$ and $\|u^{m}\|_{\mathcal{V}}$ are bounded due to coercivity of $\mathcal{R}_{0}$. Similarly boundedness of $\|f_{\theta^{m}}\|_{W^{1,\infty}(U)}$ follows using the norm equivalence of $\|\cdot\|_{W^{1,\infty}(U)}$ and $\|\cdot\|_{L^{\rho}(U)}+|\cdot|_{W^{1,\infty}(U)}$. Boundedness of $\|u_{0}^{m}\|_{H}$ follows as $\lambda^{m}\to\infty$ as $m\to\infty$ together with boundedness of $\|u^{m}(0)\|_{H}$, which holds by boundedness of $\|u^{m}\|_{\mathcal{V}}$, and continuity of the initial condition map shown in Appendix C, II. Finally by $\lambda^{m}\to\infty$ as $m\to\infty$, coercivity of $\mathcal{D}_{\text{BC}}$, boundedness of $\gamma$ and boundedness of the $u^{m}$, also boundedness of $\|g^{m}\|_{\mathcal{B}}$ can be inferred. As a consequence of reflexivity of $X_{\varphi}^{N\times L}$, $\mathcal{V}^{N\times L},H^{N\times L},\mathcal{B}^{N\times L}$ and the fact that $W^{1,\infty}(U)^{N}$ is the dualspace of a separable space, we derive that there exist weakly convergent subsequences (w.l.o.g. the whole sequences as we will see subsequently that the limit is unique) and $\tilde{\varphi}\in X_{\varphi}^{N\times L},\tilde{u}\in\mathcal{V}^{N\times L},\tilde{u}_{0}\in H^{N\times L},\tilde{g}\in\mathcal{B}^{N\times L}$ and similarly a weak-* convergent subsequence and $\tilde{f}\in W^{1,\infty}(U)^{N}$ with $\varphi^{m}\rightharpoonup\tilde{\varphi}$, $u^{m}\rightharpoonup\tilde{u}$, $u_{0}^{m}\rightharpoonup\tilde{u}_{0}$, $g^{m}\rightharpoonup\tilde{g}$, $f_{\theta^{m}}\overset{*}{\rightharpoonup}\tilde{f}$ as $m\to\infty$ (by [15, Theorem 3.18] and Banach-Alaoglu-Bourbaki e.g. in [15, Theorem 3.16]). By weak lower semicontinuity and weak-* lower semicontinuity together with the previous considerations we derive
$\mathcal{R}_{0}(\tilde{\varphi},\tilde{u},\tilde{u}_{0},\tilde{g})+\|\tilde{f}\|_{L^{\rho}(U)}^{\rho}+\|\nabla\tilde{f}\|_{L^{\infty}(U)}\\ \leq\liminf_{m\to\infty}\mathcal{R}_{0}(\varphi^{m},u^{m},u_{0}^{m},g^{m})+\|f^{m}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f^{m}\|_{L^{\infty}(U)}\\ \leq\mathcal{R}_{0}(\varphi^{\dagger},u^{\dagger},u_{0}^{\dagger},g^{\dagger})+\|f^{\dagger}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f^{\dagger}\|_{L^{\infty}(U)}$ (31)
We argue that $\tilde{u}=u^{\dagger}$: As the right hand side of (3) converges it holds true that
$K^{m}u^{m,l}-y^{m,l}\to 0~~~\text{strongly in }~\mathcal{Y}~\text{ as }~m\to\infty$ (32)
due to $\mu^{m}\to\infty$ as $m\to\infty$. The following estimation shows that $K^{m}u^{m,l}$ converges to $K^{\dagger}\tilde{u}^{l}$ as $m\to\infty$. Due to the triangle inequality we have that
$\|K^{m}u^{m,l}-K^{\dagger}\tilde{u}^{l}\|_{\mathcal{Y}}$ $\leq\|K^{m}u^{m,l}-K^{\dagger}u^{m,l}\|_{\mathcal{Y}}+\|K^{\dagger}u^{m,l}-K^{\dagger}\tilde{u}^{l}\|_{\mathcal{Y}}.$
Employing the convergence condition in (10), ensuring that the first term on the right hand side converges to zero, and weak-strong continuity of $K^{\dagger}$, implying $K^{\dagger}u^{m,l}\to K^{\dagger}\tilde{u}^{l}$ in $\mathcal{Y}$ as $m\to\infty$, we recover that indeed
$K^{m}u^{m,l}\to K^{\dagger}\tilde{u}^{l}~~~\text{strongly in }~\mathcal{Y}~\text{ as }~m\to\infty.$ (33)
Thus, by (27), the convergences (32), and (33), together with Assumption 5, v), and
$\|K^{\dagger}\tilde{u}^{l}-K^{\dagger}u^{\dagger,l}\|_{\mathcal{Y}}\leq\|K^{\dagger}\tilde{u}^{l}-K^{m}u^{m,l}\|_{\mathcal{Y}}+\|K^{m}u^{m,l}-y^{m,l}\|_{\mathcal{Y}}\\ +\|y^{m,l}-K^{m}u^{\dagger,l}\|_{\mathcal{Y}}+\|K^{m}u^{\dagger,l}-K^{\dagger}u^{\dagger,l}\|_{\mathcal{Y}}$
we derive $K^{\dagger}\tilde{u}^{l}=K^{\dagger}u^{\dagger,l}$. As a consequence of injectivity of $K^{\dagger}$ we finally derive that $\tilde{u}=u^{\dagger}$. We argue next that $\tilde{u}_{0}=u_{0}^{\dagger}$. For that, note once more that by convergence of the right hand side of (3) and $\lambda^{m}\to\infty$ as $m\to\infty$ we obtain that $u^{m}(0)-u_{0}^{m}\to 0$ in $H^{N\times L}$ as $m\to\infty$. As $u_{0}^{m}\rightharpoonup\tilde{u}_{0}$ in $H^{N\times L}$ as $m\to\infty$ we recover that $u^{m}(0)\rightharpoonup\tilde{u}_{0}$ in $H^{N\times L}$ as $m\to\infty$. Together with $u^{m}\rightharpoonup u^{\dagger}$, by what we have just shown, and weak closedness of the initial condition evaluation verified in II. of Appendix C, we obtain that indeed $\tilde{u}_{0}=u^{\dagger}(0)=u_{0}^{\dagger}$. By similar arguments and the assumption that $\mathcal{D}_{\text{BC}}(z)=0$ for $z\in\mathcal{B}^{N}$ iff $z=0$ we obtain that $\gamma(u^{m,l})-g^{m,l}\to 0$ in $\mathcal{B}^{N}$ as $m\to\infty$. As $g^{m}\rightharpoonup\tilde{g}$ and $\gamma(u^{m,l})\rightharpoonup\gamma(u^{\dagger,l})=g^{\dagger,l}$ by continuity of $\gamma$, both in $\mathcal{B}^{N}$ as $m\to\infty$, it also holds $\tilde{g}=g^{\dagger}$. It remains to show $\tilde{\varphi}=\varphi^{\dagger}$ and $\tilde{f}=f^{\dagger}$. Using the already discussed identities for $\tilde{u},\tilde{u}_{0}$ and $\tilde{g}$, estimation (31) yields
$\mathcal{R}_{0}(\tilde{\varphi},u^{\dagger},u_{0}^{\dagger},g^{\dagger})+\|\tilde{f}\|_{L^{\rho}(U)}^{\rho}+\|\nabla\tilde{f}\|_{L^{\infty}(U)}\\ \leq\mathcal{R}_{0}(\varphi^{\dagger},u^{\dagger},u_{0}^{\dagger},g^{\dagger})+\|f^{\dagger}\|_{L^{\rho}(U)}^{\rho}+\|\nabla f^{\dagger}\|_{L^{\infty}(U)}.$
Moreover, as the right hand side of (3) converges as $m\to\infty$, it holds true that
$\lim_{m\to\infty}\sum_{1\leq l\leq L}\|\frac{\partial}{\partial t}u^{m,l}-F(t,u^{m,l},\varphi^{m,l})-f_{\theta^{m}}(t,\mathcal{J}_{\kappa}u^{m,l})\|_{\mathcal{W}}^{q}=0$ (34)
due to $\lambda^{m}\to\infty$ as $m\to\infty$. We argue that
$\frac{\partial}{\partial t}u^{m,l}-F(t,u^{m,l},\varphi^{m,l})-f_{\theta^{m}}(t,\mathcal{J}_{\kappa}u^{m,l})\rightharpoonup\frac{\partial}{\partial t}u^{\dagger,l}-F(t,u^{\dagger,l},\tilde{\varphi}^{l})-\tilde{f}(t,\mathcal{J}_{\kappa}u^{\dagger,l})$
as $m\to\infty$ in $\mathcal{W}^{N}$, which together with (34) and weak lower semicontinuity of the $\|\cdot\|_{\mathcal{W}}$-norm implies that
$\frac{\partial}{\partial t}u^{\dagger,l}=F(t,u^{\dagger,l},\tilde{\varphi}^{l})+\tilde{f}(t,\mathcal{J}_{\kappa}u^{\dagger,l}).$ (35)
By Assumption 5, viii), and the considerations in Appendix C, I. showing weak continuity of the temporal derivative, it follows that
$\frac{\partial}{\partial t}u^{m,l}-F(t,u^{m,l},\varphi^{m,l})\rightharpoonup\frac{\partial}{\partial t}u^{\dagger,l}-F(t,u^{\dagger,l},\tilde{\varphi}^{l})$ (36)
[…]
$\|f_{\theta^{m}}(t,\mathcal{J}_{\kappa}u^{m,l})-\tilde{f}(t,\mathcal{J}_{\kappa}u^{m,l})\|_{\mathcal{W}}\leq c\|f_{\theta^{m}}-\tilde{f}\|_{L^{\infty}(U)}$
[…]
As a consequence, since $u^{m}\rightharpoonup u^{\dagger}$ in $\mathcal{V}^{N\times L}$ as $m\to\infty$ we derive that $u^{m,l}_{n}\to u^{\dagger,l}_{n}$ in $L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))$ strongly (w.l.o.g. for the whole sequence).
Thus, it suffices to show that $\tilde{f}(t,\mathcal{J}_{\kappa}u^{m,l})\to\tilde{f}(t,\mathcal{J}_{\kappa}u^{\dagger,l})$ in $L^{q}(0,T;L^{\hat{p}}(\Omega))^{N}\hookrightarrow\mathcal{W}^{N}$ as $m\to\infty$. Due to $\tilde{f}\in W^{1,\infty}(U)^{N}$, it induces a well-defined Nemytskii operator $\tilde{f}$ with $\tilde{f}(\mathcal{J}_{\kappa}u)](t,x)=\tilde{f}(t,\mathcal{J}_{\kappa}u(t,x))$ for $u\in L^{p}(0,T;L^{\hat{p}}(\Omega))^{N}$ and a.e. $(t,x)\in(0,T)\times\Omega$. Hence, we derive for $m$ large enough such that $(t,\mathcal{J}_{\kappa}u^{m,l})\in U$,
$\|\tilde{f}(t,\mathcal{J}_{\kappa}u^{m,l})-\tilde{f}(t,\mathcal{J}_{\kappa}u^{\dagger,l})\|_{L^{q}(0,T;L^{\hat{p}}(\Omega))}\leq c\|\tilde{f}\|_{W^{1,\infty}(U)}\|u^{m,l}-u^{\dagger,l}\|_{L^{q}(0,T;W^{\kappa,\hat{p}}(\Omega))}$
for some constant $c>0$ and thus, the left hand side approaches zero as $m\to\infty$.
With this, identity (35) follows and by (31) together with uniqueness of the solution of ($\mathcal{P}^{\dagger}$) that also $\tilde{\varphi}=\varphi^{\dagger}$ and $\tilde{f}=f^{\dagger}$, which concludes the proof. ∎
Remark 31.
Given that the last result essentially corresponds to a classical convergence result for inverse problems, an interesting future research direction is to what extent variational source conditions such as in [109] can be used to also obtain convergence rates here. Furthermore, also the viewpoint of statistical inverse problems (see e.g. [109] and [92]) on this setting is a relevant future research direction.
4 Conclusions
In this work, we have considered the problem of learning structured models from data in an all-at-once framework. That is, the state, the nonlinearity and physical parameters, constituting the unknowns of a PDE system, are identified simultaneously based on noisy measured data of the state. It is shown that the main identification problem is wellposed in a general setup. The main results of this work are i) unique reconstructability of the state, the unknown nonlinearity and the parameters of the known physical term as regularization-minimizing solutions of a limit problem with full measurements, and ii) that reconstructions of these quantities based on incomplete, noisy measurements approximate the unique regularization-minimizing solutions truth in the limit. For that, the class of functions used to approximate the unknown nonlinearity must meet a regularity and approximation capacity condition. These conditions are discussed and ensured for the case of fully connected feed forward neural networks.
The results of this work provide a general framework that guarantees unique reconstructability in the limit of a practically useful all-at-once formulation in learning PDE models. This is particularly interesting because uniqueness of the quantities of interest is not given in general, but rather under certain conditions on the class of approximating functions and for certain regularization functionals. This provides an analysis-based guideline on which minimal conditions need to be ensured by practical implementations of PDE-based model learning setups in order to expect unique recovery of regularization-minimizing solutions in the limit.
Appendix A Neural networks
In the following section we will provide proofs for Proposition 19, treating Assumption 3, and Proposition 20, addressing Assumption 5, ii) and vi), both results dealing with neural networks as introduced in Definition 16.
A.1 Proof of Proposition 19 (Assumption 3 for neural networks)
We start by proving the first part of Proposition 19 on the induction of well-defined Nemytskii operators.
Lemma 32.
Let Assumption 2 hold true. Suppose that $\sigma\in\mathcal{C}(\mathbb{R},\mathbb{R})$ is Lipschitz continuous with constant $L_{\sigma}$ (w.l.o.g. $L_{\sigma}\geq 1$). Then $\mathcal{N}_{\theta}:(0,T)\times(\otimes_{k=0}^{\kappa}\mathbb{R}^{p_{k}})^{N}\to\mathbb{R}$ induces a well-defined Nemytskii operator $\mathcal{N}_{\theta}:(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}\to L^{p}(0,T;L^{\hat{q}}(\Omega))$ via $\mathcal{N}_{\theta}(u)](t)=\mathcal{N}_{\theta}(u(t,\cdot))$. The same applies to $\mathcal{N}_{\theta}:(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}\to\mathcal{W}$.
Proof.
First note that $\mathcal{N}_{\theta}$ is Lipschitz continuous with some Lipschitz constant
$L_{\theta}\leq L_{\sigma}^{L-1}\prod_{l=1}^{L}|w^{l}|_{\infty}.$ (37)
Hereinafter for $1\leq\alpha\leq\infty$ we denote by $\alpha^{*}$ the corresponding dual exponent defined by $\alpha^{*}:=\frac{\alpha}{\alpha-1}$ if $\alpha\in(0,\infty)$, $\alpha^{*}:=1$ if $\alpha=\infty$ and $\alpha^{*}=\infty$ if $\alpha=1$. Now fixing some $c\geq\|\mathcal{N}_{\theta}(0,0)\|_{L^{\hat{q}}(\Omega)}$ we have for $u=((u_{1}^{k})_{k},\dots,(u_{N}^{k})_{k})\in(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}$ and a.e. $t\in(0,T)$ that
$\|\mathcal{N}_{\theta}(t,u(t,\cdot))\|_{L^{\hat{q}}(\Omega)}$ $\leq\|\mathcal{N}_{\theta}(0,0)\|_{L^{\hat{q}}(\Omega)}+\|\mathcal{N}_{\theta}(t,u(t,\cdot))-\mathcal{N}_{\theta}(0,0)\|_{L^{\hat{q}}(\Omega)}$
$\leq c+\sup_{\begin{subarray}{c}\varphi\in L^{\hat{q}^{*}}(\Omega),\\ \|\varphi\|_{L^{\hat{q}^{*}}(\Omega)}\leq 1\end{subarray}}\langle\mathcal{N}_{\theta}(t,u(t,\cdot))-\mathcal{N}_{\theta}(0,0),\varphi\rangle_{L^{\hat{q}}(\Omega),L^{\hat{q}^{*}}(\Omega)}$
$\leq c+\sup_{\begin{subarray}{c}\varphi\in L^{\hat{q}^{*}}(\Omega),\\ \|\varphi\|_{L^{\hat{q}^{*}}(\Omega)}\leq 1\end{subarray}}\int_{\Omega}|\mathcal{N}_{\theta}(t,u(t,x))-\mathcal{N}_{\theta}(0,0)||\varphi(x)|\mathop{}\!\mathrm{d}x$
$\leq c+L_{\theta}\sup_{\begin{subarray}{c}\varphi\in L^{\hat{q}^{*}}(\Omega),\\ \|\varphi\|_{L^{\hat{q}^{*}}(\Omega)}\leq 1\end{subarray}}\int_{\Omega}(T+|u(t,x)|_{1})|\varphi(x)|\mathop{}\!\mathrm{d}x$
$\leq c+L_{\theta}(T|\Omega|^{1/\hat{q}}+\sum_{\begin{subarray}{c}1\leq n\leq N\\ 0\leq k\leq\kappa\end{subarray}}\|u_{n}^{k}(t)\|_{L^{\hat{q}}(\Omega)^{p_{k}}})$
where the product norms correspond to the respective $\|\cdot\|_{1}$-norm. As $V\hookrightarrow L^{\hat{p}}(\Omega)\hookrightarrow L^{\hat{q}}(\Omega)$ and $\|u_{n}^{0}(t)\|_{V}<\infty$ for a.e. $t\in(0,T)$ due to $(u_{n}^{0})_{n}\in\mathcal{V}^{N}\subseteq L^{p}(0,T;V)^{N}$ it holds true that $\|u_{n}^{0}(t)\|_{L^{\hat{q}}(\Omega)}<\infty$ for $1\leq n\leq N$. The embedding $V_{k}\hookrightarrow L^{\hat{q}}(\Omega)$ implies $V_{k}^{\times}\hookrightarrow L^{\hat{q}}(\Omega)^{p_{k}}$ by which we may infer again that $\|u_{n}^{k}(t)\|_{L^{\hat{q}}(\Omega)^{p_{k}}}<\infty$ for a.e. $t\in(0,T)$ as $u_{n}^{k}\in\mathcal{V}^{\times}_{k}=L^{p}(0,T;V_{k}^{\times})$ for $1\leq n\leq N$, $1\leq k\leq\kappa$. Thus, it holds for a.e. $t\in(0,T)$ that $\mathcal{N}_{\theta}(t,u(t,\cdot))\in L^{\hat{q}}(\Omega)$ which is separable. Now $t\mapsto\mathcal{N}_{\theta}(t,u(t,\cdot))$ is weakly measurable, i.e.,
$t\mapsto\int_{\Omega}\mathcal{N}_{\theta}(t,u(t,x))w(x)\mathop{}\!\mathrm{d}x$
is Lebesgue measurable for all $w\in L^{\hat{q}^{*}}(\Omega)$ which follows by standard arguments as $\mathcal{N}_{\theta}$ is continuous, $w,u(t,\cdot)$ Lebesgue measurable and measurability is preserved under integration. Employing Pettis Theorem (see [98, Theorem 1.34]) we obtain that $t\mapsto\mathcal{N}_{\theta}(t,u(t,\cdot))\in L^{\hat{q}}(\Omega)$ is Bochner measurable. Similarly as before one can show that for $u=((u_{1}^{k})_{k},\dots,(u_{N}^{k})_{k})\in(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}$ it holds for some generic $\tilde{c}>0$,
$\|\mathcal{N}_{\theta}(u)\|_{L^{p}(0,T;L^{\hat{q}}(\Omega))}\leq\tilde{c}(1+\sum_{\begin{subarray}{c}1\leq n\leq N\\ 0\leq k\leq\kappa\end{subarray}}\|u_{n}^{k}\|_{L^{p}(0,T;L^{\hat{q}}(\Omega)^{p_{k}})})\leq\tilde{c}(1+\sum_{\begin{subarray}{c}1\leq n\leq N\\ 0\leq k\leq\kappa\end{subarray}}\|u_{n}^{k}\|_{\mathcal{V}_{k}^{\times}})<\infty$ (38)
again by $V_{k}\hookrightarrow L^{\hat{q}}(\Omega)$ using the isomorphism $L^{p}(0,T;L^{\hat{q}}(\Omega))^{p_{k}}\cong L^{p}(0,T;L^{\hat{q}}(\Omega)^{p_{k}})$ for $0\leq k\leq\kappa$. Finally, we derive by separability of $L^{\hat{q}}(\Omega)$ that $\mathcal{N}_{\theta}(u)$ is Bochner integrable (see [98, Section 1.5]) and by $p\geq q$ together with $L^{\hat{q}}(\Omega)\hookrightarrow W$ that also the Nemytskii operator $\mathcal{N}_{\theta}:(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}\to\mathcal{W}$ is well-defined. ∎
The next result addresses the remaining part of Proposition 19 on continuity.
Lemma 33 (Strong-strong continuity of $\mathcal{N}$).
Assume that $\sigma\in\mathcal{C}(\mathbb{R},\mathbb{R})$ is Lipschitz continuous with Lipschitz constant $L_{\sigma}$ (w.l.o.g. $L_{\sigma}\geq 1$). Then under Assumption 2, $\mathcal{N}:\Theta\times(\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}}))^{N}\to L^{q}(0,T;L^{\hat{q}}(\Omega))$, $(\theta,v)\mapsto\mathcal{N}_{\theta}(v)$ is strongly-strongly continuous.
Proof.
By analogous reasoning as in Lemma 32 the Nemytskii operator $\mathcal{N}$ in the assertions of this lemma is well-defined.
Let $(\theta^{m},u^{m})\to(\theta,u)$ in $\Theta\times(\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}}))^{N}$ as $m\to\infty$. We aim to show that $\mathcal{N}(\theta^{m},u^{m})\to\mathcal{N}(\theta,u)$ strongly in $L^{q}(0,T;L^{\hat{q}}(\Omega))$ as $m\to\infty$.
Note that for $z\in\mathbb{R}^{1+N\sum_{k=0}^{\kappa}p_{k}}$ it holds
$\mathcal{N}(\theta,z)$ $=(L_{\theta_{L}}\circ\dots\circ L_{\theta_{1}})(z),$
$\mathcal{N}(\theta^{m},z)$ $=(L_{\theta^{m}_{L}}\circ\dots\circ L_{\theta^{m}_{1}})(z)$
and define for $1\leq s\leq L-1$ the feed-forward neural networks $\mathcal{N}_{s}(\theta^{m},\theta,z)$ by
$\mathcal{N}_{s}(\theta^{m},\theta,z)$ $=(L_{\theta_{L}}\circ\dots\circ L_{\theta_{L-s+1}}\circ L_{\theta^{m}_{L-s}}\circ\dots\circ L_{\theta_{1}^{m}})(z),$
$\mathcal{N}_{0}(\theta^{m},\theta,z)$ $=\mathcal{N}(\theta^{m},z),$
$\mathcal{N}_{L}(\theta^{m},\theta,z)$ $=\mathcal{N}(\theta,z).$
By $\theta^{m}\to\theta$ as $m\to\infty$ and continuity of $\theta^{m}\mapsto(L_{\theta_{s}^{m}}\circ\dots\circ L_{\theta_{1}^{m}})(0)$ for all $s=1,\ldots,L$ there exists $C>0$, used generically in the estimations below, with
$|\mathfrak{L}_{s}^{m}(0)|_{\infty}<C,~~~\forall 1\leq s\leq L,$
for sufficiently large $m\in\mathbb{N}$, where we set
$\mathfrak{L}_{s}^{m}=L_{\theta_{s}^{m}}\circ\dots\circ L_{\theta_{1}^{m}}$
for $1\leq s\leq L$ and $\mathfrak{L}_{0}^{m}=\text{id}$ the identity map. Recall that we aim to estimate
$\|\mathcal{N}(\theta^{m},u^{m})-\mathcal{N}(\theta,u)\|_{L^{q}(0,T;L^{\hat{q}}(\Omega))}.$
For $M>0$ such that $L_{\sigma}^{L-1}\prod_{l=1}^{L}(\|w^{l}\|_{\infty}+1)<M$, we have for a.e. $(t,x)\in(0,T)\times\Omega$ (under abuse of notation omitting the dependence of $u,u^{m}$ on $(t,x)$) that $|\mathcal{N}(\theta^{m},t,u^{m})-\mathcal{N}(\theta,t,u)|$ is bounded by
$|\mathcal{N}(\theta^{m},t,u^{m})-\mathcal{N}(\theta^{m},t,u)|+|\mathcal{N}(\theta^{m},t,u)-\mathcal{N}(\theta,t,u)|\\ \leq M|u-u^{m}|_{1}+\sum_{s=0}^{L-1}|\mathcal{N}_{s+1}(\theta^{m},\theta,t,u)-\mathcal{N}_{s}(\theta^{m},\theta,t,u)|.$ (39)
For the second term estimate first $|\mathcal{N}_{s+1}(\theta^{m},\theta,t,u)-\mathcal{N}_{s}(\theta^{m},\theta,t,u)|$ by
$=|(L_{\theta_{L}}\circ\dots\circ L_{\theta_{L-s}}\circ\mathfrak{L}_{L-s-1}^{m})(t,u)-(L_{\theta_{L}}\circ\dots\circ L_{\theta_{L-s+1}}\circ L_{\theta_{L-s}^{m}}\circ\mathfrak{L}_{L-s-1}^{m})(t,u)|$
$\leq\left(L_{\sigma}^{s-1}\prod_{l=L-s+1}^{L}|w^{l}|_{\infty}\right)|(L_{\theta_{L-s}}\circ\mathfrak{L}_{L-s-1}^{m})(t,u)-(L_{\theta_{L-s}^{m}}\circ\mathfrak{L}_{L-s-1}^{m})(t,u)|_{\infty}$
$\leq\left(L_{\sigma}^{s}\prod_{l=L-s+1}^{L}|w^{l}|_{\infty}\right)\left[|w^{L-s}-w_{m}^{L-s}|_{\infty}|(\mathfrak{L}_{L-s-1}^{m})(t,u)|_{\infty}+|\beta^{L-s}-\beta_{m}^{L-s}|_{\infty}\right]$
$\leq\left(L_{\sigma}^{s}\prod_{l=L-s+1}^{L}|w^{l}|_{\infty}\right)|\theta^{L-s}-\theta_{m}^{L-s}|_{\infty}\left(|(\mathfrak{L}_{L-s-1}^{m})(t,u)-(\mathfrak{L}_{L-s-1}^{m})(0)|_{\infty}+C\right)$
$\leq\left(L_{\sigma}^{s}\prod_{l=L-s+1}^{L}|w^{l}|_{\infty}\right)|\theta^{L-s}-\theta_{m}^{L-s}|_{\infty}\left(L_{\sigma}^{L-s-1}\prod_{l=1}^{L-s-1}|w_{m}^{l}|_{\infty}(T+|u|_{1})+C\right)$
$\leq M|\theta^{L-s}-\theta_{m}^{L-s}|_{\infty}\left(|u|_{1}+C\right).$ (40)
Combining this with (39) it follows that
$|\mathcal{N}(\theta^{m},t,u^{m})-\mathcal{N}(\theta,t,u)|\leq M|u-u^{m}|_{1}+M(|u|_{1}+C)\sum_{s=1}^{L}|\theta^{s}-\theta^{s}_{m}|_{\infty}.$ (41)
To estimate $\|\mathcal{N}(\theta^{m},u^{m})-\mathcal{N}(\theta,u)\|_{L^{q}(0,T;L^{\hat{q}}(\Omega))}$ note that for $w^{*}\in L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))$ with $\|w^{*}\|_{L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))}\leq 1$ it holds for some generic constant $\tilde{C}>0$ by successively employing the upper bound (41), Minkowski’s inequality in $L^{\hat{q}}(\Omega)$ and Hölder’s inequality in time with $p,p^{*}$ that
$\int_{0}^{T}\|\mathcal{N}(\theta^{m},t,u^{m}(t,\cdot))-\mathcal{N}(\theta,t,u(t,\cdot))\|_{L^{\hat{q}}(\Omega)}\|w^{*}(t)\|_{L^{\hat{q}^{*}}(\Omega)}\mathop{}\!\mathrm{d}t$
$\leq$ $\tilde{C}\int_{0}^{T}\bigg\|\bigg[|u(t,\cdot)-u^{m}(t,\cdot)|_{1}+(|u(t,\cdot)|_{1}+C)\sum_{s=1}^{L}|\theta^{s}-\theta^{s}_{m}|_{\infty}\bigg]\bigg\|_{L^{\hat{q}}(\Omega)}\|w^{*}(t)\|_{L^{\hat{q}^{*}}(\Omega)}\mathop{}\!\mathrm{d}t$
$\leq$ $\tilde{C}\bigg[\|u-u^{m}\|_{(\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{q}}(\Omega)^{p_{k}}))^{N}}+(\|u\|_{(\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{q}}(\Omega)^{p_{k}}))^{N}}+C)\sum_{s=1}^{L}|\theta^{s}-\theta^{s}_{m}|_{\infty}\bigg]$
due to $\|w^{*}\|_{L^{p^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))}\leq 1$ as $p\geq q$ and $L^{\hat{p}}(\Omega)\hookrightarrow L^{\hat{q}}(\Omega)$. As the right hand side of the previous estimation is independent of $w^{*}$ we obtain that
$\|\mathcal{N}(\theta^{m},u^{m})-\mathcal{N}(\theta,u)\|_{L^{q}(0,T;L^{\hat{q}}(\Omega))}\\ \leq\tilde{C}\bigg[\|u-u^{m}\|_{(\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}}))^{N}}+(\|u\|_{(\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}}))^{N}}+C)\sum_{s=1}^{L}|\theta^{s}-\theta^{s}_{m}|_{\infty}\bigg].$
Now by $u_{m}\to u$ in $(\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}}))^{N}$, $\|u\|_{(\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}}))^{N}}<\infty$ and $\theta_{m}\to\theta$ as $m\to\infty$ we derive that the last argument converges to zero as $m\to\infty$.
Thus, it holds
$\mathcal{N}(\theta^{m},u^{m})\to\mathcal{N}(\theta,u)~~~\text{as}~~~m\to\infty~~~\text{in}~L^{q}(0,T;L^{\hat{q}}(\Omega))$
yielding strong-strong continuity of the joint operator $\mathcal{N}$ as claimed. ∎
Combining Lemma 32 and Lemma 33 concludes the result in Proposition 19.
A.2 Proof of Proposition 20 (Assumption 5 for neural networks)
First, we verify $W^{1,\infty}_{loc}-$regularity of the classes $\mathcal{F}_{n}^{m}$.
Lemma 34.
Assume that $\sigma\in\mathcal{C}(\mathbb{R},\mathbb{R})$ is locally Lipschitz continuous and let $(\mathcal{F}_{n}^{m})_{n}$ be given as in Definition 17. Then $\mathcal{F}_{n}^{m}\subseteq W^{1,\infty}_{\text{loc}}(\mathbb{R}^{1+N\sum_{k=0}^{\kappa}p_{k}})$ for $1\leq n\leq N$, $m\in\mathbb{N}$.
Proof.
Let $f\in\mathcal{F}_{n}^{m}$ for some $1\leq n\leq N$ and $m\in\mathbb{N}$. Since the activation function $\sigma$ is supposed to be locally Lipschitz continuous, $f$ is also locally Lipschitz continuous. This follows from the fact that continuous functions map bounded sets to bounded sets and thus, recalling the layer-wise structure of $f$ in Definition 16, for any bounded $U\subseteq\mathbb{R}^{1+N\sum_{k=0}^{\kappa}p_{k}}$ the instance $f$ is Lipschitz continuous on $U$ with a constant depending on local Lipschitz constants of $\sigma$ and norms of the weights. Rademacher’s Theorem finally yields $f\in W^{1,\infty}(U)$ and thus the assertion of the lemma. ∎
The next result shows for bounded $U\subseteq\mathbb{R}^{D}$ strong-strong continuity of the map $\Theta\ni\theta\mapsto\nabla\mathcal{N}_{\theta}\in L^{\infty}(U)$ for $\mathcal{C}^{1}$-regular and Lipschitz continuous activation functions, in particular implying weak lower semicontinuity as claimed in Proposition 20.
Lemma 35.
Let $U\subseteq\mathbb{R}^{D}$ be bounded. Furthermore, let the activation function $\sigma$ of the class of parameterized approximation functions fulfill $\sigma\in\mathcal{C}^{1}(\mathbb{R},\mathbb{R})$. Then the map
$\Theta\ni\theta\mapsto\nabla\mathcal{N}_{\theta}\in L^{\infty}(U)$
is strongly-strongly continuous.
Proof.
We consider first the case that $\sigma\in\mathcal{C}^{1}(\mathbb{R},\mathbb{R})$ is Lipschitz continuous with constant $L_{\sigma}$ (w.l.o.g. $L_{\sigma}\geq 1$). Let $(\theta^{m})_{m}\subseteq\Theta$ such that $\theta^{m}\to\theta\in\Theta$ as $m\to\infty$. Maintaining the notation in the proof of Lemma 33 we further set for $1\leq k\leq l\leq L$
$\mathfrak{L}_{k,l}=L_{\theta_{l}}\circ\dots\circ L_{\theta_{k}}$
with $\mathfrak{L}_{k,l}=\text{id}$ the identity map for $k>l$. Then we obtain for fixed $z\in U$ that
$|\nabla\mathcal{N}_{\theta^{m}}(z)-\nabla\mathcal{N}_{\theta}(z)|_{\infty}\leq\sum_{s=0}^{L-1}|\nabla\mathcal{N}_{s+1}(\theta^{m},\theta,z)-\nabla\mathcal{N}_{s}(\theta^{m},\theta,z)|_{\infty}\\ =\sum_{s=0}^{L-1}|\nabla[(\mathfrak{L}_{L-s,L}\circ\mathfrak{L}_{L-s-1}^{m})(z)]-\nabla[(\mathfrak{L}_{L-s+1,L}\circ\mathfrak{L}_{L-s}^{m})(z)]|_{\infty}.$
We consider a summand of the last sum for fixed $0\leq s\leq L-1$ and show convergence to zero for $m\to\infty$. For that we introduce the following simplifying notation for products of matrices $C_{0}\cdot\ldots\cdot C_{n}$ for $n\in\mathbb{N}$ where the row and column dimensions fit for the product to make sense, by
$\mathcal{P}_{l=0}^{n}C_{l}:=C_{0}\cdot\ldots\cdot C_{n}.$
Furthermore, we set $\mathcal{P}_{l=k}^{m}C_{l}:=1$ for $k>m$. Defining
$A_{l,s}^{m}(z)=\sigma^{\prime}(w^{L-l-1}(\mathfrak{L}_{L-s,L-l-2}\circ\mathfrak{L}_{L-s-1}^{m})(z)+b^{L-l-1})w^{L-l-1}~~\text{for }~0\leq l\leq s-1,$
$B_{l,s}^{m}(z)=\sigma^{\prime}(w^{L-l-1}(\mathfrak{L}_{L-s+1,L-l-2}\circ\mathfrak{L}_{L-s}^{m})(z)+b^{L-l-1})w^{L-l-1}~~\text{for }~0\leq l\leq s-2,$
$\text{and}~~B_{s-1,s}^{m}(z)=\sigma^{\prime}(w_{m}^{L-s}\mathfrak{L}_{L-s-1}^{m}(z)+b_{m}^{L-s})w_{m}^{L-s}$
for $z\in U$, we derive by the chain rule that
$|\nabla[(\mathfrak{L}_{L-s,L}\circ\mathfrak{L}_{L-s-1}^{m})(z)]-\nabla[(\mathfrak{L}_{L-s+1,L}\circ\mathfrak{L}_{L-s}^{m})(z)]|_{\infty}$
can be estimated by
$=|w^{L}(\mathcal{P}_{l=0}^{s-1}A_{l,s}^{m}(z)-\mathcal{P}_{l=0}^{s-1}B_{l,s}^{m}(z))\nabla[\mathfrak{L}_{L-s-1}^{m}(z)]|_{\infty}$
$\leq|w^{L}|_{\infty}|\nabla[\mathfrak{L}_{L-s-1}^{m}(z)]|_{\infty}\sum_{r=0}^{s-1}|(\mathcal{P}_{l=0}^{r-1}B_{l,s}^{m}(z))(A_{r,s}^{m}(z)-B_{r,s}^{m}(z))(\mathcal{P}_{l=r+1}^{s-1}A_{l,s}^{m}(z))|_{\infty}$
$\leq|w^{L}|_{\infty}|\nabla[\mathfrak{L}_{L-s-1}^{m}(z)]|_{\infty}\sum_{r=0}^{s-1}(\prod_{l=0}^{r-1}|B_{l,s}^{m}(z)|_{\infty})|A_{r,s}^{m}(z)-B_{r,s}^{m}(z)|_{\infty}(\prod_{l=r+1}^{s-1}|A_{l,s}^{m}(z)|_{\infty}).$ (42)
Let $M>0$ such that $L_{\sigma}^{L-1}\prod_{l=1}^{L}(|w^{l}|_{\infty}+1)<M$ and $m\in\mathbb{N}$ sufficiently large such that $|w^{l}_{m}-w^{l}|_{\infty}<1$ for $1\leq l\leq L$ which is possible due to $\theta^{m}\to\theta$ as $m\to\infty$. As $|A_{l,s}^{m}(z)|_{\infty},|B_{l,s}^{m}(z)|_{\infty}\leq L_{\sigma}M$ for $0\leq s\leq L-1,1\leq l\leq s-1$, $|w^{L}|_{\infty}<M$ and
$\nabla[\mathfrak{L}_{L-s-1}^{m}(z)]=\mathcal{P}_{l=0}^{L-s-2}\sigma^{\prime}(w^{L-s-l-1}_{m}\mathfrak{L}_{L-s-l-2}^{m}(z)+b^{L-s-l-1}_{m})w^{L-s-l-1}_{m}$
by the chain rule, implying $|\nabla[\mathfrak{L}_{L-s-1}^{m}(z)]|_{\infty}\leq L_{\sigma}^{L-s-1}M$, it remains to show that
$\lim_{m\to\infty}|A_{r,s}^{m}(z)-B_{r,s}^{m}(z)|_{\infty}=0.$ (43)
This follows as $\theta^{m}\to\theta$, $\mathfrak{L}_{L-s,L-l-2}\circ\mathfrak{L}_{L-s-1}^{m}\to\mathfrak{L}_{1,L-l-2}$ in $L^{\infty}(U)$ for $0\leq l\leq s-1$ and $\mathfrak{L}_{L-s+1,L-l-2}\circ\mathfrak{L}_{L-s}^{m}\to\mathfrak{L}_{1,L-l-2}$ in $L^{\infty}(U)$ for $0\leq l\leq s-2$ as $m\to\infty$ by similar considerations as in (A.1) due to continuity of $\sigma^{\prime}$. As the convergence in (43) holds uniformly for $z\in U$ we recover the assertion of the lemma that $\nabla\mathcal{N}_{\theta^{m}}\to\nabla\mathcal{N}_{\theta}\in L^{\infty}(U)$ as $m\to\infty$.
In case $\sigma\in\mathcal{C}^{1}(\mathbb{R},\mathbb{R})$ is not Lipschitz continuous (such as ReQU), the result follows by a similar strategy as above. An adaption concerns uniform boundedness of the $\sigma^{\prime}$ terms in $A_{l,s}^{m}(z),B_{l,s}^{m}(z)$ for $z\in U$, which follows from uniform convergence $\mathfrak{L}_{L-s+1,L-l-2}\circ\mathfrak{L}_{L-s}^{m}\to\mathfrak{L}_{1,L-l-2}$ in $L^{\infty}(U)$ as $m\to\infty$ and the fact that the latter map $U$ to bounded sets. ∎
Finally, in case of the activation being the Rectified Linear Unit, we recover for bounded $U\subseteq\mathbb{R}^{D}$ weak lower semicontinuity of the map $\Theta\ni\theta\mapsto\|\nabla\mathcal{N}_{\theta}\|_{L^{\infty}(U)}$ as claimed in Proposition 20.
Lemma 36.
Let $U\subseteq\mathbb{R}^{D}$ be bounded. Furthermore, let the activation function $\sigma$ of the class of parameterized approximation functions be the Rectified Linear Unit. Then for $(\theta^{m})_{m}\subseteq\Theta$ with $\theta^{m}\to\theta\in\Theta$ as $m\to\infty$ it holds
$\|\nabla\mathcal{N}_{\theta}\|_{L^{\infty}(U)}\leq\liminf_{m\to\infty}\|\nabla\mathcal{N}_{\theta^{m}}\|_{L^{\infty}(U)}.$
Proof.
Let $(\theta^{m})_{m}\subseteq\Theta$ with $\theta^{m}\to\theta\in\Theta$ as $m\to\infty$. We show that
$|\nabla\mathcal{N}_{\theta}(z)|_{\infty}\leq\liminf_{m\to\infty}|\nabla\mathcal{N}_{\theta^{m}}(z)|_{\infty}$ (44)
for a.e. $z\in U$ which further implies
$|\nabla\mathcal{N}_{\theta}(z)|_{\infty}\leq\operatorname*{ess\,sup}_{x\in U}\liminf_{m\to\infty}|\nabla\mathcal{N}_{\theta^{m}}(x)|_{\infty}\leq\liminf_{m\to\infty}\|\nabla\mathcal{N}_{\theta^{m}}\|_{L^{\infty}(U)}$
and the assertion of the lemma by taking the essential supremum over $z\in U$. Now for $z\in[(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]^{\circ}$ an inner point of the preimage of $\left\{0\right\}$ under $\nabla\mathcal{N}_{\theta}$. it holds that $\nabla\mathcal{N}_{\theta}(z)=0$ implying (44). It remains to verify (44) for $z\in[U\backslash(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]^{\circ}$ as the boundary $\partial[(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]$ is a zeroset in $\mathbb{R}^{D}$. Following the proof of Lemma 35 we recover the estimation in (A.2). Again as $\theta^{m}\to\theta$, $\mathfrak{L}_{L-s,L-l-2}\circ\mathfrak{L}_{L-s-1}^{m}\to\mathfrak{L}_{1,L-l-2}$ in $L^{\infty}(U)$ for $0\leq l\leq s-1$ and $\mathfrak{L}_{L-s+1,L-l-2}\circ\mathfrak{L}_{L-s}^{m}\to\mathfrak{L}_{1,L-l-2}$ in $L^{\infty}(U)$ for $0\leq l\leq s-2$ as $m\to\infty$ and $w^{k+1}\mathfrak{L}_{1,k}(z)+b^{k+1}\neq 0$ for $1\leq k\leq L-2$ due to $\nabla\mathcal{N}_{\theta}(z)\neq 0$, for $m$ sufficiently large we end up in the smooth regime of $\sigma^{\prime}$ such that the previous arguments yield $\lim_{m\to\infty}\nabla\mathcal{N}_{\theta^{m}}(z)=\nabla\mathcal{N}_{\theta}(z)$ for $z\in[U\backslash(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]^{\circ}$ implying (44) and concluding the assertions of the lemma. ∎
Combining the Lemmata 34, 35 and 36 concludes the result in Proposition 20.
A.3 Lifting technique
In this subsection we discuss a lifting approach that shows how approximation results, such as in Proposition 21, covering (9) can be used to infer Assumption 5, iv), i.e., also achieve the convergence $\|\nabla f_{\theta^{m}}\|_{L^{\infty}(U)}\to\|\nabla f\|_{L^{\infty}(U)}$ as $m\to\infty$. For that, one needs to impose higher regularity on $f$, such as $W^{2,\infty}$- or $\mathcal{C}^{2}$-regularity and assume that the bounded domain $U\subseteq\mathbb{R}^{D}$ of functions in $\mathcal{F}_{n}^{m}$ is star-shaped with some center given by $x_{0}\in U$. The basic idea is to approximate the partial derivatives of $f$ by the approximation result at hand (such as in Proposition 21) and lift the approximation property to the function. More concretely, let $g_{\tilde{\theta}^{m}}$ approximate $\nabla f$ uniformly on $U$ by rate $\beta>0$ and $f_{\tilde{\eta}^{m}}$ the function $f$ by rate $\gamma>0$. Then with $\text{diam}U$ denoting the diameter of $U$ we have
$\|f(x)-f_{\tilde{\eta}^{m}}(x_{0})-\int_{0}^{1}g_{\tilde{\theta}^{m}}(x_{0}+t(x-x_{0}))\cdot(x-x_{0})\mathop{}\!\mathrm{d}t\|_{L^{\infty}(U)}$
$\leq|f(x_{0})-f_{\tilde{\eta}^{m}}(x_{0})|+\operatorname*{ess\,sup}\limits_{x\in U}|\int_{0}^{1}((\nabla f-g_{\tilde{\theta}^{m}})(x_{0}+t(x-x_{0})))\cdot(x-x_{0})\mathop{}\!\mathrm{d}t|$
$\leq cm^{-\gamma}+cm^{-\beta}\text{diam}U.$
Furthermore, it holds true by the Leibniz integral rule that
$\|\nabla_{x}f-\nabla_{x}(f_{\tilde{\eta}^{m}}(x_{0})+\int_{0}^{1}g_{\tilde{\theta}^{m}}(x_{0}+t(x-x_{0}))\cdot(x-x_{0})\mathop{}\!\mathrm{d}t)\|_{L^{\infty}(U)}$
$=\|\nabla_{x}f-\int_{0}^{1}t\nabla g_{\tilde{\theta}^{m}}(x_{0}+t(x-x_{0}))\cdot(x-x_{0})+g_{\tilde{\theta}^{m}}(x_{0}+t(x-x_{0}))\mathop{}\!\mathrm{d}t\|_{L^{\infty}(U)}$
$=\|\nabla_{x}f-\int_{0}^{1}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}(tg_{\tilde{\theta}^{m}}(x_{0}+t(x-x_{0})))\mathop{}\!\mathrm{d}t\|_{L^{\infty}(U)}$
$=\|\nabla_{x}f-g_{\tilde{\theta}^{m}}(x)\|_{L^{\infty}(U)}$
$\leq cm^{-\beta}.$
Note that the Leibniz integral rule is applicable as $\int_{0}^{1}g_{\tilde{\theta}^{m}}(x_{0}+t(x-x_{0}))\cdot(x-x_{0})\mathop{}\!\mathrm{d}t$ is finite, $t\nabla g_{\tilde{\theta}^{m}}(x_{0}+t(x-x_{0}))\cdot(x-x_{0})+g_{\tilde{\theta}^{m}}(x_{0}+t(x-x_{0}))$ exists and is majorizable by $\text{diam}U\|g_{\tilde{\theta}^{m}}\|_{W^{1,\infty}(U)}$. This shows that $f$ is approximated by $f_{\tilde{\eta}^{m}}(x_{0})+\int_{0}^{1}g_{\tilde{\theta}^{m}}(x_{0}+t(x-x_{0}))\cdot(x-x_{0})\mathop{}\!\mathrm{d}t$ in $W^{1,\infty}(U)$ as $m\to\infty$ with rate given by $\min(\beta,\gamma)$.
Appendix B Physical term
In the following section we will provide proofs for Proposition 23 and Proposition 25, addressing Assumption 4 on the physical term both in the linear and nonlinear case.
B.1 Linear case
We prove Proposition 23, starting with the first part on Assumption 4, i), the induction of well-defined Nemytskii operators.
Lemma 37.
Let Assumption 2 hold true and $\tilde{V}\hookrightarrow W^{\omega,\hat{p}}(\Omega)$. Suppose that $t\mapsto\Phi_{n}(t,\varphi)$ and $t\mapsto\Psi(t,\varphi)$ are measurable for all $\varphi\in X_{\varphi}$ and $s_{\beta}$ fulfill (17). Assume that there exist functions $\mathcal{B}_{1},\mathcal{B}_{2}:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ that map bounded sets to bounded sets and $\phi\in L^{\frac{pq}{p-q}}(0,T)$ (with $\phi\in L^{\infty}(0,T)$ if $p=q$), $\psi\in L^{q}(0,T)$ such that
$\|\Phi_{n,\beta}(t,\varphi)\|_{L^{s_{\beta}}(\Omega)}\leq\phi(t)\mathcal{B}_{1}(\|\varphi\|_{X_{\varphi}}),~~~\|\Psi(t,\varphi)\|_{L^{\hat{q}}(\Omega)}\leq\psi(t)\mathcal{B}_{2}(\|\varphi\|_{X_{\varphi}}).$ (45)
Then $F$ in (16) induces a well-defined Nemytskii operator $F:\mathcal{V}^{N}\times X_{\varphi}\to\mathcal{W}$ with
$F((u_{n})_{1\leq n\leq N},\varphi)](t)=F(t,(u_{n}(t))_{1\leq n\leq N},\varphi)$
for $(u_{n})_{1\leq n\leq N}\in\mathcal{V}^{N},\varphi\in X_{\varphi}$ and $t\in(0,T)$.
Proof.
Employing similar arguments as in the proof of Lemma 32 together with measurability of $t\mapsto\Phi_{n}(t,\varphi)$ and $t\mapsto\Psi(t,\varphi)$ yields Bochner measurability of
$(0,T)\ni t\mapsto\Psi(t,\varphi(\cdot))+\sum_{n=1}^{N}\mathcal{J}_{\omega}u_{n}(t,\cdot)\cdot\Phi_{n}(t,\varphi(\cdot))\in W.$
Welldefinedness follows by the following chain of estimations for $u=(u_{n})_{1\leq n\leq N}\in\mathcal{V}^{N}$ and $\varphi\in X_{\varphi}$ for some generic constant $c>0$. By the embedding $L^{\hat{q}}(\Omega)\hookrightarrow W$ it holds $\|F(u,\varphi)\|_{\mathcal{W}}\leq c\|F(u,\varphi)\|_{L^{q}(0,T;L^{\hat{q}}(\Omega))}$ which by the definition of $F$ and the triangle inequality can be estimated by
$c\left(\sum_{n=1}^{N}\left(\int_{0}^{T}\|\mathcal{J}_{\omega}u_{n}(t,\cdot)\cdot\Phi_{n}(t,\varphi(\cdot))\|_{L^{\hat{q}}(\Omega)}^{q}\mathop{}\!\mathrm{d}t\right)^{1/q}+\left(\int_{0}^{T}\|\Psi(t,\varphi(\cdot))\|_{L^{\hat{q}}(\Omega)}^{q}\mathop{}\!\mathrm{d}t\right)^{1/q}\right).$
Due to the growth condition in (45) we may estimate the term
$\left(\int_{0}^{T}\|\Psi(t,\varphi(\cdot))\|_{L^{\hat{q}}(\Omega)}^{q}\mathop{}\!\mathrm{d}t\right)^{1/q}\leq\mathcal{B}_{2}(\|\varphi\|_{X_{\varphi}})\|\psi\|_{L^{q}(0,T)}<\infty.$
For the remaining part note that by [9, Theorem 6.1], [9, Corollary 6.3] and the choice of $s_{\beta}$ it holds true that the pointwise multiplication of functions is a continuous bilinear map
$W^{\omega-|\beta|,\hat{p}}(\Omega)\times L^{s_{\beta}}(\Omega)\to L^{\hat{q}}(\Omega).$
Thus, there exists some generic constant $c>0$ independent of $u_{n},t,\varphi,\Phi_{n}$ with
$\|\mathcal{J}_{\omega}u_{n}(t,\cdot)\cdot\Phi_{n}(t,\varphi(\cdot))\|_{L^{\hat{q}}(\Omega)}\leq c\sum_{0\leq|\beta|\leq\omega}\|D^{\beta}u_{n}(t,\cdot)\|_{W^{\omega-|\beta|,\hat{p}}(\Omega)}\|\Phi_{n,\beta}(t,\varphi(\cdot))\|_{L^{s_{\beta}}(\Omega)}.$
We employ (45) together with Hölder’s inequality to obtain
$\left(\int_{0}^{T}\|\mathcal{J}_{\omega}u_{n}(t,\cdot)\cdot\Phi_{n}(t,\varphi(\cdot))\|_{L^{\hat{q}}(\Omega)}^{q}\mathop{}\!\mathrm{d}t\right)^{1/q}\leq c\mathcal{B}_{1}(\|\varphi\|_{X_{\varphi}})\left(\int_{0}^{T}\|u_{n}\|_{W^{\omega,\hat{p}}(\Omega)}^{q}\phi(t)^{q}\mathop{}\!\mathrm{d}t\right)^{1/q}.$
Using Hölder’s inequality once more and $\mathcal{V}\hookrightarrow L^{p}(0,T;W^{\omega,\hat{p}}(\Omega))$ yields that
$\left(\int_{0}^{T}\|u_{n}\|_{W^{\omega,\hat{p}}(\Omega)}^{q}\phi(t)^{q}\mathop{}\!\mathrm{d}t\right)^{1/q}\leq c\|u_{n}\|_{L^{p}(0,T;W^{\omega,\hat{p}}(\Omega))}\|\phi\|_{L^{\frac{pq}{p-q}}(0,T)}\leq c\|u_{n}\|_{\mathcal{V}}\|\phi\|_{L^{\frac{pq}{p-q}}(0,T)}$
which is again finite by assumption. The case $s_{\beta}=\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}}$ can be covered similarly using $V\hookrightarrow W^{\omega,\hat{p}}(\Omega)$ and employing Hölder’s inequality. Finally, we derive that $\|F(u,\varphi)\|_{\mathcal{W}}<\infty$ which concludes the assertions of the lemma. ∎
The next result addresses the remaining part of Proposition 23 on continuity.
Lemma 38.
Let the assumptions of Lemma 37 hold true. Suppose that $\Psi(t,\cdot):X_{\varphi}\to L^{\hat{q}}(\Omega)$ is weakly continuous for almost every $t\in(0,T)$. Let $s_{\beta}$ be given as in Lemma 37, additionally with strict inequality $\frac{\omega-|\beta|}{d}>\frac{1}{\hat{p}}-\frac{1}{\hat{q}}+\frac{1}{s_{\beta}}$ if $\hat{q}=1$ or $s_{\beta}=\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}}$. Assume that $\Phi_{n,\beta}(t,\cdot):X_{\varphi}\to L^{s_{\beta}}(\Omega)$ is weakly continuous for a.e. $t\in(0,T)$. Furthermore, suppose that either $\omega\leq\kappa$ or otherwise in case $\omega>\kappa$ the following additional conditions hold:
[…]
    $\|\Phi_{n,\beta}(t,\varphi)\|_{\frac{c_{\beta}\hat{q}}{c_{\beta}-\hat{q}}}\leq\phi(t)\mathcal{B}_{1}(\|\varphi\|_{X_{\varphi}}).$
  • For $|\beta|=\omega$ assume that $\Phi_{n,\beta}(t,\cdot):X_{\varphi}\to L^{\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}}}(\Omega)$ is well-defined and weak-strong continuous for a.e. $t\in(0,T)$.
Then $\mathcal{V}^{N}\times X_{\varphi}\ni(u,\varphi)\mapsto F(u,\varphi)\in\mathcal{W}$ induced by (16) is weak-weak continuous.
Proof.
Let $(u^{k})_{k}\subseteq\mathcal{V}^{N},(\varphi^{k})_{k}\subseteq X_{\varphi}$ and $u\in\mathcal{V}^{N},\varphi\in X_{\varphi}$ with $u^{k}\rightharpoonup u$ in $\mathcal{V}^{N}$ and $\varphi^{k}\rightharpoonup\varphi$ in $X_{\varphi}$ as $k\to\infty$. We verify that $F(u^{k},\varphi^{k})\rightharpoonup F(u,\varphi)$ in $\mathcal{W}$ as $k\to\infty$. First, by $L^{\hat{q}}(\Omega)\hookrightarrow W$ and the growth condition in (45) it holds true for $w^{*}\in\mathcal{W}^{*}$ and a.e. $t\in[0,T]$ that
$\langle\Psi(t,\varphi^{k})-\Psi(t,\varphi),w^{*}(t)\rangle_{W,W^{*}}$ $\leq c(\|\Psi(t,\varphi^{k})\|_{L^{\hat{q}}(\Omega)}+\|\Psi(t,\varphi)\|_{L^{\hat{q}}(\Omega)})\|w^{*}(t)\|_{W^{*}}$
$\leq c(\mathcal{B}_{2}(\|\varphi^{k}\|_{X_{\varphi}})+\mathcal{B}_{2}(\|\varphi\|_{X_{\varphi}}))\psi(t)\|w^{*}(t)\|_{W^{*}}.$
By $\varphi^{k}\rightharpoonup\varphi$ in $X_{\varphi}$ the $\|\varphi^{k}\|_{X_{\varphi}}$ are uniformly bounded for all $k$. Thus, as $\mathcal{B}_{2}$ maps bounded sets to bounded sets there exists some $\tilde{c}$ such that $\mathcal{B}_{2}(\|\varphi^{k}\|_{X_{\varphi}})+\mathcal{B}_{2}(\|\varphi\|_{X_{\varphi}})\leq\tilde{c}$ for all $k$ and we derive that $\langle\Psi(t,\varphi^{k})-\Psi(t,\varphi),w^{*}(t)\rangle_{W,W^{*}}$ is majorized by the integrable function $t\mapsto\tilde{c}\psi(t)\|w^{*}(t)\|_{W^{*}}$ independently of $k$ with
$\int_{0}^{T}\langle\Psi(t,\varphi^{k})-\Psi(t,\varphi),w^{*}(t)\rangle_{W,W^{*}}\mathop{}\!\mathrm{d}t\leq\tilde{c}\|\psi\|_{L^{q}(0,T)}\|w^{*}\|_{\mathcal{W}^{*}}<\infty$
by Hölder’s inequality. The Dominated Convergence Theorem and weak-weak continuity of $\Psi$ for a.e. $t\in(0,T)$ yield that $\langle\Psi(\cdot,\varphi^{k})-\Psi(\cdot,\varphi),w^{*}\rangle_{\mathcal{W},\mathcal{W}^{*}}\to 0$ as $k\to\infty$ and hence, that $\Psi(\cdot,\varphi^{k})\rightharpoonup\Psi(\cdot,\varphi)$ in $\mathcal{W}$. Thus, by (16) it remains to show that
$\mathcal{J}_{\omega}u^{k}_{n}\cdot\Phi_{n}(\cdot,\varphi^{k})\rightharpoonup\mathcal{J}_{\omega}u_{n}\cdot\Phi_{n}(\cdot,\varphi)\qquad\text{for}~1\leq n\leq N$ (46)
in $L^{q}(0,T;L^{\hat{q}}(\Omega))$ as $k\to\infty$ which is sufficient due to $L^{q}(0,T;L^{\hat{q}}(\Omega))\hookrightarrow\mathcal{W}$. Since
$\mathcal{J}_{\omega}u^{k}_{n}\cdot\Phi_{n}(\cdot,\varphi^{k})-\mathcal{J}_{\omega}u_{n}\cdot\Phi_{n}(\cdot,\varphi)=\sum_{0\leq|\beta|\leq\omega}\left(D^{\beta}u_{n}^{k}\cdot\Phi_{n,\beta}(\cdot,\varphi^{k})-D^{\beta}u_{n}\cdot\Phi_{n,\beta}(\cdot,\varphi)\right)$
by (16) it suffices to prove that for any fixed $0\leq|\beta|\leq\omega$ and $1\leq n\leq N$
$D^{\beta}u_{n}^{k}\cdot\Phi_{n,\beta}(\cdot,\varphi^{k})\rightharpoonup D^{\beta}u_{n}\cdot\Phi_{n,\beta}(\cdot,\varphi)$ (47)
in $L^{q}(0,T;L^{\hat{q}}(\Omega))$ as $k\to\infty$. We show first that
$D^{\beta}u_{n}\cdot\left(\Phi_{n,\beta}(\cdot,\varphi^{k})-\Phi_{n,\beta}(\cdot,\varphi)\right)\rightharpoonup 0$ (48)
in $L^{q}(0,T;L^{\hat{q}}(\Omega))$ as $k\to\infty$ and then
$\left(D^{\beta}u_{n}^{k}-D^{\beta}u_{n}\right)\cdot\Phi_{n,\beta}(\cdot,\varphi^{k})\rightharpoonup 0$ (49)
in $L^{q}(0,T;L^{\hat{q}}(\Omega))$ as $k\to\infty$, proving the weak convergence in (47). For that, let $w^{*}\in L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))$. Then for a.e. $t\in(0,T)$ it holds that $D^{\beta}u_{n}(t)\in W^{\omega-|\beta|,\hat{p}}(\Omega)$ (with $W^{0,\hat{p}}(\Omega)=L^{\hat{p}}(\Omega)$) and $w^{*}(t)\in L^{\hat{q}^{*}}(\Omega)$. By [9, Theorem 6.1, Corollary 6.3] the inclusion $D^{\beta}u_{n}(t)w^{*}(t)\in L^{r_{\beta}}(\Omega)$ holds true with $\frac{\hat{p}\hat{q}}{\hat{q}-\hat{p}+\hat{p}\hat{q}}\leq r_{\beta}\leq\hat{q}^{*}$ and $r_{\beta}^{-1}\geq\frac{1}{\hat{p}}+\frac{1}{\hat{q}^{*}}-\frac{\omega-|\beta|}{d}$ (with strict inequality if $\hat{q}=1$). In particular by the requirements on $s_{\beta}$ we may choose $r_{\beta}=s_{\beta}^{*}$ (or equivalently $r_{\beta}^{*}=s_{\beta}$). Using that $D^{\beta}u_{n}\in L^{p}(0,T;W^{\omega-|\beta|,\hat{p}}(\Omega))$ and $w^{*}\in L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))$ we derive
$D^{\beta}u_{n}w^{*}\in L^{\frac{pq^{*}}{p+q^{*}}}(0,T;L^{r_{\beta}}(\Omega)).$
Thus, we obtain by the growth condition (45) that for $w^{*}\in L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))$
$\langle\Phi_{n,\beta}(t,\varphi^{k})-\Phi_{n,\beta}(t,\varphi),$ $D^{\beta}u_{n}(t)w^{*}(t)\rangle_{L^{r_{\beta}^{*}}(\Omega),L^{r_{\beta}}(\Omega)}$
$\leq\|\Phi_{n,\beta}(t,\varphi^{k})-\Phi_{n,\beta}(t,\varphi)\|_{L^{r_{\beta}^{*}}(\Omega)}\|D^{\beta}u_{n}(t)w^{*}(t)\|_{L^{r_{\beta}}(\Omega)}$
$\leq c\phi(t)\|D^{\beta}u_{n}(t)w^{*}(t)\|_{L^{r_{\beta}}(\Omega)}$
for $0\leq|\beta|\leq\omega$ and a.e. $t\in[0,T]$. Hence, independently of $k$, the term
$\langle\Phi_{n,\beta}(t,\varphi^{k})-\Phi_{n,\beta}(t,\varphi),D^{\beta}u_{n}(t)w^{*}(t)\rangle_{L^{r_{\beta}^{*}}(\Omega),L^{r_{\beta}}(\Omega)}$
is majorized by the integrable function $t\mapsto c\phi(t)\|D^{\beta}u_{n}(t)w^{*}(t)\|_{L^{r_{\beta}}(\Omega)}$ with
$\int_{0}^{T}\langle\Phi_{n,\beta}(t,\varphi^{k})-\Phi_{n,\beta}(t,\varphi),D^{\beta}u_{n}(t)w^{*}(t)\rangle_{L^{r_{\beta}^{*}}(\Omega),L^{r_{\beta}}(\Omega)}\mathop{}\!\mathrm{d}t\\ \leq c\|\phi\|_{L^{\frac{pq}{p-q}}(\Omega)}\|D^{\beta}u_{n}w^{*}\|_{L^{\frac{pq^{*}}{p+q^{*}}}(0,T;L^{r_{\beta}}(\Omega))}<\infty$
as $(\frac{pq^{*}}{p+q^{*}})^{*}=\frac{pq}{p-q}$. Employing dominated convergence together with weak continuity of $\Phi_{n,\beta}(t,\cdot):X_{\varphi}\to L^{r_{\beta}^{*}}(\Omega)=L^{s_{\beta}}(\Omega)$ for a.e. $t\in(0,T)$ concludes (48). The case that $s_{\beta}=\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}}$ can be similarly dealt with as before using that $D^{\beta}u_{n}(t)w^{*}(t)\in L^{(\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}})^{*}}(\Omega)$ for $w^{*}(t)\in L^{\hat{q}^{*}}(\Omega)$ by Hölder’s generalized inequality. Next we prove the weak convergence in (49) which follows if we can show for $w^{*}\in L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))$ that
$\int_{0}^{T}\langle\left(D^{\beta}u_{n}^{k}(t)-D^{\beta}u_{n}(t)\right)\cdot\Phi_{n,\beta}(t,\varphi^{k}),w^{*}(t)\rangle_{L^{\hat{q}}(\Omega),L^{\hat{q}^{*}}(\Omega)}\mathop{}\!\mathrm{d}t\to 0$ (50)
as $k\to\infty$. In fact due to Hölder’s inequality, the growth condition in (45) and similar arguments regarding the multiplication operator as in Lemma 37 we obtain that the integrand of (50) can be bounded from above by
$c\phi(t)\mathcal{B}_{1}(\|\varphi^{k}\|_{X_{\varphi}})\|u_{n}^{k}(t)-u_{n}(t)\|_{W^{\omega,\hat{p}}(\Omega)}\|w^{*}(t)\|_{L^{\hat{q}^{*}}(\Omega)}$ (51)
for some generic constant $c>0$ for a.e. $t\in(0,T)$. Using that $\mathcal{V}\hookrightarrow\mathcal{C}(0,T;W^{\omega,\hat{p}}(\Omega))$ by [98, Lemma 7.1] and the assumption $\tilde{V}\hookrightarrow W^{\omega,\hat{p}}(\Omega)$, we derive that $u_{n}^{k}\rightharpoonup u_{n}$ in $\mathcal{C}(0,T;W^{\omega,\hat{p}}(\Omega))$ as $k\to\infty$ which by boundedness of weakly convergent sequences in Banach spaces (see [15, Proposition 3.5 (iii)]) implies that $\|u_{n}^{k}(t)-u_{n}(t)\|_{W^{\omega,\hat{p}}(\Omega)}$ is bounded independently of $t\in(0,T)$ and $k\in\mathbb{N}$. Using uniform boundedness of $\|\varphi^{k}\|_{X_{\varphi}}$ for all $k\in\mathbb{N}$ implies that (51) and hence the integrand of (50) is majorized by the function $t\mapsto c\phi(t)\|w^{*}(t)\|_{L^{\hat{q}^{*}}(\Omega)}$ which is integrable since
$\int_{0}^{T}\phi(t)\|w^{*}(t)\|_{L^{\hat{q}^{*}}(\Omega)}\mathop{}\!\mathrm{d}t\leq T^{1/p}\|\phi\|_{L^{\frac{pq}{p-q}}(\Omega)}\|w^{*}\|_{L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))}$
by Hölder’s inequality. This can be similarly shown to hold true in case $s_{\beta}=\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}}$ using Hölder’s generalized inequality. We show that the integrand of (50) converges to zero pointwise for a.e. $t\in(0,T)$ under the following case distinction.
[…]
$c\phi(t)\mathcal{B}_{1}(\|\varphi^{k}\|_{X_{\varphi}})\|D^{\beta}u_{n}^{k}(t)-D^{\beta}u_{n}(t)\|_{L^{c_{\beta}}(\Omega)}\|w^{*}(t)\|_{L^{\hat{q}^{*}}(\Omega)}$
[…]
Thus, we recover (47) and consequently (46). Finally, this implies that $F(u^{k},\varphi^{k})\rightharpoonup F(u,\varphi)$ in $\mathcal{W}$ concluding weak continuity as stated in the assertion of the lemma. ∎
Combining Lemma 37 and Lemma 38 concludes the result in Proposition 23.
Remark 39.
[…]
B.2 Nonlinear case
We prove Proposition 25, starting with the first part on Assumption 4, i), the induction of well-defined Nemytskii operators.
Lemma 40.
Let Assumption 2 and the extended state space embedding
$\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$
hold true. Suppose that the $F_{n}(\cdot,\cdot,\varphi):(0,T)\times V^{N}\to W$ satisfy the Carathéodory condition, i.e., $t\mapsto F_{n}(t,v,\varphi)$ is measurable for $v\in V^{N}$ and $v\mapsto F_{n}(t,v,\varphi)$ is continuous for a.e. $t\in(0,T)$. Further assume that the $F_{n}$ satisfy the growth condition
$\|F_{n}(t,(v_{n})_{1\leq n\leq N},\varphi)\|_{W}\leq\mathcal{B}_{0}(\|\varphi\|_{X_{\varphi}},\sum_{n=1}^{N}\|v_{n}\|_{H})(\Gamma(t)+\sum_{n=1}^{N}\|v_{n}\|_{V})$ (52)
for some $\Gamma\in L^{q}(0,T)$ and $\mathcal{B}_{0}:\mathbb{R}^{2}\to\mathbb{R}$, increasing in the second entry and, for fixed second entry, mapping bounded sets to bounded sets. Then the $F_{n}:(0,T)\times V^{N}\times X_{\varphi}\to W$ induce well-defined Nemytskii operators $F_{n}:\mathcal{V}^{N}\times X_{\varphi}\to\mathcal{W}$ with
$[F_{n}(v,\varphi)](t)=F_{n}(t,v(t),\varphi)$ (53)
for $v\in\mathcal{V}^{N}$ and $\varphi\in X_{\varphi}$.
Proof.
The Carathéodory assumption ensures Bochner measurability of the map $t\mapsto F_{n}(t,v(t),\varphi)$ for $v\in\mathcal{V}^{N}$ and $\varphi\in X_{\varphi}$. Growth condition (52) and Hölder’s inequality imply that for $v\in\mathcal{V}^{N}$ and $\varphi\in X_{\varphi}$ the term $\int_{0}^{T}\|F_{n}(t,v(t),\varphi)\|_{W}^{q}\mathop{}\!\mathrm{d}t$ can be bounded, for $C>0$ some in the following generically used constant, by
$C\int_{0}^{T}\mathcal{B}_{0}(\|\varphi\|_{X_{\varphi}},\sum_{n=1}^{N}\|v_{n}(t)\|_{H})^{q}(|\Gamma(t)|^{q}+\sum_{n=1}^{N}\|v_{n}(t)\|_{V}^{q})\mathop{}\!\mathrm{d}t$
which may be further estimated by
$C\mathcal{B}_{0}(\|\varphi\|_{X_{\varphi}},\sum_{n=1}^{N}\|v_{n}\|_{\mathcal{C}(0,T;H)})^{q}(\|\Gamma\|_{L^{q}(0,T)}^{q}+\sum_{n=1}^{N}\int_{0}^{T}\|v_{n}(t)\|_{V}^{q}\mathop{}\!\mathrm{d}t).$ (54)
Monotonicity of $\mathcal{B}_{0}$ in its second entry, $v_{n}\in\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$, $\Gamma\in L^{q}(0,T)$ and
$\int_{0}^{T}\|v_{n}(t)\|_{V}^{q}\mathop{}\!\mathrm{d}t\leq T^{\frac{p-q}{p}}\|v_{n}\|^{q}_{L^{p}(0,T;V)}\leq T^{\frac{p-q}{p}}\|v_{n}\|^{q}_{\mathcal{V}}<\infty$
yield that (54) is finite. As a consequence, we derive that $\int_{0}^{T}\|F_{n}(t,v(t),\varphi)\|_{W}^{q}\mathop{}\!\mathrm{d}t<\infty$ and thus, that $\|F_{n}(v,\varphi)\|_{\mathcal{W}}<\infty$ which together with separability of $W$ implies Bochner integrability of $t\mapsto F_{n}(t,v(t),\varphi)$ and well-definedness of the Nemytskii operator $F_{n}:\mathcal{V}^{N}\times X_{\varphi}\to\mathcal{W}$ concluding the assertions of the lemma. ∎
The next result addresses the remaining part of Proposition 25. The proof is essentially based on [1, Lemma 5], for which the requirements of Lemma 40 are extended by a stronger growth condition.
Lemma 41.
Let Assumption 2 and the extended state space embedding
$\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$
hold true. Suppose that the $F_{n}(\cdot,\cdot,\varphi):(0,T)\times V^{N}\to W$ fulfill the Carathéodory condition as in Lemma 40 and weak-weak continuity of
$F_{n}(t,\cdot):H^{N}\times X_{\varphi}$ $\to W$
$(v_{1},\dots,v_{N},\varphi)$ $\mapsto F_{n}(t,v_{1},\dots,v_{N},\varphi)$
for a.e. $t\in(0,T)$. Further assume that the $F_{n}$ satisfy the stricter growth condition
$\|F_{n}(t,(v_{n})_{1\leq n\leq N},\varphi)\|_{W}\leq\mathcal{B}_{0}(\|\varphi\|_{X_{\varphi}},\sum_{n=1}^{N}\|v_{n}\|_{H})(\Gamma(t)+\sum_{n=1}^{N}\|v_{n}\|_{H})$ (55)
for some $\Gamma\in L^{q}(0,T)$ and $\mathcal{B}_{0}:\mathbb{R}^{2}\to\mathbb{R}$, increasing in the second entry and, for fixed second entry, mapping bounded sets to bounded sets. Then the Nemytskii operator in (53) is weak-weak continuous.
Proof.
First note that, for $(u_{n})_{n}\in\mathcal{V}^{N},\psi\in X_{\varphi}$ and $t\in(0,T)$, the growth condition (55) together with $\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$ and monotonicity of $\mathcal{B}_{0}$ yields
$\|F_{n}(u_{1},\dots,u_{N},\psi)(t)\|_{W}\leq\mathcal{B}_{0}(\|\psi\|_{X_{\varphi}},\sum_{n=1}^{N}\|u_{n}\|_{\mathcal{C}(0,T;H)})(\Gamma(t)+\sum_{n=1}^{N}\|u_{n}(t)\|_{H}).$ (56)
Now let $(v,\varphi)\in\mathcal{V}^{N}\times X_{\varphi}$ and $(v^{m})_{m}\subseteq\mathcal{V}^{N},(\varphi^{m})_{m}\subseteq X_{\varphi}$ with $v^{m}\rightharpoonup v$ in $\mathcal{V}^{N}$ and $\varphi^{m}\rightharpoonup\varphi$ in $X_{\varphi}$. We show
$F_{n}(v_{1}^{m},\dots,v_{N}^{m},\varphi^{m})\rightharpoonup F_{n}(v_{1},\dots,v_{N},\varphi)~~\text{ in }~\mathcal{W}.$ (57)
Boundedness of weakly convergent sequences (see e.g. [15, Proposition 3.5 (iii)]) and $\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$ together with the assumptions on $\mathcal{B}_{0}$ ensure the existence of $c_{\varphi},c_{v}>0$ such that both $\mathcal{B}_{0}(\|\varphi\|_{X_{\varphi}},\sum_{n=1}^{N}\|v_{n}\|_{\mathcal{C}(0,T;H)})\leq\mathcal{B}_{0}(c_{\varphi},c_{v})$ and
$\sup_{m\in\mathbb{N}}\mathcal{B}_{0}(\|\varphi^{m}\|_{X_{\varphi}},\sum_{n=1}^{N}\|v_{n}^{m}\|_{\mathcal{C}(0,T;H)})\leq\mathcal{B}_{0}(c_{\varphi},c_{v})$ (58)
hold true. Fixing $w^{*}\in\mathcal{W}^{*}$ and using (56) and (58) it follows for a.e. $t\in[0,T]$ that
$\langle F_{n}(v_{1}^{m},$ $\dots,v_{N}^{m},\varphi^{m})(t)-F_{n}(v_{1},\dots,v_{N},\varphi)(t),w^{*}(t)\rangle_{W,W^{*}}$
$\leq(\|F_{n}(v_{1}^{m},\dots,v_{N}^{m},\varphi^{m})(t)\|_{W}+\|F_{n}(v_{1},\dots,v_{N},\varphi)(t)\|_{W})\|w^{*}(t)\|_{W^{*}}$
$\leq\mathcal{B}_{0}(c_{\varphi},c_{v})(\Gamma(t)+\sum_{n=1}^{N}\|v_{n}(t)\|_{H}+\sum_{n=1}^{N}\|v_{n}^{m}(t)\|_{H})\|w^{*}(t)\|_{W^{*}}$
$\leq\mathcal{B}_{0}(c_{\varphi},c_{v})(|\Gamma(t)|+2c_{v})\|w^{*}(t)\|_{W^{*}}.$
As a consequence, for $c\geq\mathcal{B}(c_{\varphi},c_{v})$ the function
$t\mapsto\langle F_{n}(v_{1}^{m},\dots,v_{N}^{m},\varphi^{m})(t)-F_{n}(v_{1},\dots,v_{N},\varphi)(t),w^{*}(t)\rangle_{W,W^{*}}$
is majorized by the integrable function $t\mapsto c(|\Gamma(t)|+2c_{v})\|w^{*}(t)\|_{W^{*}}$ with
$\langle F_{n}(v_{1}^{m},\dots,v_{N}^{m},\varphi^{m})-F_{n}(v_{1},\dots,v_{N},\varphi),w^{*}\rangle_{\mathcal{W},\mathcal{W}^{*}}\\ \leq c\int_{0}^{T}(|\Gamma(t)|+2c_{v})\|w^{*}(t)\|_{W^{*}}\mathop{}\!\mathrm{d}t\leq c(\|\Gamma\|_{L^{q}(0,T)}+2c_{v}T^{1/q})\|w^{*}\|_{\mathcal{W}^{*}}<\infty$
as $p\geq q$. Thus, once we argue weak convergence
$F_{n}(t,v_{1}^{m}(t),\dots,v_{N}^{m}(t),\varphi^{m})\rightharpoonup F_{n}(t,v_{1}(t),\dots,v_{N}(t),\varphi)$ (59)
in $W$ for a.e. $t\in(0,T)$, weak convergence in (57) follows by the Dominated Convergence Theorem. For the former, note that by $\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$ the pointwise evaluation map realizing $u(t)\in H$ for $u\in\mathcal{V}$ is weakly closed due to
$\|u(t)\|_{H}\leq\|u\|_{\mathcal{C}(0,T;H)}\leq c\|u\|_{\mathcal{V}}$
for $t\in(0,T)$. By $v^{m}\rightharpoonup v$ in $\mathcal{C}(0,T;H)^{N}$ it holds true that $(\|v^{m}(t)\|_{H})_{m}$ is bounded for $t\in(0,T)$. Thus employing weak closedness of the evaluation map yields that every subsequence and hence, the whole sequence $v^{m}(t)$ converges weakly $v^{m}(t)\rightharpoonup v(t)$ in $H^{N}$. This together with weak-weak continuity of $F_{n}(t,\cdot):H^{N}\times X_{\varphi}\to W$ implies the convergence stated in (59) and finally, the assertion of the lemma. ∎
Combining Lemma 40 and Lemma 41 concludes the result in Proposition 25.
Appendix C Existence of minimizers
In this section we verify wellposedness of the minimization problem in (2) under the Assumptions 2, 3, 4. As first step, we show that (2) is indeed well-defined by proving that, for any $f_{\theta_{n},n}\in\mathcal{F}_{n}^{m}$, the composed function $(t,u)\mapsto f_{\theta_{n},n}(t,\mathcal{J}_{\kappa}u_{1},\dots,\mathcal{J}_{\kappa}u_{N})$ for $u\in V^{N}$ induces a well-defined Nemytskii operator on the dynamic space for $n=1,\dots,N$ and similarly the trace map $\gamma$. For that we consider first the differential operator introduced in (6).
Lemma 42.
Let Assumption 2 hold true. Then the function $\mathcal{J}_{\kappa}:W^{\kappa,\hat{p}}(\Omega)\to\otimes_{k=0}^{\kappa}L^{\hat{p}}(\Omega)^{p_{k}}$ induces a well-defined Nemytskii operator $\mathcal{J}_{\kappa}:L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))\to\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}})$ with
$\mathcal{J}_{\kappa}v](t)=\mathcal{J}_{\kappa}v(t)$
for $v\in L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))$. Furthermore, it is weak-weak continuous.
Proof.
We show first that for fixed $\beta\in\mathbb{N}_{0}^{d}$ with $0\leq k:=|\beta|\leq\kappa$ the differential operator $D^{\beta}:W^{\kappa,\hat{p}}(\Omega)\to L^{\hat{p}}(\Omega)$ induces a well-defined Nemytskii operator $D^{\beta}:L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))\to L^{p}(0,T;L^{\hat{p}}(\Omega))$ with $D^{\beta}v](t)=D^{\beta}v(t)$ for $v\in L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))$. To that end let $v\in L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))$. By Assumption 2 we derive that $v(t,\cdot)\in W^{\kappa,\hat{p}}(\Omega)$ for a.e. $t\in(0,T)$. Thus, it follows that
$\|D^{\beta}v(t,\cdot)\|_{L^{\hat{p}}(\Omega)}\leq\|v(t,\cdot)\|_{W^{\kappa,\hat{p}}(\Omega)}<\infty$ (60)
for a.e. $t\in(0,T)$. As in particular $v\in L^{1}(0,T;W^{\kappa,\hat{p}}(\Omega))$ is Bochner measurable there exist temporal simple functions $v_{k}$ approximating $v$ pointwise a.e. in $(0,T)$ in the strong sense of $W^{\kappa,\hat{p}}(\Omega)$. Employing the embedding $W^{\kappa,\hat{p}}(\Omega)\hookrightarrow L^{\hat{p}}(\Omega)$ yields that the temporal simple functions $D^{\beta}v_{k}$ approximate $D^{\beta}v$ pointwise a.e. in $(0,T)$ in the strong sense of $L^{\hat{p}}(\Omega)$ and hence, Bochner measurability of
$(0,T)\ni t\mapsto D^{\beta}v(t,\cdot)\in L^{\hat{p}}(\Omega).$
Similar to (60) well-definedness of the Nemytskii operator $D^{\beta}:L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))\to L^{p}(0,T;L^{\hat{p}}(\Omega))$ with $D^{\beta}v](t)=D^{\beta}v(t)$ for $v\in L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))$ follows.
Weak-weak continuity of $D^{\beta}:L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))\to L^{p}(0,T;L^{\hat{p}}(\Omega))$ follows by boundedness and linearity where the latter follows immediately from linearity of the differential operator $D^{\beta}$. To see boundedness let $w\in L^{p^{*}}(0,T;L^{\hat{p}^{*}}(\Omega))$. Then by (60) we derive for some $c>0$ that
$\langle D^{\beta}v,w\rangle_{L^{p}(0,T;L^{\hat{p}}(\Omega)),L^{p^{*}}(0,T;L^{\hat{p}^{*}}(\Omega))}$ $=\int_{0}^{T}\langle D^{\beta}v(t),w(t)\rangle_{L^{\hat{p}}(\Omega),L^{\hat{p}^{*}}(\Omega)}\mathop{}\!\mathrm{d}t$
$\leq c\int_{0}^{T}\|v(t)\|_{W^{\kappa,\hat{p}}(\Omega)}\|w(t)\|_{L^{\hat{p}^{*}}(\Omega)}\mathop{}\!\mathrm{d}t$
$\leq c\|v\|_{L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))}\|w\|_{L^{p^{*}}(0,T;L^{\hat{p}^{*}}(\Omega))}$
proving that $\|D^{\beta}v\|_{L^{p}(0,T;L^{\hat{p}}(\Omega))}\leq c\|v\|_{L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))}$.
As a consequence, for fixed $0\leq k\leq\kappa$ the function $J^{k}:W^{\kappa,\hat{p}}(\Omega)\to L^{\hat{p}}(\Omega)^{p_{k}}$ in (7) induces a well-defined Nemytskii operator $J^{k}:L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))\to L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}})$ with $J^{k}v](t)=J^{k}v(t)$ for $v\in L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))$ which is linear and bounded and thus, weak-weak continuous. This is straightforward as $J^{k}$ is the Cartesian product of finitely many functions which by the previous considerations induce well-defined Nemytskii operators sharing the property of weak-weak continuity, respectively. The same arguments yield the assertion of the lemma that $\mathcal{J}_{\kappa}$ induces a well-defined Nemytskii operator $\mathcal{J}_{\kappa}:L^{p}(0,T;W^{\kappa,\hat{p}}(\Omega))\to\otimes_{k=0}^{\kappa}L^{p}(0,T;L^{\hat{p}}(\Omega)^{p_{k}})$ which is weak-weak continuous. ∎
By minor adaptions of the previous proof it is straightforward to show that indeed also the Nemytskii operator $\mathcal{J}_{\kappa}:\mathcal{V}\to\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times}$ is well-defined. Employing Assumption 3, i) we obtain that $(t,u)\mapsto f_{\theta_{n},n}(t,\mathcal{J}_{\kappa}u_{1},\dots,\mathcal{J}_{\kappa}u_{N})$ for $u\in V^{N}$ induces a well-defined Nemytskii operator with
$[f_{\theta_{n},n}(\mathcal{J}_{\kappa}u_{1},\dots,\mathcal{J}_{\kappa}u_{N})](t)(x)=f_{\theta_{n},n}(t,\mathcal{J}_{\kappa}u_{1}(t,x),\dots,\mathcal{J}_{\kappa}u_{N}(t,x))$
for $u\in\mathcal{V}^{N}$ and $t\in(0,T)$. On basis of the previous considerations we recover the following continuity result.
Lemma 43.
In the setup of Assumption 2 and Assumption 3 it holds that
$\Theta_{n}^{m}\times\mathcal{V}^{N}\ni(\theta_{n},u)\mapsto f_{n}(\theta_{n},u)=:f_{\theta_{n},n}(\mathcal{J}_{\kappa}u_{1},\dots,\mathcal{J}_{\kappa}u_{N})\in\mathcal{W}$
is weak-weak continuous for $n=1,\dots,N$.
Proof.
[…]
Lastly, it remains to show that the trace map $\gamma$ induces a well-defined Nemytskii operator on the extended space.
Lemma 44.
Let Assumption 2 hold true. Then the trace map $\gamma:V\to B$ induces a well-defined Nemytskii operator $\gamma:\mathcal{V}\to\mathcal{B}$ with $\gamma(v)](t)=\gamma(v(t))$ for $v\in\mathcal{V}$. Furthermore, it is weak-weak continuous.
Proof.
By Assumption 2, iv), the map $\gamma$ is continuous. Together with separability of the spaces $V,B$ and $p\geq s$ we derive by [88, Theorem 1.43] that $\gamma$ induces a well defined Nemytskii operator $\gamma:L^{p}(0,T;V)\to L^{s}(0,T;B)=\mathcal{B}$ which is continuous. Employing $\mathcal{V}\hookrightarrow L^{p}(0,T;V)$ and linearity of $\gamma$ concludes the proof. ∎
As a consequence together with the considerations in Section 2 the terms occurring in problem (2) are well-defined. In view of wellposedness of the minimization problem (2) we follow [1]. For that purpose define for $1\leq l\leq L$ the maps $G^{l}$ by
$G^{l}:X_{\varphi}^{N\times L}\times\mathcal{V}^{N\times L}\times\otimes_{n}\Theta_{n}^{m}\times H^{N\times L}\times\mathcal{B}^{N\times L}\to\mathcal{W}^{N}\times H^{N}\times\mathcal{B}^{N}\times\mathcal{Y}$
where $(\varphi,u,\theta,u_{0},g)$ is mapped to
$(\frac{\partial}{\partial t}u^{l}-F(t,u^{l},\varphi^{l})-f_{\theta}(t,\mathcal{J}_{\kappa}u^{l}),u^{l}(0)-u_{0}^{l},\gamma(u^{l})-g^{l},K^{m}u^{l})$
with $\varphi=(\varphi_{n}^{l})_{\begin{subarray}{c}1\leq n\leq N\\ 1\leq l\leq L\end{subarray}}\subseteq X_{\varphi},u=(u_{n}^{l})_{\begin{subarray}{c}1\leq n\leq N\\ 1\leq l\leq L\end{subarray}}\subseteq\mathcal{V},u_{0}=(u_{0,n}^{l})_{\begin{subarray}{c}1\leq n\leq N\\ 1\leq l\leq L\end{subarray}}\subseteq H$ and $\theta\in\otimes_{n}\Theta_{n}^{m}$. Recall that, notation wise, we use direct vectorial extensions over $n=1,\dots,N$. Furthermore, define for the domain of definition given by $\mathbf{D}(G):=X_{\varphi}^{N\times L}\times\mathcal{V}^{N\times L}\times\otimes_{n}\Theta_{n}\times H^{N\times L}\times\mathcal{B}^{N\times L}$ the operator
$G:\hskip 21.33955pt\mathbf{D}(G)$ $\to\mathcal{W}^{N\times L}\times H^{N\times L}\times\mathcal{B}^{N\times L}\times\mathcal{Y}^{L}$ (61)
$(\varphi,u,\theta,u_{0},g)$ $\mapsto(G^{l}(\varphi,u,\theta,u_{0},g))_{1\leq l\leq L}.$
For $\lambda,\mu\in\mathbb{R}_{+}$ we define the map $\|\cdot\|_{\lambda,\mu}$ in $\mathcal{W}^{N\times L}\times H^{N\times L}\times\mathcal{B}^{N\times L}\times\mathcal{Y}^{L}$ by
$\|(w,h,b,y)\|_{\lambda,\mu}=\sum_{l=1}^{L}[\lambda(\|w^{l}\|^{q}_{\mathcal{W}}+\|h^{l}\|^{2}_{H}+\mathcal{D}_{\text{BC}}(b^{l}))+\mu\|y^{l}\|^{r}_{\mathcal{Y}}]$
for $(w,h,b,y)\in\mathcal{W}^{N\times L}\times H^{N\times L}\times\mathcal{B}^{N\times L}\times\mathcal{Y}^{L}$. Letting $\mathcal{R}$ as in Assumption 2, vi), minimization problem (2) may be equivalently rewritten by
$\min_{(\varphi,u,\theta,u_{0},g)\in\mathbf{D}(G)}\|G(\varphi,u,\theta,u_{0},g)-(0,0,0,y)\|_{\lambda,\mu}+\mathcal{R}(\varphi,u,\theta,u_{0},g).$ ($\mathcal{P}^{\prime}$)
Note that problem ($\mathcal{P}^{\prime}$) is in canonical form as the sum of a data-fidelity term and a regularization functional where $G$, given in (61), is the forward operator and $(0,0,0,y)\in\mathcal{W}^{N\times L}\times H^{N\times L}\times\mathcal{B}^{N\times L}\times\mathcal{Y}^{L}$ the measured data. We prove that problem ($\mathcal{P}^{\prime}$) admits a solution in $\mathbf{D}(G)$. If the forward operator $G$ is weakly closed then problem ($\mathcal{P}^{\prime}$) admits a minimizer due to the direct method (see e.g. [100, Chapter 3]) and Assumption 2, vi). The idea is to choose a minimizing sequence, which certainly, for indices large enough is bounded by coercivity of the regularizer, the norm in $H$ and the discrepancy term (together with boundedness of the trace map), thus, attaining a weakly convergent subsequence. Employing weak closedness of $G$, weak lower semicontinuity of the norms, the regularizing term and the discrepancy term (due to Assumption 2, i) and Lemma 44) we derive that the limit of this subsequence is a solution of the minimization problem ($\mathcal{P}^{\prime}$).
Thus, it remains to verify weak closedness of the operator $G$. This is obviously equivalent and reduces to showing weak closedness of the operators $G^{l}$ for $1\leq l\leq L$. For weak closedness of $G^{l}$ it suffices to verify that
  1. I.
    $(\varphi_{n}^{l},(u_{k}^{l})_{1\leq k\leq N},\theta_{n})\mapsto\frac{\partial}{\partial t}u_{n}^{l}-F_{n}(t,(u^{l}_{k})_{1\leq k\leq N},\varphi^{l}_{n})-f_{\theta_{n},n}(t,(\mathcal{J}_{\kappa}u^{l}_{k})_{1\leq k\leq N})$
  2. II.
    $(u_{n}^{l},u_{0,n}^{l})\mapsto u_{n}^{l}(0)-u_{0,n}^{l}$
  3. III.
    $u^{l}=(u_{n}^{l})_{1\leq n\leq N}\mapsto K^{m}u^{l}$
  4. IV.
    $(u^{l},g^{l})\mapsto\gamma(u^{l})-g^{l}$
are weakly closed in $\mathbf{D}(G)$. The weak closedness in III. and IV. follows immediately by weak-weak continuity of $K^{m}$ and continuity of $\gamma$ assumed in Assumption 2. In view of I. it suffices to verify weak closedness of the differential operator $\frac{\partial}{\partial t}:\mathcal{V}\to\mathcal{W}$ as the map $(\theta_{n},v,\varphi)\mapsto F_{n}(v_{1},\dots,v_{N},\varphi)+f_{\theta_{n},n}(\mathcal{J}_{\kappa}v_{1},\dots,\mathcal{J}_{\kappa}v_{N})\in\mathcal{W}$ for $(\theta_{n},v,\varphi)\in\Theta_{n}^{m}\times\mathcal{V}^{N}\times X_{\varphi}$ is weakly closed by Lemma 43 and Assumption 4, ii). For weak closedness of $\frac{\partial}{\partial t}:\mathcal{V}\to\mathcal{W}$ recall Assumption 2, ii) that $\tilde{V}\hookrightarrow W$, and iii) that $\mathcal{V}=L^{p}(0,T;V)\cap W^{1,p,p}(0,T;\tilde{V}),\mathcal{W}=L^{q}(0,T;W)$ with some $p\geq q$. Let $(u_{m})_{m}\subseteq\mathcal{V}$ such that $u_{m}\rightharpoonup u\in\mathcal{V}$ and $\frac{\partial}{\partial t}u_{m}\rightharpoonup v\in\mathcal{W}$. As $\frac{\partial}{\partial t}u_{m}\rightharpoonup\frac{\partial}{\partial t}u\in L^{p}(0,T;\tilde{V})\hookrightarrow\mathcal{W}$ it follows immediately that $\frac{\partial}{\partial t}u=v$, concluding weak closedness of the temporal derivative. For II., employing the embedding $\mathcal{V}\hookrightarrow\mathcal{C}(0,T;H)$ we have that the map $(\cdot)_{t=0}:\mathcal{V}\to H$ with $u\mapsto u(0)$ is weakly closed due to
$\|u(0)\|_{H}\leq\sup_{0\leq t\leq T}\|u(t)\|_{H}\leq c\|u\|_{\mathcal{V}}.$
Thus, problem ($\mathcal{P}^{\prime}$) admits a solution in $\mathbf{D}(G)$ and we conclude wellposedness of problem (2) under the Assumptions 2 to 4.
Appendix D Proof of Proposition 1
In this section we sketch the proof for the linear example in Proposition 1 showcasing our main results for the sake of completeness. For that, once Assumption 2 - 5 are verified to hold true, it follows by application of Proposition 29 and Theorem 30 under suitable choice of regularization parameters depending on the noise of the measurement data and Proposition 22.
[…]
$X_{\varphi}^{L}\times\mathcal{V}^{L}\times\Theta^{m}\ni(\varphi,u,\theta)\mapsto\sum_{l=1}^{L}(\|\varphi^{l}\|_{L^{2}(\Omega)}^{2}+\|u^{l}\|_{\mathcal{V}}^{2})+\|f_{\theta}\|_{L^{2}(U)}^{2}+\|\nabla f_{\theta}\|_{L^{\infty}(U)}+\nu^{m}\|\theta\|,$
for a sufficiently large interval $U$ and regularization parameters $\nu^{m}>0$, is coercive and weakly lower semicontinuous (which follows by Proposition 20). Note that $N=1$. Furthermore, the choice of the image space as $W=L^{2}(\Omega)$ is justified since $\partial_{t}u-\varphi\cdot\nabla u\in L^{2}(\Omega)$ for $\varphi\in H^{1}(\Omega)\hookrightarrow L^{\infty}(\Omega)$ and $u\in\mathcal{V}=W^{1,2,2}(0,T;V)$ as $\partial_{t}u\in L^{2}(\Omega)$ and $\nabla u\in L^{2}(\Omega)$ by $u\in\mathcal{V}=W^{1,2,2}(0,T;V)$. The inclusion for parameterized nonlinearities $f_{\theta}(u)\in L^{2}(\Omega)$ is fulfilled by the following considerations.
Ad Assumption 3: The extendability to a well-defined Nemytskii operator and continuity property of the parameterized nonlinearities follow as mentioned in the paragraph right before Proposition 19 by [1, Lemma 4, Lemma 5] under the regularity condition in Assumption 5, i) that is addressed below.
[…]
Appendix E Example on choice of regularization
In this section, based on a simple example, we show that choosing a $W^{1,\infty}$-type norm in accordance with (3), as opposed to an $L^{p}$-type norm, can indeed be necessary in general for the identification of a hidden physics component. For that, we consider a one dimensional time-independent equation on the unit interval with known physical term $F$ that depends on the first spatial derivative of the state. Furthermore, we suppose that the state $u$ is given via the full measurement operator (e.g. equal to the identity with noiseless measurement data) and is approximated by a known sequence of states $(u_{m})_{m}$. Note that, although this setup is simpler than the general one considered in this work, the example shows that already in this simplified situation appropriate regularization, as discussed, matters. Considering classes of continuous functions parameterized by sets $\Theta^{m}$ that can approximate more and more complicated functions better for increasing $m\in\mathbb{N}$ (according to Assumption 5, ii) and iv)), we provide an example where solutions $(f_{\theta^{m}})_{m}$ to
$\min_{\theta\in\Theta^{m}}\|f_{\theta}\|_{L^{2}(0,1)}^{2}+\lambda^{m}\|f_{\theta}(u_{m})+F(u_{m}^{\prime})\|^{2}_{L^{2}(0,1)}+\mu^{m}\|\theta\|$ ($\mathcal{P}^{m}_{L^{2}}$)
with $\theta^{m}\in\Theta^{m}$, do not converge to the unique solution $f^{\dagger}\in L^{2}(0,1)$ of
$\min_{f\in L^{2}(0,1)}\|f\|_{L^{2}(0,1)}^{2}\quad\text{s.t.}~f(u)=g.$ ($\mathcal{P}^{\dagger}_{L^{2}}$)
It is important to note that the constraint in ($\mathcal{P}^{\dagger}_{L^{2}}$) is not well-defined for general $f,u,g\in L^{2}(0,1)$ but for the setup discussed below it is. In view of well-definedness of the minimization problem ($\mathcal{P}^{m}_{L^{2}}$) we refer to the considerations below. Here $(\lambda^{m})_{m},(\mu^{m})_{m}\subset\mathbb{R}_{+}$ are suitable regularization parameters, the former monotonically increasing and divergent, and the latter a zero sequence. Finally, in contrast to above situation we show that penalizing the gradient term similarly as in (3) allows to identify the unique solution $f^{\dagger}$ of
$\min_{f\in W^{1,\infty}(0,1)}\|f\|_{L^{2}(0,1)}^{2}+\|\nabla f\|_{L^{\infty}(0,1)}\quad\text{s.t.}~f(u)=g$ ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$)
as limit of solutions $(f_{\theta^{m}})_{m}$ for $\theta^{m}\in\Theta^{m}$ to
$\min_{\theta\in\Theta^{m}}\|f_{\theta}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{\theta}\|_{L^{\infty}(0,1)}+\lambda^{m}\|f_{\theta}(u_{m})+F(u_{m}^{\prime})\|^{2}_{L^{2}(0,1)}+\mu^{m}\|\theta\|.$ ($\mathcal{P}^{m}_{W^{1,\infty}}$)
Concretely, we choose the known physical term $F$ by
$\mathbb{R}\ni v\mapsto F(v)=\begin{cases}\left(1-4(v-1)^{2}\right)^{1/2},&\text{if}~1/2\leq v\leq 3/2,\\ 0,&\text{otherwise.}\end{cases}$
Furthermore, we suppose that the state is given by the identity map $u(x)=x$ for $x\in[0,1]$ and is approximated by the sequence of states $(u_{m})_{m}$ given for $m\in\mathbb{N}$ by
$u_{m}(x)=x+\frac{1}{4\pi m}\sin(2\pi mx)\quad\text{for}~x\in[0,1].$
The sequence $(u_{m})_{m}$ converges to $u$ in $L^{\infty}(0,1)$ since $\|u_{m}-u\|_{L^{\infty}(0,1)}=(4\pi m)^{-1}$ for $m\in\mathbb{N}$ (in fact, even weakly in $H^{1}(0,1)$). Note that $u_{m}:[0,1]\to[0,1]$ is bijective and even diffeomorphic since $u_{m}^{\prime}(x)=1+\cos(2\pi mx)/2\in[1/2,3/2]$ for all $x\in[0,1]$ and $m\in\mathbb{N}$. The sequence $(g_{m})_{m}$ with $g_{m}:=-F(u_{m}^{\prime})$, i.e., for $m\in\mathbb{N}$
$g_{m}(x)=\sin(2\pi mx)\quad\text{for}~x\in[0,1],$
converges to $g=0$ weakly in $L^{2}(0,1)$. The regularization parameters are chosen by $\lambda^{m}=\lambda_{0}m^{1/2}$ for $m\in\mathbb{N}$ and fixed $\lambda_{0}>0$ whereas $(\mu^{m})_{m}$ is a zero sequence satisfying the following conditions. We assume that the zero function can be parameterized for any $m\in\mathbb{N}$ with suitable parameters $\hat{\theta}^{m}\in\Theta^{m}$. Furthermore, suppose that the periodic functions $h_{m}:=g_{m}\circ u_{m}^{-1}$ (which have period $1/m$) can be parameterized with parameters $\tilde{\theta}^{m}\in\Theta^{m}$ for $m\in\mathbb{N}$. Then we choose $(\mu^{m})_{m}$ such that both $(\mu^{m}\|\hat{\theta}^{m}\|)_{m}$ and $(\mu^{m}\|\tilde{\theta}^{m}\|)_{m}$ are zero sequences. Note that the PDE data term $\|f(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}$ is well-defined for any $f\in L^{2}(0,1)$. Indeed, since $u_{m}$ is a diffeomorphism of the interval $[0,1]$ onto itself, a change of variables yields
$\|f(u_{m})\|_{L^{2}(0,1)}^{2}=\int_{0}^{1}|f(x)|^{2}(u_{m}^{\prime}(u_{m}^{-1}(x)))^{-1}\mathop{}\!\mathrm{d}x.$ (62)
Using that $1/2\leq u_{m}^{\prime}(x)\leq 3/2$ for $x\in[0,1]$ and $m\in\mathbb{N}$ implies
$\frac{2}{3}\|f\|_{L^{2}(0,1)}^{2}\leq\|f(u_{m})\|_{L^{2}(0,1)}^{2}\leq 2\|f\|_{L^{2}(0,1)}^{2}.$ (63)
Note that by similar arguments we derive for $h_{m}=g_{m}\circ u_{m}^{-1}$ that
$\|h_{m}\|_{L^{2}(0,1)}^{2}\leq\frac{3}{2}\|g_{m}\|_{L^{2}(0,1)}^{2}=3/4$ (64)
using that $\|g_{m}\|_{L^{2}(0,1)}^{2}=1/2$ for $m\in\mathbb{N}$. For $f\in W^{1,\infty}(0,1)$ well-definedness follows from the embedding $W^{1,\infty}(0,1)\hookrightarrow\mathcal{C}(0,1)$. Furthermore, note that well-definedness of ($\mathcal{P}^{m}_{W^{1,\infty}}$) follows by the considerations in this work, whereas well-definedness of ($\mathcal{P}^{m}_{L^{2}}$) is a consequence of the direct method together with (62) applied to parameterizations instead of $f$ and continuity with respect to the parameterization similar as in Proposition 19.
We now start by considering the $L^{2}$-regularized problem. For $m\in\mathbb{N}$ the function $h_{m}=g_{m}\circ u_{m}^{-1}$ is representable by $f_{\tilde{\theta}^{m}}$ with $\tilde{\theta}^{m}\in\Theta^{m}$ such that the objective functional of ($\mathcal{P}^{m}_{L^{2}}$) in an optimum can be estimated by
$\|f_{\theta^{m}}\|_{L^{2}(0,1)}^{2}+\lambda_{0}m^{1/2}\|f_{\theta^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\theta^{m}\|\\ \leq\|f_{\tilde{\theta}^{m}}\|_{L^{2}(0,1)}^{2}+\lambda_{0}m^{1/2}\|\underbrace{f_{\tilde{\theta}^{m}}(u_{m})-g_{m}}_{=0}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\tilde{\theta}^{m}\|.$
Due to the choice of $(\mu^{m})_{m}$ and (64) the right hand side is uniformly bounded for sufficiently large $m\in\mathbb{N}$. Thus, it follows that $\|f_{\theta^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}$ converges to zero as $m\to\infty$. Hence, there exists a constant $c>0$ such that by the reverse triangle inequality for sufficiently large $m\in\mathbb{N}$
$\|f_{\theta^{m}}(u_{m})\|_{L^{2}(0,1)}\geq\underbrace{\|g_{m}\|_{L^{2}(0,1)}}_{=2^{-1/2}}-\|f_{\theta^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}\geq c>0.$ (65)
Thus, the sequence $(f_{\theta^{m}})_{m}$ cannot converge to zero in $L^{2}(0,1)$ as (63) would immediately lead to a contradiction to (65). However, the constant zero function is the unique solution to ($\mathcal{P}^{\dagger}_{L^{2}}$) (of course up to representatives in the Lebesgue sense). As a consequence, the reconstruction of the hidden physical term fails. In fact, one can show for the minimization problem similar to ($\mathcal{P}^{m}_{L^{2}}$) considered over general $L^{2}(0,1)$-functions (and without parameter regularization) that by analyzing the first variation of the resulting strictly convex objective functional under (62), the corresponding minimizer $f_{m}$ is given by
$f_{m}(x)=\lambda_{0}\left(\lambda_{0}+m^{-1/2}u_{m}^{\prime}(u_{m}^{-1}(x))\right)^{-1}g_{m}(u_{m}^{-1}(x))\quad\text{for}~x\in[0,1],$
whose $L^{2}(0,1)$-norm can be shown to be uniformly bounded from below. Another interesting point is that the $(f_{m})_{m}$ are continuously differentiable in the open unit interval and $\lim_{m\to\infty}\|\nabla f_{m}\|_{L^{\infty}(0,1)}=\infty$. Thus, also in a purely analytic setup, the reconstruction of the hidden physics fails.
Let us consider next the $W^{1,\infty}$-type regularized problems. Problem ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$) attains the unique solution $f^{\dagger}=0$. We argue that $f^{\dagger}$ is recovered by $(f_{\theta^{m}})_{m}$ solving ($\mathcal{P}^{m}_{W^{1,\infty}}$). For that, we first estimate the objective function of ($\mathcal{P}^{m}_{W^{1,\infty}}$) in the optimum by its value at $f_{\hat{\theta}^{m}}$ representing the constant zero function, yielding
$\|f_{\theta^{m}}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}+\lambda_{0}m^{1/2}\|f_{\theta^{m}}(u_{m})-g_{m}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\theta^{m}\|\\ \leq\lambda_{0}m^{1/2}\|g_{m}\|_{L^{2}(0,1)}^{2}+\mu^{m}\|\hat{\theta}^{m}\|=\lambda_{0}m^{1/2}/2+\mu^{m}\|\hat{\theta}^{m}\|.$ (66)
We derive by (66) that
$\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}\leq\lambda_{0}m^{1/2}/2+\mu^{m}\|\hat{\theta}^{m}\|,$
which implies the existence of a constant $c>0$ such that $\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}\leq cm^{1/2}$ for $m\in\mathbb{N}$ since $(\mu^{m}\|\hat{\theta}^{m}\|)_{m}$ is a zero sequence. Under this constraint on the gradient, the best approximation of the data term is bounded from below by
$\inf_{\begin{subarray}{c}\theta\in\Theta^{m},\\ \|\nabla f_{\theta}\|_{L^{\infty}(0,1)}\leq c\sqrt{m}\end{subarray}}\|f_{\theta}(u_{m})-g_{m}\|_{L^{2}(0,1)}\geq\inf_{\begin{subarray}{c}f\in W^{1,\infty}(0,1),\\ \|\nabla f\|_{L^{\infty}(0,1)}\leq c\sqrt{m}\end{subarray}}\|f(u_{m})-g_{m}\|_{L^{2}(0,1)}.$ (67)
We analyze the right hand side in more detail. By the mean value theorem it holds true for $f\in W^{1,\infty}(0,1)$ that
$\|f(u)-f(u_{m})\|_{L^{2}(0,1)}\leq\|\nabla f\|_{L^{\infty}(0,1)}\|u-u_{m}\|_{L^{2}(0,1)}=\tilde{c}\|\nabla f\|_{L^{\infty}(0,1)}m^{-1}$
with $\tilde{c}=32^{-1/2}\pi^{-1}$, such that (67) can be estimated from below by
$\inf_{\begin{subarray}{c}f\in W^{1,\infty}(0,1),\\ \|\nabla f\|_{L^{\infty}(0,1)}\leq cm^{1/2}\end{subarray}}\|f-g_{m}\|_{L^{2}(0,1)}-\tilde{c}cm^{-1/2}.$ (68)
Using a scaling argument, the value of the approximation problem in (68) equals
$cm^{1/2}\inf_{\begin{subarray}{c}f\in W^{1,\infty}(0,1),\\ \|\nabla f\|_{L^{\infty}(0,1)}\leq 1\end{subarray}}\|f-c^{-1}m^{-1/2}g_{m}\|_{L^{2}(0,1)}.$ (69)
Due to [18, Theorem 1.1] this problem attains a unique solution $\hat{f}$ which satisfies
$\hat{f}(x)$ $=\max_{y\in\partial A^{+}}\hat{f}(y)-|x-y|\qquad\text{for all}~x\in A^{+},$
$\hat{f}(x)$ $=\min_{y\in\partial A^{-}}\hat{f}(y)+|x-y|\qquad\text{for all}~x\in A^{-},$
where $A^{+}=\text{supp}(\max(\hat{f}-c^{-1}m^{-1/2}g_{m},0))$ and $A^{-}=\text{supp}(\min(\hat{f}-c^{-1}m^{-1/2}g_{m},0))$ denote the respective supports, and $\partial A^{+},\partial A^{-}$ the corresponding boundaries. With this, and the fact that $c^{-1}m^{-1/2}g_{m}$ attains $m$ full periods in $[0,1]$, it follows by a symmetry argument that $\hat{f}$ is given by the saw-tooth function with unit slope and same period and sign as $g_{m}$. As a consequence, the term in (69) equals
$cm^{1/2}\left(4m\int_{0}^{1/4m}(c^{-1}m^{-1/2}g_{m}(x)-x)^{2}\mathop{}\!\mathrm{d}x\right)^{1/2}=\left(\frac{1}{2}-\frac{2c}{\pi^{2}m^{1/2}}+\frac{c^{2}}{48m}\right)^{1/2}.$
Combining these arguments, finally, yields by (66) that
$\|f_{\theta^{m}}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}\\ \leq\lambda_{0}m^{1/2}\left[\frac{1}{2}-\left(\left(\frac{1}{2}-\frac{2c}{\pi^{2}m^{1/2}}+\frac{c^{2}}{48m}\right)^{1/2}-\frac{c}{\sqrt{32}\pi m^{1/2}}\right)^{2}\right]+\mu^{m}\|\hat{\theta}^{m}\|.$
The right hand side converges to the constant $\frac{c\lambda_{0}}{4\pi}\left(\frac{8}{\pi}+1\right)$ as $m\to\infty$. Thus, there exist $M\in\mathbb{N}$ and a constant $\alpha>0$ such that for all $m\geq M$ it holds true that
$\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}\leq\alpha.$
Repeating the arguments starting from (67), but now with the refined bound $\|\nabla f_{\theta}\|_{L^{\infty}(0,1)}\leq\alpha$ on the gradients, yields for $m\geq M$ that
$\|f_{\theta^{m}}\|_{L^{2}(0,1)}^{2}+\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}\\ \leq\lambda_{0}m^{1/2}\left[\frac{1}{2}-\left(\left(\frac{1}{2}-\frac{2\alpha}{\pi^{2}m}+\frac{\alpha^{2}}{48m^{2}}\right)^{1/2}-\frac{\alpha}{\sqrt{32}\pi m}\right)^{2}\right]+\mu^{m}\|\hat{\theta}^{m}\|,$
which converges to zero as $m\to\infty$. As a consequence, it holds true that
$\lim_{m\to\infty}\|\nabla f_{\theta^{m}}\|_{L^{\infty}(0,1)}=0\quad\text{and}\quad\lim_{m\to\infty}\|f_{\theta^{m}}\|_{L^{2}(0,1)}=0,$
implying that $(f_{\theta^{m}})_{m}$ converges uniformly to zero, the unique solution to ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$).
Numerical experiments.
To show that the above counterexample is also observable in practice, we implemented it numerically using two-hidden layers neural networks with input- and output dimension one. The first layer consists of $10$ nodes and is sine-activated with frequency $2$ (to enable the representation of high frequency oscillations). The second layer consists of $m\in\mathbb{N}$ nodes and is ReLU-activated. The network training for ($\mathcal{P}^{m}_{L^{2}}$) and ($\mathcal{P}^{m}_{W^{1,\infty}}$) was performed over $1500$ epochs for $m\in\{10,100,1000\}$ using the Adam optimizer with learning rate $3\cdot 10^{-3}$ and weight decay $\mu^{m}=0.1\cdot m^{-1}$ for $200$ uniformly sampled training points in the unit interval. Furthermore, we chose $\lambda_{0}=1/2$. The gradient $\nabla f_{\theta}$ in view of the loss calculation of ($\mathcal{P}^{m}_{W^{1,\infty}}$) is approximated via $30$ uniformly sampled approximations of the gradient using finite differences. An illustration of the results is given in Figure 1.
Figure 1: Numerical example with different regularizations. Top: $L^{2}$ regularization, bottom: $W^{1,\infty}$-type regularization.
In Figure 1 the first and second line of subplots correspond to the considerations on ($\mathcal{P}^{m}_{L^{2}}$) and ($\mathcal{P}^{m}_{W^{1,\infty}}$), respectively. The first column of subplots depicts approximations of hidden physics after $1500$ epochs for the different $m$ above. In the second column the corresponding data term losses are plotted. Finally, in the third column the $L^{2}$-deviation of the approximated hidden physics from the unique solutions of the limit problem ($\mathcal{P}^{\dagger}_{L^{2}}$) and ($\mathcal{P}^{\dagger}_{W^{1,\infty}}$), the zero function, is depicted, respectively. The main conclusion is that under $L^{2}$-regularization the physical term reconstruction fails, whereas with the $W^{1,\infty}$-type penalty the hidden physics is identified correctly. For the latter it is further clear that the unknown term is approximated better for increasing $m$ and the parameterized approximations are relatively flat due to the gradient penalty. Note that the data term loss of ($\mathcal{P}^{m}_{W^{1,\infty}}$) stagnates as predicted by our analysis above.
Notation
We briefly summarize the function spaces and embeddings, which form the basis of this work. A list of the symbols and abbreviations used is also provided.
General spaces
$\Omega$. space domain
$L^{p}(\Omega)$. Lebesgue space
$W^{\kappa,p}(\Omega)$. Sobolev space
$L^{p}(I;X),\mathcal{C}(I;X)$. Bochner space [98, Section 1.5]
$W^{1,p,q}(I;X)$. Sobolev-Bochner space [98, Section 7.1]
$W^{1,\infty}_{loc}(\mathbb{R}^{D})$. space of locally $W^{1,\infty}$-regular functions on $\mathbb{R}^{D}$
Function spaces (cf. Section 2)
For some $1\leq p,q,r,s<\infty$ with $p\geq q,p\geq s$:
$V$. state space
$\tilde{V}$. space of time derivative
[…]
$V_{k},V_{k}^{\times}$, $\mathcal{V}_{k},\mathcal{V}_{k}^{\times}$. space of spatial derivatives and dynamic extensions
$W$, $\mathcal{W}=L^{q}(0,T;W)$. image space and dynamic extension
$Y$, $\mathcal{Y}=L^{r}(0,T;Y)$. observation space and dynamic extension
$B$, $\mathcal{B}=L^{s}(0,T;B)$. boundary trace space and dynamic extension
$H$. initial trace space
$X_{\varphi}$. parameter space
$\Theta_{n}^{m}$. parameter sets
$\mathcal{F}_{n}^{m}$. approximation classes
Embeddings (cf. Assumption 2, ii))
$V\hookrightarrow H\hookrightarrow\tilde{V}\hookrightarrow W$
$L^{\hat{p}}(\Omega)\hookrightarrow V_{k}\hookrightarrow L^{\hat{q}}(\Omega)$ for $1\leq k\leq\kappa$
$W^{\kappa,\hat{p}}(\Omega)\hookrightarrow\tilde{V}$ or $\tilde{V}\hookrightarrow W^{\kappa,\hat{p}}(\Omega)$
$L^{\hat{q}}(\Omega)\hookrightarrow W$ for some $1\leq\hat{q}\leq\hat{p}$$\infty$
$V\hookrightarrow Y$
[…]
Notational conventions
$\mathcal{D}_{BC}$. discrepancy term for boundary conditions
$f,f_{\theta}$. hidden physics component and parametrization
$F$. known physical model
$\gamma$. boundary trace map
[…]
$J^{l}$, $\mathcal{J}_{\kappa}$. Jacobian mapping and derivative operator
$K^{\dagger},K^{m}$. measurement operators
$\mathcal{N}_{\theta}$. feed forward neural network
$\otimes$. Cartesian product of spaces
$\varphi$. physical parameter
$\mathcal{R}^{\dagger},\mathcal{R}_{m},\mathcal{R}_{0}$. regularization functionals
$u,u_{0}$. state, initial condition
$y,y^{m}$. measurement data